### v3

Abstract
In the past several years, the last-iterate convergence of the Stochastic Gradient Descent (SGD) algorithm has triggered people’s interest due to its good performance in practice but lack of theoretical understanding. For Lipschitz convex functions, different works have established the optimal $O(\log(1/\delta)\log T/\sqrt{T})$ or $O(\sqrt{\log(1/\delta)/T})$ high-probability convergence rates for the final iterate, where $T$ is the time horizon and $\delta$ is the failure probability. However, to prove these bounds, all the existing works are either limited to compact domains or require almost surely bounded noise. It is natural to ask whether the last iterate of SGD can still guarantee the optimal convergence rate but without these two restrictive assumptions. Besides this important question, there are still lots of theoretical problems lacking an answer. For example, compared with the last-iterate convergence of SGD for non-smooth problems, only few results for smooth optimization have yet been developed. Additionally, the existing results are all limited to a non-composite objective and the standard Euclidean norm. It still remains unclear whether the last-iterate convergence can be provably extended to wider composite optimization and non-Euclidean norms. In this work, to address the issues mentioned above, we revisit the last-iterate convergence of stochastic gradient methods and provide the first unified way to prove the convergence rates both in expectation and in high probability to accommodate general domains, composite objectives, non-Euclidean norms, Lipschitz conditions, smoothness, and (strong) convexity simultaneously. Additionally, we extend our analysis to obtain the last-iterate convergence under heavy-tailed noise.
1 Introduction
In this paper, we consider the constrained composite optimization problem $\min_{x\in\mathcal{X}}F(x)\coloneqq f(x)+h(x)$ where both $f(x)$ and $h(x)$ are convex (but possibly satisfying additional conditions such as strong convexity, smoothness, etc.) and $\mathcal{X}\subseteq\mathbb{R}^{d}$ is a nonempty closed convex set. Since a true gradient is computationally prohibitive to obtain (e.g., large-scale machine learning tasks) or even infeasible to access (e.g., streaming data), the classic Stochastic Gradient Descent (SGD) (Robbins and Monro, 1951) algorithm has emerged to be the gold standard for a light-weight yet effective computational procedure commonly adopted in production for the majority of machine learning tasks: SGD only requires a stochastic first-order oracle $\widehat{\partial}f(x)$ satisfying $\mathbb{\mathbb{E}}[\widehat{\partial}f(x)\mid x]\in\partial f(x)$ where $\partial f(x)$ denotes the set of subgradients at $x$ and guarantees provable convergence under certain conditions (e.g., Lipschitz condition for $f(x)$ and finite variance on the stochastic oracle).
A particularly important problem in this area is to understand the last-iterate convergence of SGD, which has been motivated by experimental studies suggesting that returning the final iterate of SGD (or sometimes the average of the last few iterates) – rather than a running average – often yields a solution that works well in practice (e.g., Shalev-Shwartz et al. (2007)). As such, a fruitful line of literature (Rakhlin et al., 2011; Shamir and Zhang, 2013; Harvey et al., 2019b; Orabona, 2020; Jain et al., 2021) developed an extensive theoretical understanding of the non-asymptotic last-iterate convergence rate. Loosely speaking, two optimal upper bounds, $\widetilde{O}(1/\sqrt{T})$ for Lipschitz convex functions and $\widetilde{O}(1/T)$ for Lipschitz strongly convex functions, have been established for both in-expectation and high-probability convergence when $h(x)=0$ (see Subsection 1.2 for a detailed discussion). However, to prove the high-probability rates, existing works rely on restrictive assumptions: compact domains or almost surely bounded noise (or both), which can simplify the analysis but are unrealistic in lots of problems. Until today, whether these two assumptions can be relaxed simultaneously or not still remains unclear. Naturally, we want to ask the following question:
Q1: Is it possible to prove the high-probability last-iterate convergence of SGD for Lipschitz (strongly) convex functions without the compact domain assumption and beyond the bounded noise?
Compared with the fast development of non-smooth problems, the understanding of the last-iterate convergence of SGD for smooth problems (i.e., the gradients of $f(x)$ are Lipschitz) is much slower. The best in-expectation bound for smooth convex optimization under $\mathcal{X}=\mathbb{R}^{d}$ until now is still $O(1/\sqrt[3]{T})$ due to Moulines and Bach (2011), which is far from the optimal rate $O(1/\sqrt{T})$ of the averaging output (Lan, 2020). However, temporarily suppose the domain is compact, one can immediately improve the rate from $O(1/\sqrt[3]{T})$ to $\widetilde{O}(1/\sqrt{T})$ by noticing that we can reduce the smooth problem to the Lipschitz problem [^1] To see why gradients are bounded in this case, we first fix a point $x_{0}$ in the domain. Then by smoothness, there is $\left\|\nabla f(x)-\nabla f(x_{0})\right\|_{2}=O\left(\left\|x-x_{0}\right\|_{2}\right)$ for any other point $x$, which immediately implies $\left\|\nabla f(x)\right\|_{2}=O\left(\left\|x-x_{0}\right\|_{2}+\left\|\nabla f(x_{0})\right\|_{2}\right)=O\left(D+\left\|\nabla f(x_{0})\right\|_{2}\right)$ where $D$ is the domain diameter. and use the known bounds from non-smooth convex optimization. Hence, one may expect the last-iterate convergence rate of SGD for smooth convex optimization should still be $O(1/\sqrt{T})$ for any kind of domain. If one further considers smooth and strongly convex problems, as far as we know, there is no formal result has been established for the final iterate of SGD in a general domain except for the $O(1/T)$ in-expectation rate when $\mathcal{X}=\mathbb{R}^{d}$ under the PL-condition (which is known as a relaxation for strong convexity) (Gower et al., 2021; Khaled and Richtárik, 2023). The above discussion thereby leads us to the second main question:
Q2: Does the last iterate of SGD provably converge in the rate of $O(1/\sqrt{T})$ for smooth and convex functions and $O(1/T)$ for smooth and strongly convex functions in a general domain?
Besides the two aforementioned questions, there are still several important missing parts. First, recalling that our original goal is to optimize the composite objective $F(x)=f(x)+h(x)$, it is still unclear whether – and if so, how – the last-iterate convergence of this harder problem can be proved. Moreover, the previous works are limited to the standard Euclidean norm. Whereas, in lots of specialized tasks, it may be beneficial to employ a general norm instead of the $\ell_{2}$ norm to capture the non-Euclidean structure. However, whether this extension can be done remains open. Additionally, the proof techniques in the existing works vary in different settings, which builds a barrier for researchers to better understand the convergence of the last iterate of SGD. Motivated by these challenges, we would like to ask the final question:
Q3: Is there a unified way to analyze the last-iterate convergence of stochastic gradient methods both in expectation and in high probability to accommodate general domains, composite objectives, non-Euclidean norms, Lipschitz conditions, smoothness, and (strong) convexity at once?
1.1 Our Contributions
We provide affirmative answers to the above three questions and establish several new results by revisiting a simple algorithm, Composite Stochastic Mirror Descent (CSMD) (Duchi et al., 2010), which is based on the famous Mirror Descent (MD) algorithm (Nemirovski and Yudin, 1983; Beck and Teboulle, 2003) and includes SGD as a special case. Specifically, our contributions are as follows.
  • We establish the first high-probability convergence result for the last iterate of CSMD in general domains under sub-Gaussian noise to answer Q1 affirmatively.
  • We prove the last iterate of CSMD can converge in the rate of $O(1/\sqrt{T})$ for smooth convex optimization and $\widetilde{O}(1/T)$ for smooth strongly convex problems both in expectation and in high probability for any general domain $\mathcal{X}$, hence resolving Q2.
  • We present a simple unified analysis that differs from the prior works and can be directly applied to various scenarios simultaneously, thus leading to a positive answer to Q3.
By extending the proof idea:
  • In Section 5, we provide the first in-expectation convergence bound under heavy-tailed noise for the last iterate of the CSMD algorithm.
  • In Section 6, we prove the first high-probability convergence rate under sub-Weibull noise for the last iterate of the CSMD algorithm.
1.2 Related Work
We review the literature related to the last-iterate convergence of plain stochastic gradient methods [^2] To clarify, we mean the algorithm does not contain momentum or averaging operations. measured by the function value gap (see Subsection 2.1 for why we use this criterion) for both Lipschitz and smooth (strongly) convex optimization. We only focus on the algorithms without momentum or averaging since it is already known that, without further special assumptions, both operations cannot help to improve the lower order term $O(1/\sqrt{T})$ for general convex functions and $O(1/T)$ for strongly convex functions. For the last iterate of accelerated or averaging based stochastic gradient methods, we refer the reader to Nesterov and Shikhman (2015); Lan (2020); Orabona and Pál (2021) for in-expectation rates and Davis and Drusvyatskiy (2020); Gorbunov et al. (2020); Liu et al. (2023b); Sadiev et al. (2023) for high-probability bounds. As for the last iterate of stochastic gradient methods for structured problems (e.g., linear regression), the reader can refer to Lei and Zhou (2017); Ge et al. (2019); Varre et al. (2021); Pan et al. (2022); Wu et al. (2022) for recent progress.
Last iterate for Lipschitz (strongly) convex functions: Rakhlin et al. (2011) is the first to show an in-expectation $O(1/T)$ convergence rate for strongly convex functions. But such a bound is obtained under the additional assumption, smoothness with respect to optimum [^3] This means $\exists L>0$ such that $f(x)-f(x^{*})\leq\frac{L}{2}\|x-x^{*}\|^{2},\forall x\in\mathcal{X}$ where $x^{*}\in\mathrm{argmin}_{x\in\mathcal{X}}f(x)$., meaning their result does not hold in general. Later on, Shamir and Zhang (2013) proves the first in-expectation last-iterate rates $O(\log T/\sqrt{T})$ and $O(\log T/T)$ for convex and strongly convex objectives, respectively. The high-probability bounds turn out to be much harder than the in-expectation rates. After several years, Harvey et al. (2019b) is the first to establish a high-probability bound in the rate of $O(\log(1/\delta)\log T/\sqrt{T})$ and $O(\log(1/\delta)\log T/T)$ for convex and strongly convex problems where $\delta$ is the probability of failure. Afterward, Jain et al. (2021) improves the previous two rates to $O(\sqrt{\log(1/\delta)/T})$ and $O(\log(1/\delta)/T)$ but with a non-standard step size schedule. They also prove the rates $O(1/\sqrt{T})$ and $O(1/T)$ in expectation under the new step size.
However, a main drawback for the general convex case in all the above papers is requiring a compact domain. To our best knowledge, Orabona (2020) is the first and the only work showing how to shave off this restriction, and thereby obtains an $O(\log T/\sqrt{T})$ rate in expectation for general domains yet it is unclear whether his proof can be extended to the high-probability case or not. Until recently, Zamani and Glineur (2025) exhibits a new proof on how to obtain the convergence rate for the last iterate but only for the deterministic case. Lastly, we would like to mention that all of these prior results are built for a non-composite objective $f(x)$ with the standard Euclidean norm.
Last iterate for smooth (strongly) convex functions: Compared with Lipschitz problems, much less work is done for smooth optimization. As far as we know, the only result showing a non-asymptotic rate for smooth convex functions dates back to Moulines and Bach (2011), in which the authors prove that the last iterate of SGD on $\mathbb{R}^{d}$ enjoys an in-expectation rate $O(1/\sqrt[3]{T})$ under additional restrictive assumptions (e.g., mean squared smoothness). As for the strongly convex case, the $O(1/T)$ rate in expectation under the PL-condition (which is known as a relaxation for strong convexity) has been established, but only for non-composite optimization under the Euclidean norm on the domain $\mathcal{X}=\mathbb{R}^{d}$ (Gower et al., 2021; Khaled and Richtárik, 2023).
Lower bounds for last iterate: Under the requirement $d=T$ where $d$ is the dimension of the problem, Harvey et al. (2019b) is the first to provide lower bounds $\Omega(\log T/\sqrt{T})$ under the step size $\Theta(1/\sqrt{t})$ for non-smooth convex functions and $\Omega(\log T/T)$ under the step size $\Theta(1/t)$ when strong convexity is additionally assumed. Note that these two rates are both proved for deterministic optimization meaning that they can be also applied to the in-expectation lower bounds. Subsequently, when $d<T$ holds, Liu and Lu (2021) extends the above two lower bounds to $\Omega(\log d/\sqrt{T})$ (this bound is also true for the step size $\Theta(1/\sqrt{T})$) and $\Omega(\log d/T)$ under the same step size in Harvey et al. (2019b). As a consequence, lower bounds $\Omega(\log(d\land T)/T)$ and $\Omega(\log(d\land T)/\sqrt{T})$ have been established for both convex and strongly convex problems under the Lipschitz condition. For the high-probability bounds, Harvey et al. (2019b) shows their two deterministic bounds will incur an extra multiplicative factor $\Omega(\log(1/\delta))$, namely, $\Omega(\log(1/\delta)\log T/\sqrt{T})$ and $\Omega(\log(1/\delta)\log T/T)$. However, under more sophisticated designed step sizes, better upper bounds without the $\Omega(\log T)$ factor are possible, for example, see Jain et al. (2021) as mentioned above.
Another highly related work is Liu et al. (2023b), which presents a generic approach to establish the high-probability convergence of the average iterate under sub-Gaussian noise. We will show that their idea can be further used to prove the high-probability convergence for the last iterate.
Additional works about heavy-tailed noise and sub-Weibull noise will be provided in Sections 5 and 6.
2 Preliminaries
Notations: $\mathbb{\mathbb{N}}$ is the set of natural numbers (excluding 0). $\left[d\right]\coloneqq\left\{1,2,\cdots,d\right\}$ for any $d\in\mathbb{\mathbb{N}}$. $a\lor b$ and $a\land b$ are defined as $\max\left\{a,b\right\}$ and $\min\left\{a,b\right\}$, respectively. $\left\langle\cdot,\cdot\right\rangle$ is the standard Euclidean inner product on $\mathbb{R}^{d}$. $\left\|\cdot\right\|$ represents a general norm on $\mathbb{R}^{d}$ and $\left\|\cdot\right\|_{*}$ is its dual norm. Given a set $A\subseteq\mathbb{R}^{d}$, $\mathrm{int}(A)$ stands for its interior points. For a function $f$, $\partial f(x)$ denotes the set of subgradients at $x$.
We focus on the following optimization problem in this work
$\min_{x\in\mathcal{X}}F(x)\coloneqq f(x)+h(x),$
where $f$ and $h$ are both convex. $\mathcal{X}\subseteq\mathrm{int}(\mathrm{dom}(f))\subseteq\mathbb{R}^{d}$ is a closed convex set. The requirement of $\mathcal{X}\subseteq\mathrm{int}(\mathrm{dom}(f))$ is only to guarantee the existence of $\partial f(x)$ for every point $x$ in $\mathcal{X}$ with no special reason. We emphasize that there is no compactness requirement on $\mathcal{X}$. Additionally, given $\psi$ being a differentiable and $1$-strongly convex function with respect to $\left\|\cdot\right\|$ on $\mathcal{X}$ (i.e., $\psi(x)\geq\psi(y)+\left\langle\nabla\psi(y),x-y\right\rangle+\frac{1}{2}\left\|x-y\right\|^{2},\forall x,y\in\mathcal{X}$ [^4] Rigorously speaking, $y$ should be in $\mathrm{int}(\mathcal{X})$. But one can think $\mathcal{X}\subseteq\mathrm{int}(\mathrm{dom}(\psi))$ to avoid this issue.), the Bregman divergence with respect to $\psi$ is defined as $D_{\psi}(x,y)\coloneqq\psi(x)-\psi(y)-\left\langle\nabla\psi(y),x-y\right\rangle$. Throughout this paper, we assume that $\mathrm{argmin}_{x\in\mathcal{X}}h(x)+\left\langle g,x-y\right\rangle+\frac{D_{\psi}(x,y)}{\eta}$ can be solved efficiently for any $g\in\mathbb{R}^{d}$, $y\in\mathcal{X}$, $\eta>0$.
Next, we list the assumptions used in our analysis:
  1. 1.
    Existence of a local minimizer: $\exists x^{*}\in\mathrm{argmin}_{x\in\mathcal{X}}F(x)$ satisfying $F(x^{*})>$-$\infty$. [^5] In fact, this assumption is not necessary for most of our results (except theorems and lemmas in Sections 6 and G) since we can bound $F(x^{T+1})-F(x)$ for any $x\in\mathcal{X}$. However, we keep it here and will use it when stating theorems in the main text. When it is possible, we will drop this assumption in the intermediate lemmas and theorems in the appendix.
  2. 2.
    $(\mu_{f},\mu_{h})$-strongly convex: For $k=f$ and $k=h$, $\exists\mu_{k}\geq 0$ such that $\mu_{k}D_{\psi}(x,y)\leq k(x)-k(y)-\left\langle g,x-y\right\rangle,\forall x,y\in\mathcal{X},g\in\partial k(y)$. Moreover, we assume at least one of $(\mu_{f},\mu_{h})$ is zero.
  3. 3.
    General $(L,M)$-smooth: $\exists L\geq 0,M\geq 0$ such that $f(x)-f(y)-\left\langle g,x-y\right\rangle\leq\frac{L}{2}\left\|x-y\right\|^{2}+M\left\|x-y\right\|,\forall x,y\in\mathcal{X},g\in\partial f(y)$.
  4. 4.
    Unbiased gradient estimator: For a given $x^{t}\in\mathcal{X}$ in the $t$-th iteration, we can access an unbiased gradient estimator $\widehat{g}^{t}$, i.e., $\mathbb{\mathbb{E}}\left[\widehat{g}^{t}\mid\mathcal{F}^{t-1}\right]\in\partial f(x^{t})$, where $\mathcal{F}^{t}\coloneqq\sigma\left(\widehat{g}^{s},s\in\left[t\right]\right)$ is the natural filtration.
  5. 5A.
    Finite variance: $\exists\sigma\geq 0$ denoting the noise level such that $\mathbb{\mathbb{E}}\left[\left\|\xi^{t}\right\|_{*}^{2}\mid\mathcal{F}^{t-1}\right]\leq\sigma^{2}$ where $\xi^{t}\coloneqq\widehat{g}^{t}-\mathbb{\mathbb{E}}\left[\widehat{g}^{t}\mid\mathcal{F}^{t-1}\right]$.
  6. 5B.
    Sub-Gaussian noise: $\exists\sigma\geq 0$ denoting the noise level such that $\mathbb{\mathbb{E}}\left[\exp\left(\lambda\left\|\xi^{t}\right\|_{*}^{2}\right)\mid\mathcal{F}^{t-1}\right]\leq\exp\left(\lambda\sigma^{2}\right),\forall\lambda\in\left[0,\sigma^{-2}\right]$.
We briefly discuss the assumptions here. Assumptions 1., 4., and 5A. are standard in the stochastic optimization literature. Assumption 2. is known as relative strong convexity appeared in previous works (Hazan and Kale, 2014; Lu et al., 2018). We use it here since the last-iterate convergence rate will be derived for the CSMD algorithm, which employs Bregman divergence to exploit the non-Euclidean geometry. In particular, when $\left\|\cdot\right\|$ is the standard $\ell_{2}$ norm, we can take $\psi(x)=\frac{1}{2}\left\|x\right\|^{2}$ to recover the common definition of strong convexity. Assumption 3. is borrowed from Section 4.2 in Lan (2020). Note that both $L$-smooth functions (by taking $M=0$) and $G$-Lipschitz functions (by taking $L=0$ and $M=2G$) are subclasses of Assumption 3.. Lastly, Assumption 5B. is used for the high-probability convergence bound.
Our proofs for the high-probability convergence rely on the following simple fact for the centered sub-Gaussian random vector. Similar results have been proved in prior works (Vershynin, 2018; Liu et al., 2023b). For completeness, we include the proof in Appendix A.
Lemma 2.1.
Given a sigma algebra $\mathcal{F}$ and a random vector $Z\in\mathbb{R}^{d}$ that is $\mathcal{F}$-measurable, if $\xi\in\mathbb{R}^{d}$ is a random vector satisfying $\mathbb{\mathbb{E}}\left[\xi\mid\mathcal{F}\right]=0$ and $\mathbb{\mathbb{E}}\left[\exp\left(\lambda\left\|\xi\right\|_{*}^{2}\right)\mid\mathcal{F}\right]\leq\exp\left(\lambda\sigma^{2}\right),\forall\lambda\in\left[0,\sigma^{-2}\right]$, then
$\mathbb{\mathbb{E}}\left[\exp\left(\left\langle\xi,Z\right\rangle\right)\mid\mathcal{F}\right]\leq\exp\left(\sigma^{2}\left\|Z\right\|^{2}\right).$
2.1 Convergence Criterion
We always measure the convergence via the function value gap, i.e., $F(x)-F(x^{*})$. There are several reasons to stick to this criterion. First, for the general convex case, the function value gap is the standard metric. Next, for strongly convex functions, the function value gap is always a stronger measurement than the squared distance to the optimal solution since $\left\|x-x^{*}\right\|^{2}=O\left(F(x)-F(x^{*})\right)$ holds by strong convexity. Even if $F(x)$ is additionally assumed to be ($L,0$)-smooth (e.g., $f(x)$ is ($L,0$)-smooth and $h(x)=0$), the bound on $\left\|x-x^{*}\right\|^{2}$ cannot be converted to the bound on $F(x)-F(x^{*})$ since $F(x)-F(x^{*})\leq\left\langle\nabla F(x^{*}),x-x^{*}\right\rangle+\frac{L}{2}\left\|x-x^{*}\right\|^{2}=O(\left\|\nabla F(x^{*})\right\|_{*}\left\|x-x^{*}\right\|+\left\|x-x^{*}\right\|^{2})$, which is probably worse than $O(\left\|x-x^{*}\right\|^{2})$ as $x^{*}$ is only a local minimizer meaning $\left\|\nabla F(x^{*})\right\|_{*}$ possibly to be non-zero. Moreover, the function value gap is important in both the theoretical and practical sides of modern machine learning (e.g., the generalization error).
3 Last-Iterate Convergence of Stochastic Gradient Methods
Algorithm 1 Composite Stochastic Mirror Descent (CSMD)
Input: $x^{1}\in\mathcal{X}$, $\eta_{t\in\left[T\right]}>0$.
for $t=1$ to $T$ do
$x^{t+1}=\mathrm{argmin}_{x\in\mathcal{X}}h(x)+\left\langle\widehat{g}^{t},x-x^{t}\right\rangle+\frac{D_{\psi}(x,x^{t})}{\eta_{t}}$
Return $x^{T+1}$
The algorithm, Composite Stochastic Mirror Descent, is presented in Algorithm 1. When $h(x)=0$, Algorithm 1 degenerates to the standard Stochastic Mirror Descent algorithm. If we further consider the case $\left\|\cdot\right\|=\left\|\cdot\right\|_{2}$, Algorithm 1 can recover the standard projected SGD by taking $\psi(x)=\frac{1}{2}\left\|x\right\|_{2}^{2}$. We assume $T\geq 2$ throughout the following paper to avoid some algebraic issues in the proof. The full version of every following theorem with its proof is deferred into the appendix.
3.1 General Convex Functions
In this section, we focus on the last-iterate convergence of Algorithm 1 for general convex functions (i.e., $\mu_{f}=\mu_{h}=0$). First, the in-expectation convergence rates are shown in Theorem 3.1.
Theorem 3.1.
Under Assumptions 1.-4. and 5A. with $\mu_{f}=\mu_{h}=0$:
If $T$ is unknown, by taking $\eta_{t\in\left[T\right]}=\frac{1}{2L}\land\frac{\eta}{\sqrt{t}}$ with $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x^{*},x^{1})}{M^{2}+\sigma^{2}}}\right)$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x^{*})\right]\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{T}+\frac{(M+\sigma)\sqrt{D_{\psi}(x^{*},x^{1})}\log T}{\sqrt{T}}\right).$
If $T$ is known, by taking $\eta_{t\in\left[T\right]}=\frac{1}{2L}\land\frac{\eta}{\sqrt{T}}$ with $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x^{*},x^{1})}{(M^{2}+\sigma^{2})\log T}}\right)$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x^{*})\right]\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{T}+\frac{(M+\sigma)\sqrt{D_{\psi}(x^{*},x^{1})\log T}}{\sqrt{T}}\right).$
Before moving on to the high-probability bounds, we would like to talk more about these in-expectation convergence results. First, the constant $\eta$ here is optimized to obtain the best dependence on the parameters $M,\sigma$ and $D_{\psi}(x^{*},x^{1})$. Indeed, the last iterate provably converges for arbitrary $\eta>0$, but with a worse dependence on $M,\sigma$ and $D_{\psi}(x^{*},x^{1})$. We refer the reader to Theorem C.1 in the appendix for a full version of Theorem 3.1 with any $\eta>0$.
Next, by taking $L=0$, we immediately get the (nearly) optimal $\widetilde{O}(1/\sqrt{T})$ convergence rate of the last iterate for non-smooth functions. Note that our bounds are better than Shamir and Zhang (2013) since it only works for bounded domains and non-composite optimization. Besides, when considering smooth problems (taking $M=0$), to our best knowledge, our $\widetilde{O}(L/T+\sigma/\sqrt{T})$ bound is the first improvement since the $O(1/\sqrt[3]{T})$ rate by Moulines and Bach (2011). Moreover, compared to Moulines and Bach (2011), Theorem 3.1 does not rely on some restrictive assumptions like bounded stochastic gradients or $x^{*}$ being a global optimal point but is able to be used for the more general composite problems. Additionally, it is worth remarking that the $\widetilde{O}(L/T+\sigma/\sqrt{T})$ rate matches the optimal $O(L/T+\sigma/\sqrt{T})$ rate for the averaged output $x_{avg}^{T+1}=(\sum_{t=2}^{T+1}x^{t})/T$ (Lan, 2020) up to an extra logarithmic factor. Notably, our bounds are also adaptive to the noise $\sigma$ in this case. In other words, we can recover the well-known $O(L/T)$ rate for the last iterate of the GD algorithm in the noiseless case. Last but most importantly, our proof is unified and thus can be applied to different settings (e.g., general domains, ($L,M$)-smoothness, non-Euclidean norms, etc.) simultaneously.
Remark 3.2.
Orabona (2020) exhibited a circuitous method based on comparing the last iterate with the averaged output to show the in-expectation last-iterate convergence for non-composite non-smooth convex optimization in general domains. However, it did not explicitly generalize to the broader problems considered in this paper. Moreover, our method is done in a direct manner (see Section 4).
Theorem 3.3.
Under Assumptions 1.-4. and 5B. with $\mu_{f}=\mu_{h}=0$ and let $\delta\in(0,1)$:
If $T$ is unknown, by taking $\eta_{t\in\left[T\right]}=\frac{1}{2L}\land\frac{\eta}{\sqrt{t}}$ with $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x^{*},x^{1})}{M^{2}+\sigma^{2}\log\frac{1}{\delta}}}\right)$, then with probability at least $1-\delta$, there is
$F(x^{T+1})-F(x^{*})\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{T}+\frac{(M+\sigma\sqrt{\log\frac{1}{\delta}})\sqrt{D_{\psi}(x^{*},x^{1})}\log T}{\sqrt{T}}\right).$
If $T$ is known, by taking $\eta_{t\in\left[T\right]}=\frac{1}{2L}\land\frac{\eta}{\sqrt{T}}$ with $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x^{*},x^{1})}{(M^{2}+\sigma^{2}\log\frac{1}{\delta})\log T}}\right)$, then with probability at least $1-\delta$, there is
$F(x^{T+1})-F(x^{*})\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{T}+\frac{(M+\sigma\sqrt{\log\frac{1}{\delta}})\sqrt{D_{\psi}(x^{*},x^{1})\log T}}{\sqrt{T}}\right).$
In Theorem 3.3, we present the high-probability bounds for $(L,M)$-smooth functions. Again, the constant $\eta$ is picked to get the best dependence on the parameters $M,\sigma,D_{\psi}(x^{*},x^{1})$ and $\log(1/\delta)$. The full version of Theorem 3.3 with arbitrary $\eta$, Theorem C.2, is deferred into the appendix. Compared with Theorem 3.1, the high-probability rates only incur an extra $O(\sqrt{\log(1/\delta)})$ factor (or $O(\log(1/\delta))$ for arbitrary $\eta$, which is known to be optimal for $L=0$ (Harvey et al., 2019b)).
In contrast to the previous bounds (Harvey et al., 2019b; Jain et al., 2021) that only work for Lipschitz functions in a compact domain, our results are the first to describe the high-probability behavior of Algorithm 1 for the wider $(L,M)$-smooth function class in a general domain even with sub-Gaussian noise, not to mention composite objectives and non-Euclidean norms. Even in the special smooth case (setting $M=0$), as far as we know, this is also the first last-iterate high-probability bound being adaptive to the noise $\sigma$ at the same time for plain stochastic gradient methods. Unlike the previous proofs employing some new probability tools (e.g., the generalized Freedman’s inequality in Harvey et al. (2019b)), our high-probability argument is simple and only based on the basic property of sub-Gaussian random vectors (see Lemma 2.1). Therefore, we believe our work can bring some new insights to researchers to gain a better understanding of the last-iterate convergence of stochastic gradient methods.
3.1.1 Optimal Rates via the Linear Decay Step Size
In this section, we show that the $O(\sqrt{\log T})$ factor appeared in Theorems 3.1 and 3.3 for the case of known $T$ can be totally removed. Even restricted to the special case of Lipschitz objectives (i.e., $L=0$ in Assumption 3.), this is a strict improvement compared to Jain et al. (2021), in which the $O(1/\sqrt{T})$ rate is obtained under the bounded noise assumption on a compact domain $\mathcal{X}$.
Theorem 3.4.
Under Assumptions 1.-4. and 5A. with $\mu_{f}=\mu_{h}=0$, if $T$ is known, by taking $\eta_{t\in\left[T\right]}=\frac{T-t+1}{2LT}\land\frac{\eta(T-t+1)}{T^{\frac{3}{2}}}$ with $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x^{*},x^{1})}{M^{2}+\sigma^{2}}}\right)$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x^{*})\right]\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{T}+\frac{(M+\sigma)\sqrt{D_{\psi}(x^{*},x^{1})}}{\sqrt{T}}\right).$
Theorem 3.5.
Under Assumptions 1.-4. and 5B. with $\mu_{f}=\mu_{h}=0$ and let $\delta\in(0,1)$, if $T$ is known, by taking $\eta_{t\in\left[T\right]}=\frac{T-t+1}{2LT}\land\frac{\eta(T-t+1)}{T^{\frac{3}{2}}}$ with $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x^{*},x^{1})}{M^{2}+\sigma^{2}\log\frac{1}{\delta}}}\right)$, then with probability at least $1-\delta$, there is
$F(x^{T+1})-F(x^{*})\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{T}+\frac{(M+\sigma\sqrt{\log\frac{1}{\delta}})\sqrt{D_{\psi}(x^{*},x^{1})}}{\sqrt{T}}\right).$
The step size decaying linearly was introduced by Zamani and Glineur (2025), in which the authors proved that it ensures the last iterate of the subgradient method converges at the rate $O(1/\sqrt{T})$ for non-composite non-smooth convex objectives with the Euclidean norm. Here, we show it can also guarantee the optimal $O(L/T+(M+\sigma)/\sqrt{T})$ rate in our setting. Again, the constant $\eta$ can be set to any positive number. The full statements of Theorems 3.4 and 3.5 are provided in Appendix C.1.
3.2 Strongly Convex Functions
Now we turn our attention to strongly convex functions. Due to the space limitation, we only provide the results for the case of $\mu_{f}>0$ and $\mu_{h}=0$. The other case, $\mu_{f}=0$ and $\mu_{h}>0$, will be delivered in Appendix D.2.
Theorem 3.6.
Under Assumptions 1.-4. and 5A. with $\mu_{f}>0$ and $\mu_{h}=0$, let $\kappa_{f}\coloneqq\frac{L}{\mu_{f}}\geq 0$:
If $T$ is unknown, by taking either $\eta_{t\in\left[T\right]}=\frac{1}{\mu_{f}(t+2\kappa_{f})}$ or $\eta_{t\in\left[T\right]}=\frac{2}{\mu_{f}(t+1+4\kappa_{f})}$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x^{*})\right]\leq\begin{cases}O\left(\frac{LD_{\psi}(x^{*},x^{1})}{T}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{f}(T+\kappa_{f})}\right)&\eta_{t\in\left[T\right]}=\frac{1}{\mu_{f}(t+2\kappa_{f})}\\ O\left(\frac{L(1+\kappa_{f})D_{\psi}(x^{*},x^{1})}{T(T+\kappa_{f})}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{f}(T+\kappa_{f})}\right)&\eta_{t\in\left[T\right]}=\frac{2}{\mu_{f}(t+1+4\kappa_{f})}\end{cases}.$
If $T$ is known, by taking $\eta_{t\in\left[T\right]}=\begin{cases}\frac{1}{\mu_{f}(1+2\kappa_{f})}&t=1\\ \frac{1}{\mu_{f}(\eta+2\kappa_{f})}&2\leq t\leq\tau\\ \frac{2}{\mu_{f}(t-\tau+2+4\kappa_{f})}&t\geq\tau+1\end{cases}$ with $\eta\coloneqq 1.5$ and $\tau\coloneqq\left\lceil\frac{T}{2}\right\rceil$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x^{*})\right]\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{\exp\left(\frac{T}{3+4\kappa_{f}}\right)}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{f}(T+\kappa_{f})}\right).$
The in-expectation rates are stated in Theorem 3.6, where the constant $\eta=1.5$ is chosen without any special reason. Generally speaking, it can be any non-negative number that satisfies $\eta+\kappa_{f}>1$. The interested reader could refer to Theorem D.1 in the appendix for a completed version of Theorem 3.6. We would like to remind the reader that $\kappa_{f}\geq 1$ is not necessary, as we are considering the general $(L,M)$-smooth functions. Hence, it can be zero.
As before, we first take $L=0$ to consider the special Lipschitz case. Due to $\kappa_{f}=0$ now, all bounds will degenerate to $O(\log T/T)$, which is known to be optimal for the step size $\nicefrac{{1}}{{\mu_{f}t}}$ (Harvey et al., 2019b) and only incurs an extra $O(\log T)$ factor compared with the best $O(1/T)$ bound when $T$ is known (Jain et al., 2021). We would also like to mention that Theorem 3.6 is the first to give the in-expectation last-iterate bound for the step size $\nicefrac{{2}}{{\mu_{f}(t+1)}}$. Interestingly, the extra $O(\log T)$ factor appears again compared to the known $O(1/T)$ bound on the function value gap for the non-uniform averaging strategy under this step size (Lacoste-Julien et al., 2012). Besides, Lacoste-Julien et al. (2012) also shows $\mathbb{\mathbb{E}}\left[\left\|x^{T+1}-x^{*}\right\|_{2}^{2}\right]=O(1/T)$. Whereas it is currently unknown whether our $\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x^{*})\right]=O(\log T/T)$ bound can be improved to match the $O(1/T)$ rate.
For the general $(L,M)$-smooth case (even for $(L,0)$-smoothness), our bounds are the first convergence results for the last iterate of stochastic gradient methods with respect to the function value gap [^6] Note that the rates under the PL-condition (e.g., Gower et al. (2021); Khaled and Richtárik (2023)) are incompatible with our settings since they can be only applied to non-constrained, non-composite and $(L,0)$-smooth optimization problems with the Euclidean norm.. Remarkably, all of these rates do not require prior knowledge of $M$ or $\sigma$ to set the step size. In particular, the bound for known $T$ is adaptive to $\sigma$ when $M=0$, i.e., it can recover the well-known linear convergence rate $O(\exp(-T/\kappa_{f}))$ when $\sigma=0$.
Theorem 3.7.
Under Assumptions 1.-4. and 5B. with $\mu_{f}>0$ and $\mu_{h}=0$, let $\kappa_{f}\coloneqq\frac{L}{\mu_{f}}\geq 0$ and $\delta\in(0,1)$:
If $T$ is unknown, by taking either $\eta_{t\in\left[T\right]}=\frac{1}{\mu_{f}(t+2\kappa_{f})}$ or $\eta_{t\in\left[T\right]}=\frac{2}{\mu_{f}(t+1+4\kappa_{f})}$, then with probability at least $1-\delta$, there is
$F(x^{T+1})-F(x^{*})\leq\begin{cases}O\left(\frac{\mu_{f}(1+\kappa_{f})D_{\psi}(x^{*},x^{1})}{T}+\frac{(M^{2}+\sigma^{2}\log\frac{1}{\delta})\log T}{\mu_{f}(T+\kappa_{f})}\right)&\eta_{t\in\left[T\right]}=\frac{1}{\mu_{f}(t+2\kappa_{f})}\\ O\left(\frac{\mu_{f}(1+\kappa_{f})^{2}D_{\psi}(x^{*},x^{1})}{T(T+\kappa_{f})}+\frac{(M^{2}+\sigma^{2}\log\frac{1}{\delta})\log T}{\mu_{f}(T+\kappa_{f})}\right)&\eta_{t\in\left[T\right]}=\frac{2}{\mu_{f}(t+1+4\kappa_{f})}\end{cases}.$
If $T$ is known, by taking $\eta_{t\in\left[T\right]}=\begin{cases}\frac{1}{\mu_{f}(1+2\kappa_{f})}&t=1\\ \frac{1}{\mu_{f}(\eta+2\kappa_{f})}&2\leq t\leq\tau\\ \frac{2}{\mu_{f}(t-\tau+2+4\kappa_{f})}&t\geq\tau+1\end{cases}$ with $\eta\coloneqq 1.5$ and $\tau\coloneqq\left\lceil\frac{T}{2}\right\rceil$, then with probability at least $1-\delta$, there is
$F(x^{T+1})-F(x^{*})\leq O\left(\frac{\mu_{f}(1+\kappa_{f})D_{\psi}(x^{*},x^{1})}{\exp\left(\frac{T}{3+4\kappa_{f}}\right)}+\frac{(M^{2}+\sigma^{2}\log\frac{1}{\delta})\log T}{\mu_{f}(T+\kappa_{f})}\right).$
To finish this section, we provide the high-probability convergence results in Theorem 3.7. Again, the constant $\eta=1.5$ is set without any particular reason. The full statement with general $\eta$, Theorem D.2, can be found in the appendix. Besides, $\kappa_{f}$ is possible to be zero as mentioned above. Compared with Theorem 3.6, only an additional $O(\log(1/\delta))$ factor appears. Such an extra loss is known to be inevitable for $L=0$ due to Harvey et al. (2019b).
For the Lipschitz case (i.e., $L=\kappa_{f}=0$), by noticing $D_{\psi}(x^{*},x^{1})=O(M^{2}/\mu_{f}^{2})$ [^7] This holds now due to $\mu_{f}\left\|x^{*}-x^{1}\right\|^{2}/2\leq\mu_{f}D_{\psi}(x^{*},x^{1})\leq M\left\|x^{*}-x^{1}\right\|$., all of these bounds will degenerate to $O(\log(1/\delta)\log T/T)$ matching the best-known last-iterate bound proved by Harvey et al. (2019b) for the step size $\nicefrac{{1}}{{\mu_{f}t}}$. For the step size $\nicefrac{{2}}{{\mu_{f}(t+1)}}$, Harvey et al. (2019a) has proved the high-probability bound $O(\log(1/\delta)/T)$ for the non-uniform averaging output instead of the last iterate. Hence, as far as we know, our high-probability rate for the step size $\nicefrac{{2}}{{\mu_{f}(t+1)}}$ is new. However, we would like to mention that our bound for known $T$ is worse by a logarithmic factor than that of Jain et al. (2021), though, which assumes bounded noise.
Finally, let us go back to the general $(L,M)$-smooth case. To our best knowledge, our results are the first to prove that the last iterate of plain stochastic gradient methods enjoys the provable high-probability convergence (even in the smooth case, i.e., $M=0$). Hence, we believe our work closes the gap between the lack of theoretical understanding and good performance of the last iterate of SGD for smooth and strongly convex functions. Lastly, the same as the in-expectation bound for known $T$ in Theorem 3.6, our high-probability bound is also adaptive to $\sigma$ when $M=0$.
4 Unified Theoretical Analysis
In this section, we introduce the ideas in our analysis and present three important lemmas, all the missing proofs of which are deferred into Appendix B.
The key insight in our proofs is to utilize the convexity of $F(x)$, which is highly inspired by the recent work of Zamani and Glineur (2025). To be more precise, using the classic convergence analysis for non-composite Lipschitz convex problems as an example, people always consider to upper bound the function value gap $f(x^{t})-f(x^{*})$ (probably with some weight before it) then sum them over time to obtain the ergodic rate. Whereas, in such an argument, convexity is not necessary in fact (except if one wants to bound the average iterate in the last step). Hence, if the convexity of $f$ can be utilized somewhere, it is reasonable to expect a last-iterate convergence guarantee. Actually, this thought is possible as shown by Zamani and Glineur (2025), in which the authors upper bound the quantity $f(x^{t})-f(z^{t})$ where $z^{t}$ is a carefully chosen convex combination of other points and finally obtain the last-iterate rate by lower bounding $f(z^{t})$ via convexity.
Though Zamani and Glineur (2025) only shows how to prove the last-iterate convergence for deterministic non-composite Lipschitz convex optimization under the Euclidean norm, we can catch the most important message conveyed by their paper and apply it to our settings. Formally speaking, we will upper bound the term $F(x^{t+1})-F(z^{t})$ for a well-designed $z^{t}$ rather than directly bound the function value gap $F(x^{t+1})-F(x^{*})$. This idea can finally help us construct a unified proof and obtain several novel results without prior restrictive assumptions. Through careful calculations, the new analysis leads us to the following most important and unified result, Lemma 4.1.
Lemma 4.1.
Under Assumptions 2. and 3., suppose $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}$ and let $\gamma_{t\in\left[T\right]}\coloneqq\eta_{t}\prod_{s=2}^{t}\frac{1+\mu_{h}\eta_{s-1}}{1-\mu_{f}\eta_{s}}$, if $w_{t\in\left[T\right]}\geq 0$ is a non-increasing sequence and $v_{t\in\left\{0\right\}\cup\left[T\right]}>0$ is defined as $v_{t}\coloneqq\frac{w_{T}\gamma_{T}}{\sum_{s=t\lor 1}^{T}w_{s}\gamma_{s}}$, then for any $x\in\mathcal{X}$, we have
$w_{T}\gamma_{T}v_{T}\left(F(x^{T+1})-F(x)\right)$
$\leq$ $w_{1}(1-\mu_{f}\eta_{1})v_{0}D_{\psi}(x,x^{1})+\sum_{t=1}^{T}2w_{t}\gamma_{t}\eta_{t}v_{t}(M^{2}+\left\|\xi^{t}\right\|_{*}^{2})$
$+\sum_{t=1}^{T}w_{t}\gamma_{t}v_{t-1}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle+\sum_{t=2}^{T}(w_{t}-w_{t-1})\gamma_{t}(\eta_{t}^{-1}-\mu_{f})v_{t-1}D_{\psi}(z^{t-1},x^{t}),$
where $\xi^{t}\coloneqq\widehat{g}^{t}-\mathbb{\mathbb{E}}\left[\widehat{g}^{t}\mid\mathcal{F}^{t-1}\right],\forall t\in\left[T\right]$ and $z^{t}\coloneqq\frac{v_{0}}{v_{t}}x+\sum_{s=1}^{t}\frac{v_{s}-v_{s-1}}{v_{t}}x^{s},\forall t\in\left\{0\right\}\cup\left[T\right]$.
Let us discuss Lemma 4.1 more here. The requirement of the step size $\eta_{t}$ having an upper bound $\nicefrac{{1}}{{2L\lor\mu_{f}}}$ is common in the optimization literature. $\gamma_{t}$ is used to ensure we can telescope sum some terms. For the special case $\mu_{f}=\mu_{h}=0$, it degenerates to $\eta_{t}$. $\xi^{t}$ naturally shows up as we are considering stochastic optimization. The most important sequences are $w_{t},v_{t}$ and $z^{t}$. As mentioned above, the appearance of $z^{t}$ is to make sure to get the last-iterate convergence. For how to find such a sequence, we refer the reader to our proofs in Appendix B for details.
We would like to say more about the sequence $w_{t}$ before moving on. Suppose we are in the deterministic case temporarily, i.e., $\xi^{t}=0$, then a natural choice is to set $w_{t}=1,\forall t\in\left[T\right]$ to remove the last residual summation. It turns out this is the correct choice even for the following in-expectation bound in Lemma 4.2. So why do we still need this redundant $w_{t}$? The reason is that setting $w_{t}$ to be one is not enough for the high-probability analysis. More precisely, if we still choose $w_{t}=1,\forall t\in\left[T\right]$, then there will be some extra positive terms after the concentration argument in the R.H.S. of the inequality in Lemma 4.1. To deal with this issue, we borrow the idea recently developed by Liu et al. (2023b), in which the authors employ an extra sequence $w_{t}$ to give a clear proof for the high-probability bound for stochastic gradient methods. We refer the reader to Liu et al. (2023b) for a detailed explanation of this technique.
Lemma 4.2.
Under Assumptions 2.-4. and 5A., suppose $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}$ and let $\gamma_{t\in\left[T\right]}\coloneqq\eta_{t}\prod_{s=2}^{t}\frac{1+\mu_{h}\eta_{s-1}}{1-\mu_{f}\eta_{s}}$, then for any $x\in\mathcal{X}$, we have
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq\frac{(1-\mu_{f}\eta_{1})D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}.$
Suppose Lemma 4.1 holds, Lemma 4.2 is immediately obtained by setting $w_{t}=1,\forall t\in\left[T\right]$ and using Assumptions 4. and 5A.. This unified result for the in-expectation last-iterate convergence can be applied to many different settings, such as composite optimization and non-Euclidean norms, without any restrictive assumptions.
Lemma 4.3.
Under Assumptions 2.-4. and 5B., suppose $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}$ and let $\gamma_{t\in\left[T\right]}\coloneqq\eta_{t}\prod_{s=2}^{t}\frac{1+\mu_{h}\eta_{s-1}}{1-\mu_{f}\eta_{s}}$, then for any $x\in\mathcal{X}$ and $\delta\in(0,1)$, with probability at least $1-\delta$, we have
$F(x^{T+1})-F(x)\leq 2\left(1+\max_{2\leq t\leq T}\frac{1}{1-\mu_{f}\eta_{t}}\right)\left[\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+\left(M^{2}+\sigma^{2}\left(1+2\log\frac{2}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}\right].$
To get Lemma 4.3, we need some extra effort to find the correct $w_{t}$ and invoke a simple property of sub-Gaussian random vectors (Lemma 2.1). The details can be found in Appendix B. Compared with prior works, this unified high-probability bound can be applied to various scenarios including general domains and sub-Gaussian noise.
Equipped with Lemma 4.2 and Lemma 4.3, we can prove all the theorems provided in Section 3 by plugging in different step sizes for different cases.
5 Last-Iterate Convergence under Heavy-Tailed Noise
In Section 3, we have shown how to obtain the convergence rate of the last iterate of CSMD in a unified manner. However, the in-expectation bounds require the finite variance assumption, which may be too optimistic in modern machine learning. As pointed out by recent studies, e.g., Mirek (2011); Simsekli et al. (2019); Zhang et al. (2020); Zhou et al. (2020); Gurbuzbalaban and Hu (2021); Camuto et al. (2021); Hodgkinson and Mahoney (2021), noises in different tasks are observed to have heavy tails, meaning that they only have a finite $p$-th moment where $p\in\left(1,2\right)$.
Naturally, one would ask whether the last iterate of stochastic gradient methods can still converge under this new challenging assumption. Surprisingly, the answer is still yes. In this section, we extend the idea used before to obtain the in-expectation last-iterate convergence rate of Algorithm 1 under heavy-tailed noise (see Assumption 5C. below). As far as we know, this is a new result in the related area.
5.1 Additional Related Work
To the best of our knowledge, for stochastic gradient methods without momentum or averaging operations, there exists no work that proves a non-asymptotic last-iterate convergence rate for the function value gap. Hence, we only review prior works related to the average iterate.
Average-iterate convergence under heavy-tailed noise: Zhang et al. (2020) proves that SGD with clipping guarantees an $O(T^{\frac{2}{p}-2})$ rate in expectation for Lipschitz strongly convex objectives. Without clipping and strong convexity, Vural et al. (2022) establishes the $O(T^{\frac{1}{p}-1})$ in-expectation convergence for the stochastic mirror descent algorithm. For the smooth problem, Sadiev et al. (2023) shows that SGD with clipping also admits nearly optimal high-probability convergence rates $\widetilde{O}(T^{\frac{1}{p}-1})$ and $\widetilde{O}(T^{\frac{2}{p}-2})$ for convex and strongly convex functions, respectively. For the non-smooth problem, Liu and Zhou (2023) is the first to prove that the clipping stochastic gradient method ensures the optimal rates $O(T^{\frac{1}{p}-1})$ under convexity and $O(T^{\frac{2}{p}-2})$ under strong convexity (for both in-expectation and high-probability convergence). Nguyen et al. (2023) later obtains the optimal high-probability rate $O(T^{\frac{1}{p}-1})$ for the smooth problem.
Lower bounds under heavy-tailed noise: For Lipschitz convex functions, Nemirovski and Yudin (1983); Raginsky and Rakhlin (2009); Vural et al. (2022) have proved a lower bound of $\Omega(T^{\frac{1}{p}-1})$. For smooth and strongly convex functions (in the sense of the Euclidean norm), a lower bound of $\Omega(T^{\frac{2}{p}-2})$ is provided in Zhang et al. (2020).
5.2 New Assumption and Some Discussions
We formally introduce Assumption 5C. to describe the noise that has a heavy tail. Note that Assumption 5C. will be the same as Assumption 5A. if we enforce $p=2$.
  1. 5C.
    Finite $p$-th moment: $\exists p\in\left(1,2\right)$, $\sigma\geq 0$ denoting the noise level such that $\mathbb{\mathbb{E}}\left[\left\|\xi^{t}\right\|_{*}^{p}\mid\mathcal{F}^{t-1}\right]\leq\sigma^{p}$.
In order to fit the new assumption, we also need to make a slight change to the mirror map $\psi$. Instead of being $1$-strongly convex, we now require $\psi$ to be $(1,\frac{p}{p-1})$ uniformly convex, i.e.,
$D_{\psi}(x,y)=\psi(x)-\psi(y)-\left\langle\nabla\psi(y),x-y\right\rangle\geq\frac{p-1}{p}\left\|x-y\right\|^{\frac{p}{p-1}},\forall x,y\in\mathcal{X}.$ (1)
Note that when $p=2$, this condition reduces to the standard $1$-strong convexity. The idea of using a uniformly convex mirror map was proposed in Vural et al. (2022) before.
Moreover, we will only consider general convex functions (i.e., $\mu_{f}=\mu_{h}=0$ in Assumption 2.) in this section. The main reason is that if considering the case $\mu_{f}>0$ in Assumption 2., we can immediately observe the following fact,
$\frac{\mu_{f}(p-1)}{p}\left\|x-y\right\|^{\frac{p}{p-1}}\leq\mu_{f}D_{\psi}(x,y)\leq f(x)-f(y)-\left\langle g,x-y\right\rangle,\forall x,y\in\mathcal{X},g\in\partial f(y),$
which implies $f$ itself should be $(\mu_{f},\frac{p}{p-1})$ uniformly convex. However, this condition is less common and cannot even be satisfied in the case of a strongly convex $f$ when considering the Euclidean norm. The other point is that even if $f$ indeed satisfies Assumption 2. with $\mu_{f}>0$, then by Assumption 3., there is
$\frac{\mu_{f}(p-1)}{p}\left\|x-y\right\|^{\frac{p}{p-1}}\leq\frac{L}{2}\left\|x-y\right\|^{2}+M\left\|x-y\right\|,\forall x,y\in\mathcal{X},$
which implies $\sup_{x,y\in\mathcal{X}}\left\|x-y\right\|<\infty$ due to $\frac{p}{p-1}>2$. Note that, with a compact domain, it is possible to obtain convergence bounds but in a non-unified way. However, this is contrary to the purpose of this paper. Due to these two points, we do not consider the case of $\mu_{f}>0$. We remark that $\mu_{h}>0$ will not lead to a result of a bounded domain. But it may also suffer from the first issue (e.g., when $h(x)=\frac{1}{2}\left\|x\right\|_{2}^{2}$). Therefore, the case of $\mu_{h}>0$ is also discarded.
Remark 5.1.
We clarify that if $\mu_{f}>0$ or $\mu_{h}>0$ indeed holds (for example, taking $h(x)=\psi(x)$ implies $\mu_{h}=1$), one can still apply our analysis framework to obtain the last-iterate convergence rate of the CSMD algorithm. This potential work is left to the interested reader.
Lastly, as mentioned before, we will only provide in-expectation convergence rates for Algorithm 1. The key reason is that we do not consider the clipping technique. As far as we know, all the high-probability results for the averaged output are established with clipping (even under the finite variance assumption). How to obtain the $O(\mathrm{polylog}\frac{1}{\delta})$ dependence for an algorithm without clipping in a single run remains unknown. Moreover, there is some evidence showing that the dependence of $\Omega(\mathrm{poly}\frac{1}{\delta})$ is inevitable in certain cases (e.g., Theorem 2.1 in Sadiev et al. (2023)). Hence, we leave this challenging problem (finding the high-probability last-iterate bound without clipping) as a future direction.
5.3 General Convex Functions under Heavy-Tailed Noise
Now we are ready to provide the convergence results under heavy-tailed noise.
Theorem 5.2.
Under Assumptions 1.-4. and 5C. with $\mu_{f}=\mu_{h}=0$:
If $T$ is unknown, by taking $\eta_{t\in\left[T\right]}=\frac{\eta_{*}}{t^{\frac{2-p}{p}}}\land\frac{\eta}{t^{\frac{1}{p}}}$ with $\eta_{*}=\Theta\left(\frac{\left[D_{\psi}(x^{*},x^{1})\right]^{\frac{2-p}{p}}}{L}\right)$ and $\eta=\Theta\left(\left[\frac{D_{\psi}(x^{*},x^{1})}{M^{p}+\sigma^{p}}\right]^{\frac{1}{p}}\right)$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x^{*})\right]\leq O\left(\frac{L\left[D_{\psi}(x^{*},x^{1})\right]^{2-\frac{2}{p}}\log T}{T^{2-\frac{2}{p}}}+\frac{(M+\sigma)\left[D_{\psi}(x^{*},x^{1})\right]^{1-\frac{1}{p}}\log T}{T^{1-\frac{1}{p}}}\right).$
If $T$ is known, by taking $\eta_{t\in\left[T\right]}=\frac{\eta_{*}}{T^{\frac{2-p}{p}}}\land\frac{\eta}{T^{\frac{1}{p}}}$ with $\eta_{*}=\Theta\left(\frac{1}{L}\left[\frac{D_{\psi}(x^{*},x^{1})}{\log T}\right]^{\frac{2-p}{p}}\right)$ and $\eta=\Theta\left(\left[\frac{D_{\psi}(x^{*},x^{1})}{(M^{p}+\sigma^{p})\log T}\right]^{\frac{1}{p}}\right)$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x^{*})\right]\leq O\left(\frac{L\left[D_{\psi}(x^{*},x^{1})\right]^{2-\frac{2}{p}}(\log T)^{\frac{2}{p}-1}}{T^{2-\frac{2}{p}}}+\frac{(M+\sigma)\left[D_{\psi}(x^{*},x^{1})\right]^{1-\frac{1}{p}}(\log T)^{\frac{1}{p}}}{T^{1-\frac{1}{p}}}\right).$
Theorem 5.2 is the first theoretical result proving the last-iterate convergence of stochastic gradient methods under heavy-tailed noise. Note that the rate is near optimal, matching the lower bound $\Omega(1/T^{1-\frac{1}{p}})$ (Nemirovski and Yudin, 1983; Raginsky and Rakhlin, 2009; Vural et al., 2022) up to a logarithmic factor. In Theorem 5.3, we further show how to remove the unsatisfied extra logarithmic factor when $T$ is known. Moreover, it is worth pointing out that Theorem 5.2 recovers Theorem 3.1 perfectly if enforcing $p=2$. In Theorem F.1 in the appendix, we provide the convergence theory for arbitrary positive $\eta_{*}$ and $\eta$.
5.3.1 Optimal Rate under Heavy-Tailed Noise
Theorem 5.3.
Under Assumptions 1.-4. and 5C. with $\mu_{f}=\mu_{h}=0$, if $T$ is known, by taking $\eta_{t\in\left[T\right]}=\frac{\eta_{*}(T-t+1)}{T^{\frac{2}{p}}}\land\frac{\eta(T-t+1)}{T^{\frac{1}{p}+1}}$ with $\eta_{*}=\Theta\left(\frac{\left[D_{\psi}(x^{*},x^{1})\right]^{\frac{2-p}{p}}}{L}\right)$ and $\eta=\Theta\left(\left[\frac{D_{\psi}(x^{*},x^{1})}{M^{p}+\sigma^{p}}\right]^{\frac{1}{p}}\right)$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x^{*})\right]\leq O\left(\frac{L\left[D_{\psi}(x^{*},x^{1})\right]^{2-\frac{2}{p}}}{T^{2-\frac{2}{p}}}+\frac{(M+\sigma)\left[D_{\psi}(x^{*},x^{1})\right]^{1-\frac{1}{p}}}{T^{1-\frac{1}{p}}}\right).$
The optimal convergence rate for known $T$ is presented in Theorem 5.3. The new step size generalizes the learning rate used in Theorem 3.4 (but still exhibits a linear decay) to handle heavy-tailed noise. Hence, for stochastic convex optimization under heavy-tailed noise, we show that the convergence rate of the last iterate can match the lower bound perfectly. The full version with any positive $\eta_{*}$ and $\eta$ is provided in Theorem F.2 in the appendix.
6 Last-Iterate Convergence under Sub-Weibull Noise
In Section 3, we establish high-probability last-iterate convergence under the assumption of sub-Gaussian noise. It is natural to ask whether it is possible to relax it to, for example, sub-exponential noise. In this section, we provide an affirmative answer by considering a general family of noise, sub-Weibull noise with a parameter $p\in\left(0,2\right)$ (see Assumption 5D. below), which contains sub-exponential noise as a special case ($p=1$). The sub-Weibull distribution was originally proposed by Vladimirova et al. (2020); Kuchibhotla and Chakrabortty (2022). It has already been studied under different problems (e.g., bandit problems (Hao et al., 2019; Kim and Tewari, 2019)).
6.1 Additional Related Work
Convergence under sub-Weibull noise: As far as we know, when considering convex optimization, only few convergence results have been established under sub-Weibull noise. Hence, we include references that also consider non-convex optimization or online optimization, e.g., Madden et al. (2024); Wood et al. (2022); Li and Liu (2022); Ospina et al. (2022); Liu et al. (2023a).
6.2 New Assumption and Some Discussions
The following assumption describes what sub-Weibull noise is:
  1. 5D.
    Sub-Weibull noise: $\exists p\in\left(0,2\right)$, $\sigma\geq 0$ denoting the noise level such that $\mathbb{\mathbb{E}}\left[\exp\left(\lambda\left\|\xi^{t}\right\|_{*}^{p}\right)\mid\mathcal{F}^{t-1}\right]\leq\exp\left(\lambda\sigma^{p}\right),\forall\lambda\in\left[0,\sigma^{-p}\right]$.
If we enforce $p=2$ in Assumption 5D., it recovers Assumption 5B. (i.e., sub-Gaussian noise). Intuitively, one can think of the tail of sub-Weibull noise with $p<2$ decays strictly slower than the tail of sub-Gaussian noise. We emphasize that it is also possible to obtain the high-probability bound for $p>2$. However, the case of $p>2$ is not harder than the sub-Gaussian case ($p=2$). So, we specify $p\in(0,2)$ in this section.
Next, we only provide the convergence theorems for general convex functions in this section for simplicity. In the unified analysis (see Section G in the appendix), the strongly convex case is stated as well. Therefore, the interested reader can still apply our Lemma G.2 to obtain the high-probability bound for the strongly convex case, following similar steps used in proving Theorem 3.7.
Moreover, under the new Assumption 5D., certain barriers will arise when applying the technique of Liu et al. (2023b), i.e., the auxiliary weight sequence $w_{t\in\left[T\right]}$ used in Lemmas 4.1 and 4.3. To deal with this newly risen issue in the high-probability analysis, we will introduce a different technique, borrowed from Ivgi et al. (2023) and Liu and Zhou (2023), to replace the weight sequence $w_{t\in\left[T\right]}$ previously used. A detailed discussion is provided in Appendix G.
Lastly, we state a useful technical tool, Lemma 6.1, the proof of which essentially follows the same way as proving concentration bounds for sub-Gaussian and sub-exponential random variables in Vershynin (2018). For details, see Section A in the appendix.
Lemma 6.1.
Given a sigma algebra $\mathcal{F}$ and a random vector $Z\in\mathbb{R}^{d}$ that is $\mathcal{F}$-measurable, if $\xi\in\mathbb{R}^{d}$ is a random vector satisfying $\mathbb{\mathbb{E}}\left[\xi\mid\mathcal{F}\right]=0$ and $\mathbb{\mathbb{E}}\left[\exp\left(\lambda\left\|\xi\right\|_{*}^{p}\right)\mid\mathcal{F}\right]\leq\exp\left(\lambda\sigma^{p}\right),\forall\lambda\in\left[0,\sigma^{-p}\right]$ for $p\in\left[1,2\right)$, then
$\mathbb{\mathbb{E}}\left[\exp\left(\left\langle\xi,Z\right\rangle\right)\mid\mathcal{F}\right]\leq\begin{cases}\exp\left(6\sigma^{2}\left\|Z\right\|^{2}\right)&\text{if }\left\|Z\right\|\leq\frac{1}{2\sigma}\text{ almost surely}\\ \exp\left(6\sigma^{2}\left\|Z\right\|^{2}+2^{q-1}\sigma^{q}\left\|Z\right\|^{q}\right)&p\in\left(1,2\right),q\coloneqq\frac{p}{p-1}\end{cases}.$
6.3 General Convex Functions under Sub-Weibull Noise
We first introduce a function $C(\delta,p)=\begin{cases}O\left(1+\log^{\frac{2}{p}}\frac{1}{\delta}\right)&p\in\left[1,2\right)\\ O\left(1+\log^{1+\frac{2}{p}}\frac{1}{\delta}\right)&p\in\left(0,1\right)\end{cases}$, which is used to reflect the dependence on $\log\frac{1}{\delta}$ for different $p$, the detailed definition is provided in (69) in Appendix G. When $p\in\left(0,1\right)$, it is possible to replace $C(\delta,p)=O\left(1+\log^{1+\frac{2}{p}}\frac{1}{\delta}\right)$ by $C(\delta,p)=O\left(1+\log^{\frac{2}{p}}\frac{T}{\delta}\right)$ by applying the concentration bound from Madden et al. (2024). However, we keep the factor $\mathrm{polylog\frac{1}{\delta}}$ to get rid of $T$. Now we can state the convergence results.
Theorem 6.2.
Under Assumptions 1.-4. and 5D. with $\mu_{f}=\mu_{h}=0$ and let $\delta\in(0,1)$:
If $T$ is unknown, by taking $\eta_{t\in\left[T\right]}=\frac{1}{2L}\land\frac{\eta}{\sqrt{t}}$ with $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x^{*},x^{1})}{M^{2}+\sigma^{2}C(\delta,p)}}\right)$, then with probability at least $1-\delta$, there is
$F(x^{T+1})-F(x^{*})\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{T}+\frac{(M+\sigma\sqrt{C(\delta,p)})\sqrt{D_{\psi}(x^{*},x^{1})}\log T}{\sqrt{T}}\right).$
If $T$ is known, by taking $\eta_{t\in\left[T\right]}=\frac{1}{2L}\land\frac{\eta}{\sqrt{T}}$ with $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x^{*},x^{1})}{(M^{2}+\sigma^{2}C(\delta,p))\log T}}\right)$, then with probability at least $1-\delta$, there is
$F(x^{T+1})-F(x^{*})\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{T}+\frac{(M+\sigma\sqrt{C(\delta,p)})\sqrt{D_{\psi}(x^{*},x^{1})\log T}}{\sqrt{T}}\right).$
As far as we know, Theorem 6.2 is the first to give the last-iterate convergence under sub-Weibull noise. Its full statement is provided in Appendix H. Compared with Theorem 3.3, we only incur an extra $O\left(\mathrm{polylog}\frac{1}{\delta}\right)$ factor. However, we want to point out that $\sqrt{C(\delta,p)}$ is not continuous in $p$, as there is a difference when $p$ is approaching $1$ from two sides.
6.3.1 Optimal Rate under Sub-Weibull Noise
Theorem 6.3.
Under Assumptions 1.-4. and 5D. with $\mu_{f}=\mu_{h}=0$ and let $\delta\in(0,1)$, if $T$ is known, by taking $\eta_{t\in\left[T\right]}=\frac{T-t+1}{2LT}\land\frac{\eta(T-t+1)}{T^{\frac{3}{2}}}$ with $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x^{*},x^{1})}{M^{2}+\sigma^{2}C(\delta,p)}}\right)$, then with probability at least $1-\delta$, there is
$F(x^{T+1})-F(x^{*})\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{T}+\frac{(M+\sigma\sqrt{C(\delta,p)})\sqrt{D_{\psi}(x^{*},x^{1})}}{\sqrt{T}}\right).$
With almost the same step size in Theorem 3.5, we can also get rid of the extra $O\left(\sqrt{\log T}\right)$ factor in Theorem 6.2 to obtain the optimal dependence in time $T$. The detailed version for any positive $\eta$ can be found in Appendix H.1.
7 Conclusion
In this work, we present a unified analysis for the last-iterate convergence of stochastic gradient methods and obtain several new results. More specifically, we establish the (nearly) optimal convergence of the last iterate of the CSMD algorithm both in expectation and in high probability. Our proofs can not only handle different function classes simultaneously but also be applied to composite problems with non-Euclidean norms on general domains. We believe our work develops a deeper understanding of stochastic gradient methods. However, there still remain many directions worth exploring. For example, it could be interesting to see whether our proof can be extended to adaptive gradient methods like AdaGrad (McMahan and Streeter, 2010; Duchi et al., 2011). We leave this important question as future work and expect it to be addressed.
Appendix A Technical Lemmas
A.1 Proof of Lemma 2.1
Before giving the proof of Lemma 2.1, we need the following property of sub-Gaussian vectors. This result is already known before (see Vershynin (2018)). We provide a proof here to make the paper self-consistent.
Lemma A.1.
Given a sigma algebra $\mathcal{F}$, if $\xi\in\mathbb{R}^{d}$ is a random vector satisfying $\mathbb{\mathbb{E}}\left[\exp\left(\lambda\left\|\xi\right\|_{*}^{2}\right)\mid\mathcal{F}\right]\leq\exp\left(\lambda\sigma^{2}\right),\forall\lambda\in\left[0,\sigma^{-2}\right]$, then for any integer $k\geq 1$, we have
$\mathbb{\mathbb{E}}\left[\left\|\xi\right\|_{*}^{2k}\mid\mathcal{F}\right]\leq\begin{cases}\sigma^{2}&k=1\\ e(k!)\sigma^{2k}&k\geq 2\end{cases}.$
Proof.
For the case $k=1$, given any $\lambda\in\left[0,\sigma^{-2}\right]$, there is
$\exp\left(\mathbb{\mathbb{E}}\left[\lambda\left\|\xi\right\|_{*}^{2}\mid\mathcal{F}\right]\right)\leq\mathbb{\mathbb{E}}\left[\exp\left(\lambda\left\|\xi\right\|_{*}^{2}\right)\mid\mathcal{F}\right]\leq\exp\left(\lambda\sigma^{2}\right)\Rightarrow\mathbb{\mathbb{E}}\left[\left\|\xi\right\|_{*}^{2}\mid\mathcal{F}\right]\leq\sigma^{2}.$
For $k\geq 2,$ we have
$\mathbb{\mathbb{E}}\left[\left\|\xi\right\|_{*}^{2k}\mid\mathcal{F}\right]$ $=\mathbb{\mathbb{E}}\left[\int_{0}^{\infty}2kt^{2k-1}\mathds{1}\left[\left\|\xi\right\|_{*}\geq t\right]\mathrm{d}t\mid\mathcal{F}\right]=\int_{0}^{\infty}2kt^{2k-1}\mathbb{\mathbb{E}}\left[\mathds{1}\left[\left\|\xi\right\|_{*}\geq t\right]\mid\mathcal{F}\right]\mathrm{d}t$
$\leq\int_{0}^{\infty}2kt^{2k-1}\mathbb{\mathbb{E}}\left[\frac{\exp\left(\sigma^{-2}\left\|\xi\right\|_{*}^{2}\right)}{\exp(\sigma^{-2}t^{2})}\mid\mathcal{F}\right]\mathrm{d}t\overset{(a)}{\leq}\int_{0}^{\infty}2ekt^{2k-1}\exp(-\sigma^{-2}t^{2})\mathrm{d}t$
$\overset{(b)}{=}\int_{0}^{\infty}ek\sigma^{2k}s^{k-1}\exp(-s)\mathrm{d}s=ek\sigma^{2k}\Gamma(k)=e(k!)\sigma^{2k},$
where $(a)$ is by $\mathbb{\mathbb{E}}\left[\exp\left(\sigma^{-2}\left\|\xi\right\|_{*}^{2}\right)\mid\mathcal{F}\right]\leq\exp(\sigma^{-2}\sigma^{2})=e$ and $(b)$ is by the change of variable $t=\sigma\sqrt{s}$. ∎
Now we are ready to prove Lemma 2.1.
Proof of Lemma 2.1.
Note that
$\mathbb{\mathbb{E}}\left[\exp\left(\left\langle\xi,Z\right\rangle\right)\mid\mathcal{F}\right]$ $=\mathbb{\mathbb{E}}\left[1+\left\langle\xi,Z\right\rangle+\sum_{k=2}^{\infty}\frac{\left(\left\langle\xi,Z\right\rangle\right)^{k}}{k!}\mid\mathcal{F}\right]\leq\mathbb{\mathbb{E}}\left[\left\langle\xi,Z\right\rangle\mid\mathcal{F}\right]+\mathbb{\mathbb{E}}\left[1+\sum_{k=2}^{\infty}\frac{\left\|\xi\right\|_{*}^{k}\left\|Z\right\|^{k}}{k!}\mid\mathcal{F}\right]$
$\overset{(a)}{=}\mathbb{\mathbb{E}}\left[1+\sum_{k=1}^{\infty}\frac{\left\|\xi\right\|_{*}^{2k}\left\|Z\right\|^{2k}}{(2k)!}+\sum_{k=1}^{\infty}\frac{\left\|\xi\right\|_{*}^{2k+1}\left\|Z\right\|^{2k+1}}{(2k+1)!}\mid\mathcal{F}\right]$
$\overset{(b)}{\leq}\mathbb{\mathbb{E}}\left[1+\sum_{k=1}^{\infty}\frac{\left\|\xi\right\|_{*}^{2k}\left\|Z\right\|^{2k}}{(2k)!}+\sum_{k=1}^{\infty}\frac{\left\|\xi\right\|_{*}^{2k}\left\|Z\right\|^{2k}+\left\|\xi\right\|_{*}^{2k+2}\left\|Z\right\|^{2k+2}/4}{(2k+1)!}\mid\mathcal{F}\right]$
$=\mathbb{\mathbb{E}}\left[1+\frac{2\left\|\xi\right\|_{*}^{2}\left\|Z\right\|^{2}}{3}+\sum_{k=2}^{\infty}\left\|\xi\right\|_{*}^{2k}\left\|Z\right\|^{2k}\left(\frac{1}{4(2k-1)!}+\frac{1}{(2k)!}+\frac{1}{(2k+1)!}\right)\mid\mathcal{F}\right]$
$=\mathbb{\mathbb{E}}\left[1+\frac{2\left\|\xi\right\|_{*}^{2}\left\|Z\right\|^{2}}{3}+\sum_{k=2}^{\infty}\left\|\xi\right\|_{*}^{2k}\left\|Z\right\|^{2k}\frac{1+k/2+1/(2k+1)}{(2k)!}\mid\mathcal{F}\right]$
$\overset{(c)}{\leq}1+\frac{2\sigma^{2}\left\|Z\right\|^{2}}{3}+\sum_{k=2}^{\infty}\frac{\sigma^{2k}\left\|Z\right\|^{2k}}{k!}\cdot\frac{e(1+k/2+1/(2k+1))}{\binom{2k}{k}}$
$\overset{(d)}{\leq}1+\sigma^{2}\left\|Z\right\|^{2}+\sum_{k=2}^{\infty}\frac{\sigma^{2k}\left\|Z\right\|^{2k}}{k!}=\exp\left(\sigma^{2}\left\|Z\right\|^{2}\right),$
where $(a)$ is by $\mathbb{\mathbb{E}}\left[\left\langle\xi,Z\right\rangle\mid\mathcal{F}\right]=\left\langle\mathbb{\mathbb{E}}\left[\xi\mid\mathcal{F}\right],Z\right\rangle=0$, $(b)$ holds due to AM-GM inequality, $(c)$ is by applying Lemma A.1 to $\mathbb{\mathbb{E}}\left[\left\|\xi\right\|_{*}^{2k}\left\|Z\right\|^{2k}\mid\mathcal{F}\right]=\left\|Z\right\|^{2k}\mathbb{\mathbb{E}}\left[\left\|\xi\right\|_{*}^{2k}\mid\mathcal{F}\right]$ and $(d)$ is by $2/3<1$ and
$\max_{k\geq 2,k\in\mathbb{\mathbb{N}}}\frac{e(1+k/2+1/(2k+1))}{\binom{2k}{k}}=\frac{e(1+1+1/5)}{6}<1.$
A.2 Proof of Lemma 6.1
To prove Lemma 6.1, we need the following moment bound on sub-Weibull random variables, which could be viewed as a generalization of Lemma A.1. Similar results can also be found in, for example, Vladimirova et al. (2020).
Lemma A.2.
Given a sigma algebra $\mathcal{F}$, if $\xi\in\mathbb{R}^{d}$ is a random vector satisfying $\mathbb{\mathbb{E}}\left[\exp\left(\lambda\left\|\xi\right\|_{*}^{p}\right)\mid\mathcal{F}\right]\leq\exp\left(\lambda\sigma^{p}\right),\forall\lambda\in\left[0,\sigma^{-p}\right]$ for some $p\in\left(0,2\right)$, then for any $k\geq 1$ (not necessarily an integer), we have
$\mathbb{\mathbb{E}}\left[\left\|\xi\right\|_{*}^{k}\mid\mathcal{F}\right]\leq\begin{cases}e(k!)\sigma^{k}&k\in\mathbb{\mathbb{N}}\text{ and }p\in\left[1,2\right)\\ e\left(\frac{k}{p}\right)^{\frac{k}{p}}\sigma^{k}&k\geq p\end{cases}.$
Proof.
Following a similar way of proving Lemma A.1, we can obtain that for any $k\geq 1$,
$\mathbb{\mathbb{E}}\left[\left\|\xi\right\|_{*}^{k}\mid\mathcal{F}\right]\leq e\sigma^{k}\Gamma\left(\frac{k}{p}+1\right).$
By the recursive definition of the Gamma function, there is
$\Gamma\left(\frac{k}{p}+1\right)=\Gamma\left(\frac{k}{p}-\left\lfloor\frac{k}{p}\right\rfloor+1\right)\prod_{i=0}^{\left\lfloor\frac{k}{p}\right\rfloor-1}\left(\frac{k}{p}-i\right)\leq\prod_{i=0}^{\left\lfloor\frac{k}{p}\right\rfloor-1}\left(\frac{k}{p}-i\right),$
where the last step is by $\Gamma\left(\frac{k}{p}-\left\lfloor\frac{k}{p}\right\rfloor+1\right)\leq 1$ due to $1\leq\frac{k}{p}-\left\lfloor\frac{k}{p}\right\rfloor+1<2$. Finally, when $p\in\left[1,2\right)$ and $k\in\mathbb{\mathbb{N}}$, we can bound
$\prod_{i=0}^{\left\lfloor\frac{k}{p}\right\rfloor-1}\left(\frac{k}{p}-i\right)\leq\prod_{i=0}^{\left\lfloor\frac{k}{p}\right\rfloor-1}\left(k-i\right)\leq\prod_{i=0}^{k-1}\left(k-i\right)=k!.$
If $k\geq p$, we can bound
$\prod_{i=0}^{\left\lfloor\frac{k}{p}\right\rfloor-1}\left(\frac{k}{p}-i\right)\leq\left(\frac{k}{p}\right)^{\left\lfloor\frac{k}{p}\right\rfloor}\leq\left(\frac{k}{p}\right)^{\frac{k}{p}}.$
Now we are ready to prove Lemma 6.1.
Proof of Lemma 6.1.
Note that
$\mathbb{\mathbb{E}}\left[\exp\left(\left\langle\xi,Z\right\rangle\right)\mid\mathcal{F}\right]$ $=\mathbb{\mathbb{E}}\left[1+\left\langle\xi,Z\right\rangle+\sum_{k=2}^{\infty}\frac{\left(\left\langle\xi,Z\right\rangle\right)^{k}}{k!}\mid\mathcal{F}\right]\overset{(a)}{\leq}\mathbb{\mathbb{E}}\left[1+\sum_{k=2}^{\infty}\frac{\left\|\xi\right\|_{*}^{k}\left\|Z\right\|^{k}}{k!}\mid\mathcal{F}\right]$
$\overset{(b)}{\leq}1+e\sum_{k=2}^{\infty}\sigma^{k}\left\|Z\right\|^{k}\leq\exp\left(e\sum_{k=2}^{\infty}\sigma^{k}\left\|Z\right\|^{k}\right),$ (2)
where $(a)$ is due to $\mathbb{\mathbb{E}}\left[\left\langle\xi,Z\right\rangle\mid\mathcal{F}\right]=\left\langle\mathbb{\mathbb{E}}\left[\xi\mid\mathcal{F}\right],Z\right\rangle=0$ and $\left(\left\langle\xi,Z\right\rangle\right)^{k}\leq\left(\left|\left\langle\xi,Z\right\rangle\right|\right)^{k}\leq\left\|\xi\right\|_{*}^{k}\left\|Z\right\|^{k}$, and $(b)$ is by applying Lemma A.2 to $\mathbb{\mathbb{E}}\left[\left\|\xi\right\|_{*}^{k}\left\|Z\right\|^{k}\mid\mathcal{F}\right]=\left\|Z\right\|^{k}\mathbb{\mathbb{E}}\left[\left\|\xi\right\|_{*}^{k}\mid\mathcal{F}\right]$.
Therefore, if $\left\|Z\right\|\leq\frac{1}{2\sigma}$ almost surely, we have
$e\sum_{k=2}^{\infty}\sigma^{k}\left\|Z\right\|^{k}\leq 2e\sigma^{2}\left\|Z\right\|^{2}\leq 6\sigma^{2}\left\|Z\right\|^{2}\Rightarrow\exp\left(e\sum_{k=2}^{\infty}\sigma^{k}\left\|Z\right\|^{k}\right)\leq\exp\left(6\sigma^{2}\left\|Z\right\|^{2}\right).$ (3)
When $p\in\left(1,2\right)$, recall that $q=\frac{p}{p-1}$, we can also bound
$\left\langle\xi,Z\right\rangle\leq\left\|\xi\right\|_{*}\left\|Z\right\|\leq\frac{\left\|\xi\right\|_{*}^{p}}{2p\sigma^{p}}+\frac{2^{q/p}\sigma^{q}\left\|Z\right\|^{q}}{q},$
where the first step is due to Cauchy-Schwarz inequality and the second step is by Young’s inequality. Therefore, it can be obtained
$\mathbb{\mathbb{E}}\left[\exp\left(\left\langle\xi,Z\right\rangle\right)\mid\mathcal{F}\right]\leq\mathbb{\mathbb{E}}\left[\exp\left(\frac{\left\|\xi\right\|_{*}^{p}}{2p\sigma^{p}}+\frac{2^{q/p}\sigma^{q}\left\|Z\right\|^{q}}{q}\right)\mid\mathcal{F}\right]\leq\exp\left(\frac{1}{2p}+\frac{2^{q/p}\sigma^{q}\left\|Z\right\|^{q}}{q}\right),$ (4)
where the last step is due to $Z\in\mathcal{F}$ and $\mathbb{\mathbb{E}}\left[\exp\left(\frac{\left\|\xi\right\|_{*}^{p}}{2p\sigma^{p}}\right)\mid\mathcal{F}\right]\leq\exp\left(\frac{1}{2p}\right)$ since $\frac{1}{2p}\leq 1$. Now, let us define the event $A\coloneqq\left\{\left\|Z\right\|\leq\frac{1}{2\sigma}\right\}$ and bound
$\mathbb{\mathbb{E}}\left[\exp\left(\left\langle\xi,Z\right\rangle\right)\mid\mathcal{F}\right]$ $=\mathbb{\mathbb{E}}\left[\exp\left(\left\langle\xi,Z\right\rangle\right)\mid\mathcal{F}\right]\left(\mathds{1}\left[A\right]+\mathds{1}\left[A^{c}\right]\right)$
[…]
$\overset{(c)}{\leq}\exp\left(6\sigma^{2}\left\|Z\right\|^{2}\right)\mathds{1}\left[A\right]+\exp\left(2^{q-1}\sigma^{q}\left\|Z\right\|^{q}\right)\mathds{1}\left[A^{c}\right]\leq\exp\left(6\sigma^{2}\left\|Z\right\|^{2}+2^{q-1}\sigma^{q}\left\|Z\right\|^{q}\right),$
where in $(c)$ we use (3) when $\left\|Z\right\|\leq\frac{1}{2\sigma}$, and $\frac{1}{2p}+\frac{2^{q/p}\sigma^{q}\left\|Z\right\|^{q}}{q}\leq\frac{2^{q}\sigma^{q}\left\|Z\right\|^{q}}{2p}+\frac{2^{q/p}\sigma^{q}\left\|Z\right\|^{q}}{q}=2^{q-1}\sigma^{q}\left\|Z\right\|^{q}$ when $\left\|Z\right\|>\frac{1}{2\sigma}$, where the last equation is by q/p=q1q/p=q-1 and $1/p+1/q=1$. ∎
Appendix B Missing Proofs in Section 4
In this section, we provide the missing proofs of the three most important lemmas.
B.1 Proof of Lemma 4.1
Proof of Lemma 4.1.
Inspired by Zamani and Glineur (2025), we first introduce the following auxiliary sequence
$z^{t}=\begin{cases}\left(1-\frac{v_{t-1}}{v_{t}}\right)x^{t}+\frac{v_{t-1}}{v_{t}}z^{t-1}&t\in\left[T\right]\\ x&t=0\end{cases}\Leftrightarrow z^{t}=\frac{v_{0}}{v_{t}}x+\sum_{s=1}^{t}\frac{v_{s}-v_{s-1}}{v_{t}}x^{s},\forall t\in\left\{0\right\}\cup\left[T\right],$
where we recall that $v_{t\in\left\{0\right\}\cup\left[T\right]}=\frac{w_{T}\gamma_{T}}{\sum_{s=t\lor 1}^{T}w_{s}\gamma_{s}}\geq 0$ is non-decreasing. Note that $z^{t}$ always falls in the domain $\mathcal{X}$ by its definition.
Now, we start the proof from the $(L,M)$-smoothness of $f$,
$f(x^{t+1})-f(x^{t})\leq$ $\left\langle g^{t},x^{t+1}-x^{t}\right\rangle+\frac{L}{2}\left\|x^{t+1}-x^{t}\right\|^{2}+M\left\|x^{t+1}-x^{t}\right\|$
$=$ $\left\langle\xi^{t},z^{t}-x^{t}\right\rangle+\underbrace{\left\langle\xi^{t},x^{t}-x^{t+1}\right\rangle}_{\text{I}}+\underbrace{\left\langle\widehat{g}^{t},x^{t+1}-z^{t}\right\rangle}_{\text{II}}$
$+\underbrace{\left\langle g^{t},z^{t}-x^{t}\right\rangle}_{\text{III}}+\underbrace{\frac{L}{2}\left\|x^{t+1}-x^{t}\right\|^{2}+M\left\|x^{t+1}-x^{t}\right\|}_{\text{IV}},$ (5)
where $g^{t}\coloneqq\mathbb{\mathbb{E}}\left[\widehat{g}^{t}|\mathcal{F}^{t-1}\right]\in\partial f(x^{t})$ and $\xi^{t}=\widehat{g}^{t}-g^{t}$. Next, we bound these four terms respectively.
  • For term I, by applying Cauchy-Schwarz inequality, the $1$-strong convexity of $\psi$ and AM-GM inequality, we can get the following upper bound
    $\text{I}\leq\left\|\xi^{t}\right\|_{*}\left\|x^{t}-x^{t+1}\right\|\leq\left\|\xi^{t}\right\|_{*}\sqrt{2D_{\psi}(x^{t+1},x^{t})}\leq 2\eta_{t}\left\|\xi^{t}\right\|_{*}^{2}+\frac{D_{\psi}(x^{t+1},x^{t})}{4\eta_{t}}.$ (6)
  • For term II, we recall that the update rule is $x^{t+1}=\mathrm{argmin}_{x\in\mathcal{X}}h(x)+\left\langle\widehat{g}^{t},x-x^{t}\right\rangle+\frac{D_{\psi}(x,x^{t})}{\eta_{t}}$. Hence, by the optimality condition of $x^{t+1}$, there exists $h^{t+1}\in\partial h(x^{t+1})$ such that for any $y\in\mathcal{X}$
    $\left\langle h^{t+1}+\widehat{g}^{t}+\frac{\nabla\psi(x^{t+1})-\nabla\psi(x^{t})}{\eta_{t}},x^{t+1}-y\right\rangle\leq 0,$
    which implies
    $\left\langle\widehat{g}^{t},x^{t+1}-y\right\rangle$ $\leq\frac{\left\langle\nabla\psi(x^{t})-\nabla\psi(x^{t+1}),x^{t+1}-y\right\rangle}{\eta_{t}}+\left\langle h^{t+1},y-x^{t+1}\right\rangle$
    $\leq\frac{D_{\psi}(y,x^{t})-D_{\psi}(y,x^{t+1})-D_{\psi}(x^{t+1},x^{t})}{\eta_{t}}+h(y)-h(x_{t+1})-\mu_{h}D_{\psi}(y,x^{t+1}),$
    where the last inequality holds due to $\left\langle\nabla\psi(x^{t})-\nabla\psi(x^{t+1}),x^{t+1}-y\right\rangle=D_{\psi}(y,x^{t})-D_{\psi}(y,x^{t+1})-D_{\psi}(x^{t+1},x^{t})$ and $\left\langle h^{t+1},y-x^{t+1}\right\rangle\leq h(y)-h(x_{t+1})-\mu_{h}D_{\psi}(y,x^{t+1})$ by the $\mu_{h}$-strong convexity of $h$. We substitute $y$ with $z^{t}$ to obtain
    $\text{II}\leq\frac{D_{\psi}(z^{t},x^{t})-D_{\psi}(z^{t},x^{t+1})-D_{\psi}(x^{t+1},x^{t})}{\eta_{t}}+h(z_{t})-h(x_{t+1})-\mu_{h}D_{\psi}(z^{t},x^{t+1}).$ (7)
  • For term III, we simply use the $\mu_{f}$-strong convexity of $f$ to get
    $\text{III}\leq f(z^{t})-f(x^{t})-\mu_{f}D_{\psi}(z^{t},x^{t}).$ (8)
  • For term IV, we have
    IV $\leq LD_{\psi}(x^{t+1},x^{t})+M\sqrt{2D_{\psi}(x^{t+1},x^{t})}$
    $\leq LD_{\psi}(x^{t+1},x^{t})+2\eta_{t}M^{2}+\frac{D_{\psi}(x^{t+1},x^{t})}{4\eta_{t}},$ (9)
    where the first inequality holds by the $1$-strong convexity of $\psi$ again and the second one is due to AM-GM inequality.
By plugging the bounds (6), (7), (8) and (9) into (5), we obtain
$f(x^{t+1})-f(x^{t})\leq$ $\left\langle\xi^{t},z^{t}-x^{t}\right\rangle+2\eta_{t}\left\|\xi^{t}\right\|_{*}^{2}+\frac{D_{\psi}(x^{t+1},x^{t})}{4\eta_{t}}$
$+\frac{D_{\psi}(z^{t},x^{t})-D_{\psi}(z^{t},x^{t+1})-D_{\psi}(x^{t+1},x^{t})}{\eta_{t}}+h(z_{t})-h(x_{t+1})-\mu_{h}D_{\psi}(z^{t},x^{t+1})$
$+f(z^{t})-f(x^{t})-\mu_{f}D_{\psi}(z^{t},x^{t})+LD_{\psi}(x^{t+1},x^{t})+2\eta_{t}M^{2}+\frac{D_{\psi}(x^{t+1},x^{t})}{4\eta_{t}}.$
Rearranging the terms to get
$F(x^{t+1})-F(z^{t})$
$\leq$ $\left\langle\xi^{t},z^{t}-x^{t}\right\rangle+(\eta_{t}^{-1}-\mu_{f})D_{\psi}(z^{t},x^{t})-(\eta_{t}^{-1}+\mu_{h})D_{\psi}(z^{t},x^{t+1})$
$+\left(L-\frac{1}{2\eta_{t}}\right)D_{\psi}(x^{t+1},x^{t})+2\eta_{t}(M^{2}+\left\|\xi^{t}\right\|_{*}^{2})$
$\overset{(a)}{\leq}$ $\left\langle\xi^{t},z^{t}-x^{t}\right\rangle+(\eta_{t}^{-1}-\mu_{f})D_{\psi}(z^{t},x^{t})-(\eta_{t}^{-1}+\mu_{h})D_{\psi}(z^{t},x^{t+1})+2\eta_{t}(M^{2}+\left\|\xi^{t}\right\|_{*}^{2})$
$\overset{(b)}{=}$ $\frac{v_{t-1}}{v_{t}}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle+(\eta_{t}^{-1}-\mu_{f})D_{\psi}(z^{t},x^{t})-(\eta_{t}^{-1}+\mu_{h})D_{\psi}(z^{t},x^{t+1})+2\eta_{t}(M^{2}+\left\|\xi^{t}\right\|_{*}^{2})$
$\overset{(c)}{\leq}$ $\frac{v_{t-1}}{v_{t}}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle+(\eta_{t}^{-1}-\mu_{f})\frac{v_{t-1}}{v_{t}}D_{\psi}(z^{t-1},x^{t})-(\eta_{t}^{-1}+\mu_{h})D_{\psi}(z^{t},x^{t+1})+2\eta_{t}(M^{2}+\left\|\xi^{t}\right\|_{*}^{2}),$ (10)
where $(a)$ is by $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}\leq\frac{1}{2L}\Rightarrow L-\frac{1}{2\eta_{t}}\leq 0$, $(b)$ holds due to the definition of $z^{t}=\left(1-\frac{v_{t-1}}{v_{t}}\right)x^{t}+\frac{v_{t-1}}{v_{t}}z^{t-1}$ implying $z^{t}-x^{t}=\frac{v_{t-1}}{v_{t}}\left(z^{t-1}-x^{t}\right)$, $(c)$ is by noticing $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}\leq\frac{1}{\mu_{f}}\Rightarrow\eta_{t}^{-1}-\mu_{f}\geq 0$ and
$D_{\psi}(z^{t},x^{t})\overset{(d)}{\leq}\left(1-\frac{v_{t-1}}{v_{t}}\right)D_{\psi}(x^{t},x^{t})+\frac{v_{t-1}}{v_{t}}D_{\psi}(z^{t-1},x^{t})=\frac{v_{t-1}}{v_{t}}D_{\psi}(z^{t-1},x^{t}),$
where $(d)$ is by the convexity of the first argument in $D_{\psi}(\cdot,\cdot)$.
Multiplying both sides of (10) by $w_{t}\gamma_{t}v_{t}$ (all of these three terms are non-negative) and summing up from $t=1$ to $T$, we obtain
$\sum_{t=1}^{T}w_{t}\gamma_{t}v_{t}\left(F(x^{t+1})-F(z^{t})\right)$
$\leq$ $\sum_{t=1}^{T}w_{t}\gamma_{t}v_{t-1}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle+\sum_{t=1}^{T}2w_{t}\gamma_{t}\eta_{t}v_{t}(M^{2}+\left\|\xi^{t}\right\|_{*}^{2})$
$+\sum_{t=1}^{T}w_{t}\gamma_{t}(\eta_{t}^{-1}-\mu_{f})v_{t-1}D_{\psi}(z^{t-1},x^{t})-w_{t}\gamma_{t}(\eta_{t}^{-1}+\mu_{h})v_{t}D_{\psi}(z^{t},x^{t+1})$
$=$ $w_{1}\gamma_{1}(\eta_{1}^{-1}-\mu_{f})v_{0}D_{\psi}(z^{0},x^{1})-w_{T}\gamma_{T}(\eta_{T}^{-1}+\mu_{h})v_{T}D_{\psi}(z^{T},x^{T+1})+\sum_{t=1}^{T}2w_{t}\gamma_{t}\eta_{t}v_{t}(M^{2}+\left\|\xi^{t}\right\|_{*}^{2})$
$+\sum_{t=1}^{T}w_{t}\gamma_{t}v_{t-1}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle+\sum_{t=2}^{T}\left(w_{t}\gamma_{t}(\eta_{t}^{-1}-\mu_{f})-w_{t-1}\gamma_{t-1}(\eta_{t-1}^{-1}+\mu_{h})\right)v_{t-1}D_{\psi}(z^{t-1},x^{t})$
$\overset{(e)}{=}$ $w_{1}(1-\mu_{f}\eta_{1})v_{0}D_{\psi}(x,x^{1})-w_{T}\gamma_{T}(\eta_{T}^{-1}+\mu_{h})v_{T}D_{\psi}(z^{T},x^{T+1})+\sum_{t=1}^{T}2w_{t}\gamma_{t}\eta_{t}v_{t}(M^{2}+\left\|\xi^{t}\right\|_{*}^{2})$
$+\sum_{t=1}^{T}w_{t}\gamma_{t}v_{t-1}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle+\sum_{t=2}^{T}(w_{t}-w_{t-1})\gamma_{t}(\eta_{t}^{-1}-\mu_{f})v_{t-1}D_{\psi}(z^{t-1},x^{t}),$ (11)
where $(e)$ holds due to $\gamma_{1}(\eta_{1}^{-1}-\mu_{f})=\eta_{1}(\eta_{1}^{-1}-\mu_{f})=1-\mu_{f}\eta_{1}$, $z^{0}=x$ and $\gamma_{t}(\eta_{t}^{-1}-\mu_{f})=\eta_{t}(\eta_{t}^{-1}-\mu_{f})\prod_{s=2}^{t}\frac{1+\mu_{h}\eta_{s-1}}{1-\mu_{f}\eta_{s}}=(\eta_{t-1}^{-1}+\mu_{h})\eta_{t-1}\prod_{s=2}^{t-1}\frac{1+\mu_{h}\eta_{s-1}}{1-\mu_{f}\eta_{s}}=\gamma_{t-1}(\eta_{t-1}^{-1}+\mu_{h}),\forall t\geq 2$.
By the convexity of $F$ and the definition of $z^{t}=\frac{v_{0}}{v_{t}}x+\sum_{s=1}^{t}\frac{v_{s}-v_{s-1}}{v_{t}}x^{s}$, we have
$F(z^{t})\leq\sum_{s=1}^{t}\frac{v_{s}-v_{s-1}}{v_{t}}F(x^{s})+\frac{v_{0}}{v_{t}}F(x),$
which implies
$\sum_{t=1}^{T}w_{t}\gamma_{t}v_{t}\left(F(x^{t+1})-F(z^{t})\right)\geq$ $\sum_{t=1}^{T}\left[w_{t}\gamma_{t}v_{t}F(x^{t+1})-w_{t}\gamma_{t}\left(\sum_{s=1}^{t}(v_{s}-v_{s-1})F(x^{s})+v_{0}F(x)\right)\right]$
$=$ $\sum_{t=1}^{T}\left[w_{t}\gamma_{t}v_{t}\left(F(x^{t+1})-F(x)\right)-w_{t}\gamma_{t}\sum_{s=1}^{t}(v_{s}-v_{s-1})\left(F(x^{s})-F(x)\right)\right]$
$=$ $w_{T}\gamma_{T}v_{T}\left(F(x^{T+1})-F(x)\right)-\left(\sum_{t=1}^{T}w_{t}\gamma_{t}\right)(v_{1}-v_{0})\left(F(x^{1})-F(x)\right)$
$+\sum_{t=2}^{T}\left[w_{t-1}\gamma_{t-1}v_{t-1}-\left(\sum_{s=t}^{T}w_{s}\gamma_{s}\right)(v_{t}-v_{t-1})\right]\left(F(x^{t})-F(x)\right).$
Now, by the definition of $v_{t\in\left\{0\right\}\cup\left[T\right]}=\frac{w_{T}\gamma_{T}}{\sum_{s=t\lor 1}^{T}w_{s}\gamma_{s}}$, we observe that
$\left(\sum_{t=1}^{T}w_{t}\gamma_{t}\right)(v_{1}-v_{0})=\left(\sum_{t=1}^{T}w_{t}\gamma_{t}\right)\left(\frac{w_{T}\gamma_{T}}{\sum_{s=1}^{T}w_{s}\gamma_{s}}-\frac{w_{T}\gamma_{T}}{\sum_{s=1}^{T}w_{s}\gamma_{s}}\right)=0,$
and for $2\leq t\leq T$,
$w_{t-1}\gamma_{t-1}v_{t-1}-\left(\sum_{s=t}^{T}w_{s}\gamma_{s}\right)(v_{t}-v_{t-1})=\left(\sum_{s=t-1}^{T}w_{s}\gamma_{s}\right)v_{t-1}-\left(\sum_{s=t}^{T}w_{s}\gamma_{s}\right)v_{t}$
$=$ $\left(\sum_{s=t-1}^{T}w_{s}\gamma_{s}\right)\frac{w_{T}\gamma_{T}}{\sum_{s=t-1}^{T}w_{s}\gamma_{s}}-\left(\sum_{s=t}^{T}w_{s}\gamma_{s}\right)\frac{w_{T}\gamma_{T}}{\sum_{s=t}^{T}w_{s}\gamma_{s}}=0.$
These two equations immediately imply
$\sum_{t=1}^{T}w_{t}\gamma_{t}v_{t}\left(F(x^{t+1})-F(z^{t})\right)\geq w_{T}\gamma_{T}v_{T}\left(F(x^{T+1})-F(x)\right).$ (12)
Plugging (12) into (11), we finally get
$w_{T}\gamma_{T}v_{T}\left(F(x^{T+1})-F(x)\right)$
$\leq$ $w_{1}(1-\mu_{f}\eta_{1})v_{0}D_{\psi}(x,x^{1})-w_{T}\gamma_{T}(\eta_{T}^{-1}+\mu_{h})v_{T}D_{\psi}(z^{T},x^{T+1})+\sum_{t=1}^{T}2w_{t}\gamma_{t}\eta_{t}v_{t}(M^{2}+\left\|\xi^{t}\right\|_{*}^{2})$
$+\sum_{t=1}^{T}w_{t}\gamma_{t}v_{t-1}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle+\sum_{t=2}^{T}(w_{t}-w_{t-1})\gamma_{t}(\eta_{t}^{-1}-\mu_{f})v_{t-1}D_{\psi}(z^{t-1},x^{t})$
$\leq$ $w_{1}(1-\mu_{f}\eta_{1})v_{0}D_{\psi}(x,x^{1})+\sum_{t=1}^{T}2w_{t}\gamma_{t}\eta_{t}v_{t}(M^{2}+\left\|\xi^{t}\right\|_{*}^{2})$
$+\sum_{t=1}^{T}w_{t}\gamma_{t}v_{t-1}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle+\sum_{t=2}^{T}(w_{t}-w_{t-1})\gamma_{t}(\eta_{t}^{-1}-\mu_{f})v_{t-1}D_{\psi}(z^{t-1},x^{t}).$
B.2 Proof of Lemma 4.2
Proof of Lemma 4.2.
We invoke Lemma 4.1 with $w_{t\in\left[T\right]}=1$ to get
$\gamma_{T}v_{T}\left(F(x^{T+1})-F(x)\right)\leq(1-\mu_{f}\eta_{1})v_{0}D_{\psi}(x,x^{1})+\sum_{t=1}^{T}2\gamma_{t}\eta_{t}v_{t}(M^{2}+\left\|\xi^{t}\right\|_{*}^{2})+\sum_{t=1}^{T}\gamma_{t}v_{t-1}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle.$
Take expectations on both sides to obtain
$\gamma_{T}v_{T}\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$
$\leq$ $(1-\mu_{f}\eta_{1})v_{0}D_{\psi}(x,x^{1})+\sum_{t=1}^{T}2\gamma_{t}\eta_{t}v_{t}(M^{2}+\mathbb{\mathbb{E}}\left[\left\|\xi^{t}\right\|_{*}^{2}\right])+\sum_{t=1}^{T}\gamma_{t}v_{t-1}\mathbb{\mathbb{E}}\left[\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle\right]$
$\leq$ $(1-\mu_{f}\eta_{1})v_{0}D_{\psi}(x,x^{1})+\sum_{t=1}^{T}2\gamma_{t}\eta_{t}v_{t}(M^{2}+\sigma^{2}),$
where the last line is due to $\mathbb{\mathbb{E}}\left[\left\|\xi^{t}\right\|_{*}^{2}\right]=\mathbb{\mathbb{E}}\left[\mathbb{\mathbb{E}}\left[\left\|\xi^{t}\right\|_{*}^{2}\mid\mathcal{F}^{t-1}\right]\right]\leq\sigma^{2}$ (Assumption 5A.) and $\mathbb{\mathbb{E}}\left[\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle\right]=\mathbb{\mathbb{E}}\left[\left\langle\mathbb{\mathbb{E}}\left[\xi^{t}|\mathcal{F}^{t-1}\right],z^{t-1}-x^{t}\right\rangle\right]=0$ ($z^{t-1}-x^{t}\in\mathcal{F}^{t-1}$ and Assumption 4.). Finally, we divide both sides by $\gamma_{T}v_{T}$ and plug in $v_{t\in\left\{0\right\}\cup\left[T\right]}=\frac{w_{T}\gamma_{T}}{\sum_{s=t\lor 1}^{T}w_{s}\gamma_{s}}=\frac{\gamma_{T}}{\sum_{s=t\lor 1}^{T}\gamma_{s}}$ to finish the proof. ∎
B.3 Proof of Lemma 4.3
Proof of Lemma 4.3.
We invoke Lemma 4.1 to get
$w_{T}\gamma_{T}v_{T}\left(F(x^{T+1})-F(x)\right)\leq$ $w_{1}(1-\mu_{f}\eta_{1})v_{0}D_{\psi}(x,x^{1})+\sum_{t=1}^{T}2w_{t}\gamma_{t}\eta_{t}v_{t}(M^{2}+\left\|\xi^{t}\right\|_{*}^{2})$
$+\sum_{t=1}^{T}w_{t}\gamma_{t}v_{t-1}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle+\sum_{t=2}^{T}(w_{t}-w_{t-1})\gamma_{t}(\eta_{t}^{-1}-\mu_{f})v_{t-1}D_{\psi}(z^{t-1},x^{t}).$ (13)
Let $w_{t\in\left[T\right]}$ be defined as follows
$w_{t\in\left[T\right]}\coloneqq\frac{1}{\sum_{s=2}^{t}\frac{2\gamma_{s}\eta_{s}\bar{v}_{s}\sigma^{2}}{1-\mu_{f}\eta_{s}}+\sum_{s=1}^{T}2\gamma_{s}\eta_{s}\bar{v}_{s}\sigma^{2}},$ (14)
where
$\bar{v}_{t\in\left\{0\right\}\cup\left[T\right]}\coloneqq\frac{\gamma_{T}}{\sum_{s=t\lor 1}^{T}\gamma_{s}}.$ (15)
Note that $w_{t\in\left[T\right]}\geq 0$ is non-increasing, from the definition of $v_{t\in\left\{0\right\}\cup\left[T\right]}=\frac{w_{T}\gamma_{T}}{\sum_{s=t\lor 1}^{T}w_{s}\gamma_{s}}$, there is always
$v_{t}=\frac{w_{T}\gamma_{T}}{\sum_{s=t\lor 1}^{T}w_{s}\gamma_{s}}\leq\frac{\gamma_{T}}{\sum_{s=t\lor 1}^{T}\gamma_{s}}=\bar{v}_{t},\forall t\in\left\{0\right\}\cup\left[T\right].$ (16)
Now we consider the following non-negative sequence with $U_{0}\coloneqq 1$ and
$U_{s}\coloneqq\exp\left(\sum_{t=1}^{s}2w_{t}\gamma_{t}\eta_{t}v_{t}\left\|\xi^{t}\right\|_{*}^{2}-2w_{t}\gamma_{t}\eta_{t}v_{t}\sigma^{2}\right)\in\mathcal{F}^{s},\forall s\in\left[T\right].$
We claim $U_{t}$ is a supermartingale by observing that
$\mathbb{\mathbb{E}}\left[U_{t}\mid\mathcal{F}^{t-1}\right]$ $=U_{t-1}\mathbb{\mathbb{E}}\left[\exp\left(2w_{t}\gamma_{t}\eta_{t}v_{t}\left\|\xi^{t}\right\|_{*}^{2}-2w_{t}\gamma_{t}\eta_{t}v_{t}\sigma^{2}\right)\mid\mathcal{F}^{t-1}\right]$
$\overset{(a)}{\leq}U_{t-1}\exp\left(2w_{t}\gamma_{t}\eta_{t}v_{t}\sigma^{2}-2w_{t}\gamma_{t}\eta_{t}v_{t}\sigma^{2}\right)=U_{t-1},$
where $(a)$ holds due to Assumption 5B. by noticing
[…]
Hence, we know $\mathbb{\mathbb{E}}\left[U_{T}\right]\leq U_{0}=1$. Thus, there is
$\Pr\left[U_{T}>\frac{2}{\delta}\right]\overset{(b)}{\leq}\frac{\delta}{2}\mathbb{\mathbb{E}}\left[U_{T}\right]\leq\frac{\delta}{2}$
$\Rightarrow$ $\Pr\left[\sum_{t=1}^{T}2w_{t}\gamma_{t}\eta_{t}v_{t}\left\|\xi^{t}\right\|_{*}^{2}\leq\sum_{t=1}^{T}2w_{t}\gamma_{t}\eta_{t}v_{t}\sigma^{2}+\log\frac{2}{\delta}\right]\geq 1-\frac{\delta}{2},$ (17)
where we use Markov’s inequality in $(b)$.
Next, we consider another non-negative sequence with $R_{0}\coloneqq 1$ and
$R_{s}\coloneqq\exp\left(\sum_{t=1}^{s}w_{t}\gamma_{t}v_{t-1}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle-w_{t}^{2}\gamma_{t}^{2}v_{t-1}^{2}\sigma^{2}\left\|z^{t-1}-x^{t}\right\|^{2}\right)\in\mathcal{F}^{s},\forall s\in\left[T\right].$
We prove that $R_{t}$ is also a supermartingale by
$\mathbb{\mathbb{E}}\left[R_{t}\mid\mathcal{F}^{t-1}\right]$ $=R_{t-1}\mathbb{\mathbb{E}}\left[\exp\left(w_{t}\gamma_{t}v_{t-1}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle-w_{t}^{2}\gamma_{t}^{2}v_{t-1}^{2}\sigma^{2}\left\|z^{t-1}-x^{t}\right\|^{2}\right)\mid\mathcal{F}^{t-1}\right]$
$\overset{(c)}{\leq}R_{t-1}\exp\left(w_{t}^{2}\gamma_{t}^{2}v_{t-1}^{2}\sigma^{2}\left\|z^{t-1}-x^{t}\right\|^{2}-w_{t}^{2}\gamma_{t}^{2}v_{t-1}^{2}\sigma^{2}\left\|z^{t-1}-x^{t}\right\|^{2}\right)=R_{t-1},$
where $(c)$ is by applying Lemma 2.1 (note that $z^{t-1}-x^{t}\in\mathcal{F}^{t-1}=\sigma(\widehat{g}^{s},s\in\left[t-1\right])$). Hence, we have $\mathbb{\mathbb{E}}\left[R_{T}\right]\leq R_{0}=1$, which immediately implies
$\Pr\left[R_{T}>\frac{2}{\delta}\right]\overset{(d)}{\leq}\frac{\delta}{2}\mathbb{\mathbb{E}}\left[R_{T}\right]\leq\frac{\delta}{2}$
$\Rightarrow$ $\Pr\left[\sum_{t=1}^{T}w_{t}\gamma_{t}v_{t-1}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle\leq\sum_{t=1}^{T}w_{t}^{2}\gamma_{t}^{2}v_{t-1}^{2}\sigma^{2}\left\|z^{t-1}-x^{t}\right\|^{2}+\log\frac{2}{\delta}\right]\geq 1-\frac{\delta}{2}$
$\Rightarrow$ $\Pr\left[\sum_{t=1}^{T}w_{t}\gamma_{t}v_{t-1}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle\leq\sum_{t=1}^{T}2w_{t}^{2}\gamma_{t}^{2}v_{t-1}^{2}\sigma^{2}D_{\psi}(z^{t-1},x^{t})+\log\frac{2}{\delta}\right]\geq 1-\frac{\delta}{2},$ (18)
where $(d)$ is by Markov’s inequality and the last line is due to $\left\|z^{t-1}-x^{t}\right\|^{2}\leq 2D_{\psi}(z^{t-1},x^{t})$ from the $1$-strong convexity of $\psi$.
Combining (13), (17) and (18), with probability at least $1-\delta$, there is
$w_{T}\gamma_{T}v_{T}\left(F(x^{T+1})-F(x)\right)\leq$ $w_{1}(1-\mu_{f}\eta_{1})v_{0}D_{\psi}(x,x^{1})+2\log\frac{2}{\delta}+\sum_{t=1}^{T}2w_{t}\gamma_{t}\eta_{t}v_{t}(M^{2}+\sigma^{2})$
$+\sum_{t=1}^{T}2w_{t}^{2}\gamma_{t}^{2}v_{t-1}^{2}\sigma^{2}D_{\psi}(z^{t-1},x^{t})+\sum_{t=2}^{T}(w_{t}-w_{t-1})\gamma_{t}(\eta_{t}^{-1}-\mu_{f})v_{t-1}D_{\psi}(z^{t-1},x^{t})$
$=$ $\left[w_{1}(1-\mu_{f}\eta_{1})v_{0}+2w_{1}^{2}\gamma_{1}^{2}v_{0}^{2}\sigma^{2}\right]D_{\psi}(x,x^{1})+2\log\frac{2}{\delta}+\sum_{t=1}^{T}2w_{t}\gamma_{t}\eta_{t}v_{t}(M^{2}+\sigma^{2})$
$+\sum_{t=2}^{T}\left[(w_{t}-w_{t-1})(\eta_{t}^{-1}-\mu_{f})+2w_{t}^{2}\gamma_{t}v_{t-1}\sigma^{2}\right]\gamma_{t}v_{t-1}D_{\psi}(z^{t-1},x^{t}).$
Observe that for $t\geq 2$
$(w_{t}-w_{t-1})(\eta_{t}^{-1}-\mu_{f})+2w_{t}^{2}\gamma_{t}v_{t-1}\sigma^{2}$
[…]
$=$ $2w_{t}^{2}\gamma_{t}v_{t-1}\sigma^{2}-w_{t}w_{t-1}\times\frac{2\gamma_{t}\eta_{t}\bar{v}_{t}\sigma^{2}}{1-\mu_{f}\eta_{t}}\times(\eta_{t}^{-1}-\mu_{f})=2w_{t}(w_{t}v_{t-1}-w_{t-1}\bar{v}_{t})\gamma_{t}\sigma^{2}\leq 0,$
where the last line holds due to $w_{t}\leq w_{t-1}$ and $v_{t-1}\leq v_{t}\leq\bar{v}_{t}$. So, we know
$w_{T}\gamma_{T}v_{T}\left(F(x^{T+1})-F(x)\right)$ $\leq\left[w_{1}(1-\mu_{f}\eta_{1})v_{0}+2w_{1}^{2}\gamma_{1}^{2}v_{0}^{2}\sigma^{2}\right]D_{\psi}(x,x^{1})+2\log\frac{2}{\delta}+\sum_{t=1}^{T}2w_{t}\gamma_{t}\eta_{t}v_{t}(M^{2}+\sigma^{2})$
$\overset{(e)}{\leq}w_{1}\left(1-\mu_{f}\eta_{1}+2w_{1}\gamma_{1}^{2}v_{0}\sigma^{2}\right)v_{0}D_{\psi}(x,x^{1})+2\log\frac{2}{\delta}+w_{1}\sum_{t=1}^{T}2\gamma_{t}\eta_{t}v_{t}(M^{2}+\sigma^{2}),$
where $(e)$ is by $w_{t}\leq w_{1},\forall t\in\left[T\right]$. Dividing both sides by $w_{T}\gamma_{T}v_{T}$, we get
$F(x^{T+1})-F(x)$ $\leq\frac{w_{1}}{w_{T}}\left[(1-\mu_{f}\eta_{1}+2w_{1}\gamma_{1}^{2}v_{0}\sigma^{2})\frac{v_{0}}{\gamma_{T}v_{T}}D_{\psi}(x,x^{1})+\frac{2}{w_{1}\gamma_{T}v_{T}}\log\frac{2}{\delta}+2\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}v_{t}}{\gamma_{T}v_{T}}(M^{2}+\sigma^{2})\right]$
$\overset{(f)}{\leq}\left(1+\max_{2\leq t\leq T}\frac{1}{1-\mu_{f}\eta_{t}}\right)\left[\frac{(2-\mu_{f}\eta_{1})D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+2\left(M^{2}+\sigma^{2}\left(1+2\log\frac{2}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}\right]$
$\leq 2\left(1+\max_{2\leq t\leq T}\frac{1}{1-\mu_{f}\eta_{t}}\right)\left[\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+\left(M^{2}+\sigma^{2}\left(1+2\log\frac{2}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}\right],$
where $(f)$ holds due to the following calculations
[…]
Hence, the proof is completed. ∎
Appendix C General Convex Functions
In this section, we present the full version of the theorems for general convex functions (i.e., $\mu_{f}=\mu_{h}=0$) with their proofs.
Theorem C.1.
Under Assumptions 2.-4. and 5A. with $\mu_{f}=\mu_{h}=0$, for any $x\in\mathcal{X}$:
If $T$ is unknown, by taking $\eta_{t\in\left[T\right]}=\frac{1}{2L}\land\frac{\eta}{\sqrt{t}}$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x,x^{1})}{\eta}+\eta(M^{2}+\sigma^{2})\log T\right]\right).$
In particular, by choosing $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x,x^{1})}{M^{2}+\sigma^{2}}}\right)$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{(M+\sigma)\sqrt{D_{\psi}(x,x^{1})}\log T}{\sqrt{T}}\right).$
If $T$ is known, by taking $\eta_{t\in\left[T\right]}=\frac{1}{2L}\land\frac{\eta}{\sqrt{T}}$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x,x^{1})}{\eta}+\eta(M^{2}+\sigma^{2})\log T\right]\right).$
In particular, by choosing $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x,x^{1})}{(M^{2}+\sigma^{2})\log T}}\right)$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{(M+\sigma)\sqrt{D_{\psi}(x,x^{1})\log T}}{\sqrt{T}}\right).$
Proof.
From Lemma 4.2, if $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq\frac{(1-\mu_{f}\eta_{1})D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}},$ (19)
where $\gamma_{t\in\left[T\right]}=\eta_{t}\prod_{s=2}^{t}\frac{1+\mu_{h}\eta_{s-1}}{1-\mu_{f}\eta_{s}}$. Note that $\mu_{f}=\mu_{h}=0$ now, so both $\eta_{t\in\left[T\right]}=\frac{1}{2L}\land\frac{\eta}{\sqrt{t}}$ and $\eta_{t\in\left[T\right]}=\frac{1}{2L}\land\frac{\eta}{\sqrt{T}}$ satisfy $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}=\frac{1}{2L}$. Besides, $\gamma_{t\in\left[T\right]}$ will degenerate to $\eta_{t\in\left[T\right]}$. Therefore, (19) can be simplified into
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\eta_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\sum_{s=t}^{T}\eta_{s}}.$ (20)
Before proving convergence rates for these two different step sizes, we first recall some standard results.
$\sum_{t=1}^{T}\frac{1}{\sqrt{t}}$ $=\sum_{t=1}^{T}\sqrt{t}-\frac{t-1}{\sqrt{t}}=\sqrt{T}+\sum_{t=1}^{T-1}\sqrt{t}-\frac{t}{\sqrt{t+1}}\geq\sqrt{T};$ (21)
$\sum_{s=t}^{T}\frac{1}{\sqrt{s}}$ $\geq\int_{t}^{T+1}\frac{1}{\sqrt{s}}\mathrm{d}s=2(\sqrt{T+1}-\sqrt{t}),\forall t\in\left[T\right];$ (22)
$\sum_{t=1}^{T}\frac{1}{t}$ $\leq 1+\int_{1}^{T}\frac{1}{t}\mathrm{d}t=1+\log T.$ (23)
If $\eta_{t\in\left[T\right]}=\frac{1}{2L}\land\frac{\eta}{\sqrt{t}}$, we consider the following three cases:
  • $\eta<\frac{1}{2L}$: In this case, we have $\eta_{t\in\left[T\right]}=\frac{\eta}{\sqrt{t}}$ and
    $\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{D_{\psi}(x,x^{1})}{\eta\sum_{t=1}^{T}1/\sqrt{t}}+2\eta(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{1}{t\sum_{s=t}^{T}1/\sqrt{s}}$
[…]
    $\overset{(a)}{\leq}\frac{D_{\psi}(x,x^{1})}{\eta\sqrt{T}}+\frac{4\eta(M^{2}+\sigma^{2})(1+\log T)}{\sqrt{T}}$
    $=\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x,x^{1})}{\eta}+4\eta(M^{2}+\sigma^{2})(1+\log T)\right],$ (24)
    where $(a)$ is by
    $\sum_{t=1}^{T}\frac{1}{t(\sqrt{T+1}-\sqrt{t})}$ $=\sum_{t=1}^{T}\frac{\sqrt{T+1}+\sqrt{t}}{t(T+1-t)}\leq\sum_{t=1}^{T}\frac{2\sqrt{T+1}}{t(T+1-t)}$
[…]
  • $\eta\geq\frac{\sqrt{T}}{2L}$: In this case, we have $\eta_{t\in\left[T\right]}=\frac{1}{2L}$ and
    $\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{D_{\psi}(x,x^{1})}{T/2L}+\frac{M^{2}+\sigma^{2}}{L}\sum_{t=1}^{T}\frac{1}{T-t+1}=\frac{2LD_{\psi}(x,x^{1})}{T}+\frac{M^{2}+\sigma^{2}}{L}\sum_{t=1}^{T}\frac{1}{t}$
[…]
    $\overset{(b)}{\leq}\frac{2LD_{\psi}(x,x^{1})}{T}+\frac{2\eta(M^{2}+\sigma^{2})(1+\log T)}{\sqrt{T}},$ (25)
    where $(b)$ is by $\frac{1}{L}\leq\frac{2\eta}{\sqrt{T}}$.
  • $\eta\in[\frac{1}{2L},\frac{\sqrt{T}}{2L})$: In this case, we define $\tau=\lfloor 4\eta^{2}L^{2}\rfloor$ where $\lfloor\cdot\rfloor$ is the floor function. Note that
    $4\eta^{2}L^{2}\in[1,T)\Rightarrow\tau=\lfloor 4\eta^{2}L^{2}\rfloor\in\left[T-1\right].$
    By observing $\frac{\eta}{\sqrt{t}}\geq\frac{1}{2L}\Leftrightarrow t\in\left[1,\tau\right]$, we can calculate
    $\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\eta_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\sum_{s=t}^{T}\eta_{s}}$
    $\overset{(c)}{\leq}\frac{D_{\psi}(x,x^{1})}{T^{2}}\underbrace{\sum_{t=1}^{T}\frac{1}{\eta_{t}}}_{\text{I}}+2(M^{2}+\sigma^{2})\left(\underbrace{\sum_{t=1}^{\tau}\frac{\eta_{t}^{2}}{\sum_{s=t}^{T}\eta_{s}}}_{\text{II}}+\underbrace{\sum_{t=\tau+1}^{T}\frac{\eta_{t}^{2}}{\sum_{s=t}^{T}\eta_{s}}}_{\text{III}}\right),$
    where $(c)$ is by $T^{2}\leq\left(\sum_{t=1}^{T}\eta_{t}\right)\left(\sum_{t=1}^{T}\frac{1}{\eta_{t}}\right)$. Now we bound terms I, II and III as follows
    $\text{I}=\sum_{t=1}^{T}2L\lor\frac{\sqrt{t}}{\eta}\leq\sum_{t=1}^{T}2L+\frac{\sqrt{t}}{\eta}\leq 2LT+\frac{\sqrt{T}+\int_{1}^{T}\sqrt{t}\mathrm{d}t}{\eta}=2LT+\frac{\sqrt{T}+\frac{2}{3}(T^{\frac{3}{2}}-1)}{\eta}\leq 2LT+\frac{5T^{\frac{3}{2}}}{3\eta}.$
    II $=\sum_{t=1}^{\tau}\frac{\eta_{t}^{2}}{\sum_{s=t}^{\tau}\eta_{s}+\sum_{s=\tau+1}^{T}\eta_{s}}=\sum_{t=1}^{\tau}\frac{1/(4L^{2})}{(\tau-t+1)/2L+\sum_{s=\tau+1}^{T}\eta/\sqrt{s}}$
[…]
    $\leq\begin{cases}\frac{1}{2L}\sum_{t=1}^{\tau}\frac{1}{\tau-t+1}\leq\frac{1}{2L}\left(1+\int_{1}^{\tau}\frac{1}{t}\mathrm{d}t\right)=\frac{1+\log\tau}{2L}\overset{(d)}{\leq}\frac{\eta(1+\log T)}{\sqrt{\tau}}\\ \frac{1}{2L}\sum_{t=1}^{\tau}\frac{1}{4\eta L(\sqrt{T+1}-\sqrt{\tau+1})}=\frac{\tau}{8\eta L^{2}(\sqrt{T+1}-\sqrt{\tau+1})}\overset{(e)}{\leq}\frac{\eta}{2(\sqrt{T+1}-\sqrt{\tau+1})}\end{cases}$
    $\Rightarrow\text{II}$ $\leq\eta(1+\log T)\left(\frac{1}{\sqrt{\tau}}\land\frac{1}{2(\sqrt{T+1}-\sqrt{\tau+1})}\right)\leq\frac{2\eta(1+\log T)}{\sqrt{\tau}+2(\sqrt{T+1}-\sqrt{\tau+1})}\overset{(f)}{\leq}\frac{2\eta(1+\log T)}{\sqrt{T}},$
    where $(d)$ is due to $\tau\leq T$ and $\sqrt{\tau}\leq 2\eta L$, $(e)$ holds by $\tau\leq 4\eta^{2}L^{2}$ and $(f)$ is by, for $\tau\in\left[T-1\right]$ and $T\geq 2$,
    $\sqrt{\tau}+2(\sqrt{T+1}-\sqrt{\tau+1})\geq\sqrt{T-1}+2\sqrt{T+1}-2\sqrt{T}\geq\sqrt{T}.$
[…]
    Thus, we have
    $\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{D_{\psi}(x,x^{1})}{T^{2}}\left(2LT+\frac{5T^{\frac{3}{2}}}{3\eta}\right)+2(M^{2}+\sigma^{2})\left[\frac{2\eta(1+\log T)}{\sqrt{T}}+\frac{2\eta(1+\log T)}{\sqrt{T}}\right]$
    $\leq\frac{2LD_{\psi}(x,x^{1})}{T}+\frac{1}{\sqrt{T}}\left(\frac{5D_{\psi}(x,x^{1})}{3\eta}+8\eta(M^{2}+\sigma^{2})(1+\log T)\right).$ (26)
Combining (24), (25) and (26), we know
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq$ $\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x,x^{1})}{\eta}+4\eta(M^{2}+\sigma^{2})(1+\log T)\right]\lor\left[\frac{2LD_{\psi}(x,x^{1})}{T}+\frac{2\eta(M^{2}+\sigma^{2})(1+\log T)}{\sqrt{T}}\right]$
$\lor\left[\frac{2LD_{\psi}(x,x^{1})}{T}+\frac{1}{\sqrt{T}}\left(\frac{5D_{\psi}(x,x^{1})}{3\eta}+8\eta(M^{2}+\sigma^{2})(1+\log T)\right)\right]$
$\leq$ $\frac{2LD_{\psi}(x,x^{1})}{T}+\frac{1}{\sqrt{T}}\left(\frac{5D_{\psi}(x,x^{1})}{3\eta}+8\eta(M^{2}+\sigma^{2})(1+\log T)\right)$
$=$ $O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x,x^{1})}{\eta}+\eta(M^{2}+\sigma^{2})\log T\right]\right).$ (27)
By plugging in $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x,x^{1})}{M^{2}+\sigma^{2}}}\right)$, we get the desired bound.
If $\eta_{t\in\left[T\right]}=\frac{1}{2L}\land\frac{\eta}{\sqrt{T}}$, we will obtain
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{D_{\psi}(x,x^{1})}{T}\left(2L\lor\frac{\sqrt{T}}{\eta}\right)+2\left(\frac{1}{2L}\land\frac{\eta}{\sqrt{T}}\right)(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{1}{T-t+1}$
$=\frac{D_{\psi}(x,x^{1})}{T}\left(2L\lor\frac{\sqrt{T}}{\eta}\right)+2\left(\frac{1}{2L}\land\frac{\eta}{\sqrt{T}}\right)(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{1}{t}$
[…]
$\leq\frac{2LD_{\psi}(x,x^{1})}{T}+\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x,x^{1})}{\eta}+2\eta(M^{2}+\sigma^{2})(1+\log T)\right]$
$=O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x,x^{1})}{\eta}+\eta(M^{2}+\sigma^{2})\log T\right]\right).$ (28)
By plugging in $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x,x^{1})}{(M^{2}+\sigma^{2})\log T}}\right)$, we get the desired bound. ∎
Theorem C.2.
Under Assumptions 2.-4. and 5B. with $\mu_{f}=\mu_{h}=0$ and let $\delta\in(0,1)$, for any $x\in\mathcal{X}$:
If $T$ is unknown, by taking $\eta_{t\in\left[T\right]}=\frac{1}{2L}\land\frac{\eta}{\sqrt{t}}$, then with probability at least $1-\delta$, there is
$F(x^{T+1})-F(x)\leq O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x,x^{1})}{\eta}+\eta\left(M^{2}+\sigma^{2}\log\frac{1}{\delta}\right)\log T\right]\right).$
In particular, by choosing $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x,x^{1})}{M^{2}+\sigma^{2}\log\frac{1}{\delta}}}\right)$, there is
$F(x^{T+1})-F(x)\leq O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{(M+\sigma\sqrt{\log\frac{1}{\delta}})\sqrt{D_{\psi}(x,x^{1})}\log T}{\sqrt{T}}\right).$
If $T$ is known, by taking $\eta_{t\in\left[T\right]}=\frac{1}{2L}\land\frac{\eta}{\sqrt{T}}$, then with probability at least $1-\delta$, there is
$F(x^{T+1})-F(x)\leq O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x,x^{1})}{\eta}+\eta\left(M^{2}+\sigma^{2}\log\frac{1}{\delta}\right)\log T\right]\right).$
In particular, by choosing $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x,x^{1})}{(M^{2}+\sigma^{2}\log\frac{1}{\delta})\log T}}\right)$, there is
$F(x^{T+1})-F(x)\leq O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{(M+\sigma\sqrt{\log\frac{1}{\delta}})\sqrt{D_{\psi}(x,x^{1})\log T}}{\sqrt{T}}\right).$
Proof.
From Lemma 4.3, if $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}$, with probability at least $1-\delta$, there is
$F(x^{T+1})-F(x)\leq 2\left(1+\max_{2\leq t\leq T}\frac{1}{1-\mu_{f}\eta_{t}}\right)\left[\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+\left(M^{2}+\sigma^{2}\left(1+2\log\frac{2}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}\right],$ (29)
where $\gamma_{t\in\left[T\right]}=\eta_{t}\prod_{s=2}^{t}\frac{1+\mu_{h}\eta_{s-1}}{1-\mu_{f}\eta_{s}}$. Note that $\mu_{f}=\mu_{h}=0$ now, so both $\eta_{t\in\left[T\right]}=\frac{1}{2L}\land\frac{\eta}{\sqrt{t}}$ and $\eta_{t\in\left[T\right]}=\frac{1}{2L}\land\frac{\eta}{\sqrt{T}}$ satisfy $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}=\frac{1}{2L}$. Besides, $\gamma_{t\in\left[T\right]}$ will degenerate to $\eta_{t\in\left[T\right]}$. Then we can simplify (29) into
$F(x^{T+1})-F(x)\leq\frac{4D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\eta_{t}}+4\left(M^{2}+\sigma^{2}\left(1+2\log\frac{2}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\sum_{s=t}^{T}\eta_{s}}.$ (30)
If $\eta_{t\in\left[T\right]}=\frac{1}{2L}\land\frac{\eta}{\sqrt{t}}$, similar to (27), we will have
$F(x^{T+1})-F(x)\leq O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x,x^{1})}{\eta}+\eta\left(M^{2}+\sigma^{2}\log\frac{1}{\delta}\right)\log T\right]\right).$
By plugging in $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x,x^{1})}{M^{2}+\sigma^{2}\log\frac{1}{\delta}}}\right)$, we get the desired bound.
If $\eta_{t\in\left[T\right]}=\frac{1}{2L}\land\frac{\eta}{\sqrt{T}}$, similar to (28), we will get
$F(x^{T+1})-F(x)\leq O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x,x^{1})}{\eta}+\eta\left(M^{2}+\sigma^{2}\log\frac{1}{\delta}\right)\log T\right]\right).$
By plugging in $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x,x^{1})}{(M^{2}+\sigma^{2}\log\frac{1}{\delta})\log T}}\right)$, we get the desired bound. ∎
C.1 Optimal Rates via the Linear Decay Step Size
In this section, we present the full version of Theorems 3.4 and 3.5.
Theorem C.3.
Under Assumptions 2.-4. and 5A. with $\mu_{f}=\mu_{h}=0$, for any $x\in\mathcal{X}$, if $T$ is known, by taking $\eta_{t\in\left[T\right]}=\frac{T-t+1}{2LT}\land\frac{\eta(T-t+1)}{T^{\frac{3}{2}}}$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x,x^{1})}{\eta}+\eta(M^{2}+\sigma^{2})\right]\right).$
In particular, by choosing $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x,x^{1})}{M^{2}+\sigma^{2}}}\right)$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{(M+\sigma)\sqrt{D_{\psi}(x,x^{1})}}{\sqrt{T}}\right).$
Proof.
From Lemma 4.2, if $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq\frac{(1-\mu_{f}\eta_{1})D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}},$ (31)
where $\gamma_{t\in\left[T\right]}=\eta_{t}\prod_{s=2}^{t}\frac{1+\mu_{h}\eta_{s-1}}{1-\mu_{f}\eta_{s}}$. Note that $\mu_{f}=\mu_{h}=0$ now, so $\eta_{t\in\left[T\right]}=\frac{T-t+1}{2LT}\land\frac{\eta(T-t+1)}{T^{\frac{3}{2}}}$ satisfies $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}=\frac{1}{2L}$. Besides, $\gamma_{t\in\left[T\right]}$ will degenerate to $\eta_{t\in\left[T\right]}$. Therefore, (31) can be simplified as follows
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\eta_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\sum_{s=t}^{T}\eta_{s}}.$ (32)
Plugging $\eta_{t\in\left[T\right]}=\frac{T-t+1}{2LT}\land\frac{\eta(T-t+1)}{T^{\frac{3}{2}}}$ into (32), we have
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}T-t+1}\left(2LT\lor\frac{T^{\frac{3}{2}}}{\eta}\right)+\frac{2\eta(M^{2}+\sigma^{2})}{T^{\frac{3}{2}}}\sum_{t=1}^{T}\frac{(T-t+1)^{2}}{\sum_{s=t}^{T}T-s+1}$
$\leq\frac{4LD_{\psi}(x,x^{1})}{T}+\frac{2D_{\psi}(x,x^{1})}{\eta\sqrt{T}}+\frac{4\eta(M^{2}+\sigma^{2})}{\sqrt{T}}$
$=O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x,x^{1})}{\eta}+\eta(M^{2}+\sigma^{2})\right]\right).$ (33)
By plugging in $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x,x^{1})}{M^{2}+\sigma^{2}}}\right)$, we get the desired bound. ∎
Theorem C.4.
Under Assumptions 2.-4. and 5B. with $\mu_{f}=\mu_{h}=0$ and let $\delta\in(0,1)$, for any $x\in\mathcal{X}$, if $T$ is known, by taking $\eta_{t\in\left[T\right]}=\frac{T-t+1}{2LT}\land\frac{\eta(T-t+1)}{T^{\frac{3}{2}}}$, then with probability at least $1-\delta$, there is
$F(x^{T+1})-F(x)\leq O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x,x^{1})}{\eta}+\eta(M^{2}+\sigma^{2}\log\frac{1}{\delta})\right]\right).$
In particular, by choosing $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x,x^{1})}{M^{2}+\sigma^{2}\log\frac{1}{\delta}}}\right)$, there is
$F(x^{T+1})-F(x)\leq O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{(M+\sigma\sqrt{\log\frac{1}{\delta}})\sqrt{D_{\psi}(x,x^{1})}}{\sqrt{T}}\right).$
Proof.
From Lemma 4.3, if $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}$, with probability at least $1-\delta$, there is
$F(x^{T+1})-F(x)\leq 2\left(1+\max_{2\leq t\leq T}\frac{1}{1-\mu_{f}\eta_{t}}\right)\left[\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+\left(M^{2}+\sigma^{2}\left(1+2\log\frac{2}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}\right],$ (34)
where $\gamma_{t\in\left[T\right]}=\eta_{t}\prod_{s=2}^{t}\frac{1+\mu_{h}\eta_{s-1}}{1-\mu_{f}\eta_{s}}$. Note that $\mu_{f}=\mu_{h}=0$ now, hence, $\eta_{t\in\left[T\right]}=\frac{T-t+1}{2LT}\land\frac{\eta(T-t+1)}{T^{\frac{3}{2}}}$ satisfies $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}=\frac{1}{2L}$. Besides, $\gamma_{t\in\left[T\right]}$ will degenerate to $\eta_{t\in\left[T\right]}$. Then we can simplify (34) into
$F(x^{T+1})-F(x)\leq\frac{4D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\eta_{t}}+4\left(M^{2}+\sigma^{2}\left(1+2\log\frac{2}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\sum_{s=t}^{T}\eta_{s}}.$ (35)
Similar to (33), we will have
$F(x^{T+1})-F(x)\leq O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x,x^{1})}{\eta}+\eta\left(M^{2}+\sigma^{2}\log\frac{1}{\delta}\right)\right]\right).$
By plugging in $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x,x^{1})}{M^{2}+\sigma^{2}\log\frac{1}{\delta}}}\right)$, we get the desired bound. ∎
Appendix D Strongly Convex Functions
In this section, we present the full version of the theorems for strongly convex functions (i.e., $\mu_{f}+\mu_{h}>0$) with their proofs.
D.1 The Case of $\mu_{f}>0$
Theorem D.1.
Under Assumptions 2.-4. and 5A. with $\mu_{f}>0$ and $\mu_{h}=0$, let $\kappa_{f}\coloneqq\frac{L}{\mu_{f}}\geq 0$, for any $x\in\mathcal{X}$:
If $T$ is unknown, by taking either $\eta_{t\in\left[T\right]}=\frac{1}{\mu_{f}(t+2\kappa_{f})}$ or $\eta_{t\in\left[T\right]}=\frac{2}{\mu_{f}(t+1+4\kappa_{f})}$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq\begin{cases}O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{f}(T+\kappa_{f})}\right)&\eta_{t\in\left[T\right]}=\frac{1}{\mu_{f}(t+2\kappa_{f})}\\ O\left(\frac{L(1+\kappa_{f})D_{\psi}(x,x^{1})}{T(T+\kappa_{f})}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{f}(T+\kappa_{f})}\right)&\eta_{t\in\left[T\right]}=\frac{2}{\mu_{f}(t+1+4\kappa_{f})}\end{cases}.$
If $T$ is known, by taking $\eta_{t\in\left[T\right]}=\begin{cases}\frac{1}{\mu_{f}(1+2\kappa_{f})}&t=1\\ \frac{1}{\mu_{f}(\eta+2\kappa_{f})}&2\leq t\leq\tau\\ \frac{2}{\mu_{f}(t-\tau+2+4\kappa_{f})}&t\geq\tau+1\end{cases}$ where $\eta\geq 0$ can be any number satisfying $\eta+\kappa_{f}>1$ and $\tau\coloneqq\left\lceil\frac{T}{2}\right\rceil$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{LD_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2\eta+4\kappa_{f}}\right)}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{f}(T+\kappa_{f})}\right).$
Proof.
When $\mu_{f}>0$ and $\mu_{h}=0$, suppose the condition of $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}$ in Lemma 4.2 holds, we have
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq\frac{(1-\mu_{f}\eta_{1})D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}.$ (36)
Now observe that
$\gamma_{t\in\left[T\right]}=\eta_{t}\prod_{s=2}^{t}\frac{1+\mu_{h}\eta_{s-1}}{1-\mu_{f}\eta_{s}}=\eta_{t}\prod_{s=2}^{t}\frac{1}{1-\mu_{f}\eta_{s}}=\eta_{t}\Gamma_{t}=\begin{cases}\frac{\Gamma_{t}-\Gamma_{t-1}}{\mu_{f}}&t\geq 2\\ \eta_{1}&t=1\end{cases},$
where $\Gamma_{t\in\left[T\right]}\coloneqq\prod_{s=2}^{t}\frac{1}{1-\mu_{f}\eta_{s}}$. Hence, (36) can be rewritten as
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq\frac{(1-\mu_{f}\eta_{1})D_{\psi}(x,x^{1})}{\eta_{1}+\frac{\Gamma_{T}-1}{\mu_{f}}}+2(M^{2}+\sigma^{2})\left[\frac{\eta_{1}^{2}}{\eta_{1}+\frac{\Gamma_{T}-1}{\mu_{f}}}+\sum_{t=2}^{T}\frac{\mu_{f}\eta_{t}^{2}}{(\Gamma_{T}/\Gamma_{t-1}-1)(1-\mu_{f}\eta_{t})}\right].$ (37)
Now let us check the condition of $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}$ for our three choices respectively:
$\eta_{t\in\left[T\right]}$ $=\frac{1}{\mu_{f}(t+2\kappa_{f})}\leq\frac{1}{\mu_{f}+2L}\leq\frac{1}{2L\lor\mu_{f}};$
$\eta_{t\in\left[T\right]}$ $=\frac{2}{\mu_{f}(t+1+4\kappa_{f})}\leq\frac{1}{\mu_{f}+2L}\leq\frac{1}{2L\lor\mu_{f}};$
$\eta_{t\in\left[T\right]}$ $=\begin{cases}\frac{1}{\mu_{f}(1+2\kappa_{f})}=\frac{1}{\mu_{f}+2L}\leq\frac{1}{2L\lor\mu_{f}}&t=1\\ \frac{1}{\mu_{f}(\eta+2\kappa_{f})}\leq\frac{1}{\mu_{f}(2\kappa_{f}\lor(\eta+\kappa_{f}))}\leq\frac{1}{2L\lor\mu_{f}}&2\leq t\leq\tau\\ \frac{2}{\mu_{f}(t-\tau+2+4\kappa_{f})}\leq\frac{1}{\mu_{f}+2L}\leq\frac{1}{2L\lor\mu_{f}}&t\geq\tau+1\end{cases}.$
Therefore, (37) holds for all cases.
First, we consider $\eta_{t\in\left[T\right]}=\frac{1}{\mu_{f}(t+2\kappa_{f})}$. We can calculate $\eta_{1}=\frac{1}{\mu_{f}(1+2\kappa_{f})}$ and
$\Gamma_{t\in\left[T\right]}=\prod_{s=2}^{t}\frac{1}{1-\mu_{f}\eta_{s}}=\prod_{s=2}^{t}\frac{s+2\kappa_{f}}{s-1+2\kappa_{f}}=\frac{t+2\kappa_{f}}{1+2\kappa_{f}}.$
Hence, using (37), we have
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$
$\leq$ $\frac{(1-\frac{1}{1+2\kappa_{f}})D_{\psi}(x,x^{1})}{\frac{1}{\mu_{f}(1+2\kappa_{f})}+\frac{\Gamma_{T}-1}{\mu_{f}}}+2(M^{2}+\sigma^{2})\left[\frac{\frac{1}{\mu_{f}^{2}(1+2\kappa_{f})^{2}}}{\frac{1}{\mu_{f}(1+2\kappa_{f})}+\frac{\Gamma_{T}-1}{\mu_{f}}}+\sum_{t=2}^{T}\frac{\mu_{f}\cdot\frac{1}{\mu_{f}^{2}(t+2\kappa_{f})^{2}}}{(\Gamma_{T}/\Gamma_{t-1}-1)(1-\mu_{f}\cdot\frac{1}{\mu_{f}(t+2\kappa_{f})})}\right]$
$=$ $\frac{2LD_{\psi}(x,x^{1})}{(1+2\kappa_{f})\Gamma_{T}-2\kappa_{f}}+\frac{2(M^{2}+\sigma^{2})}{\mu_{f}}\left[\frac{1}{(1+2\kappa_{f})((1+2\kappa_{f})\Gamma_{T}-2\kappa_{f})}+\sum_{t=2}^{T}\frac{1}{(\Gamma_{T}/\Gamma_{t-1}-1)(t-1+2\kappa_{f})(t+2\kappa_{f})}\right]$
$=$ $\frac{2LD_{\psi}(x,x^{1})}{T}+\frac{2(M^{2}+\sigma^{2})}{\mu_{f}}\left[\frac{1}{(1+2\kappa_{f})T}+\sum_{t=2}^{T}\frac{1}{(T-t+1)(t+2\kappa_{f})}\right]$
$=$ $\frac{2LD_{\psi}(x,x^{1})}{T}+\frac{2(M^{2}+\sigma^{2})}{\mu_{f}}\sum_{t=1}^{T}\frac{1}{(T-t+1)(t+2\kappa_{f})}$
$=$ $\frac{2LD_{\psi}(x,x^{1})}{T}+\frac{2(M^{2}+\sigma^{2})}{\mu_{f}}\sum_{t=1}^{T}\frac{1}{T+1+2\kappa_{f}}\left(\frac{1}{T-t+1}+\frac{1}{t+2\kappa_{f}}\right)$
$\leq$ $\frac{2LD_{\psi}(x,x^{1})}{T}+\frac{2(M^{2}+\sigma^{2})}{\mu_{f}}\cdot\frac{1+\log T+\frac{1}{1+2\kappa_{f}}+\log\frac{T+2\kappa_{f}}{1+2\kappa_{f}}}{T+1+2\kappa_{f}}$
$\leq$ $\frac{2LD_{\psi}(x,x^{1})}{T}+\frac{4(M^{2}+\sigma^{2})(1+\log T)}{\mu_{f}(T+2\kappa_{f})}=O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{f}(T+\kappa_{f})}\right).$ (38)
Next, for the case of $\eta_{t\in\left[T\right]}=\frac{2}{\mu_{f}(t+1+4\kappa_{f})}$, there are $\eta_{1}=\frac{1}{\mu_{f}(1+2\kappa_{f})}$ and
$\Gamma_{t\in\left[T\right]}=\prod_{s=2}^{t}\frac{1}{1-\mu_{f}\eta_{s}}=\prod_{s=2}^{t}\frac{s+1+4\kappa_{f}}{s-1+4\kappa_{f}}=\frac{(t+4\kappa_{f})(t+1+4\kappa_{f})}{(1+4\kappa_{f})(2+4\kappa_{f})}.$
Thus, we have
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$
$\leq$ $\frac{(1-\frac{1}{1+2\kappa_{f}})D_{\psi}(x,x^{1})}{\frac{1}{\mu_{f}(1+2\kappa_{f})}+\frac{\Gamma_{T}-1}{\mu_{f}}}+2(M^{2}+\sigma^{2})\left[\frac{\frac{1}{\mu_{f}^{2}(1+2\kappa_{f})^{2}}}{\frac{1}{\mu_{f}(1+2\kappa_{f})}+\frac{\Gamma_{T}-1}{\mu_{f}}}+\sum_{t=2}^{T}\frac{\mu_{f}\cdot\frac{4}{\mu_{f}^{2}(t+1+4\kappa_{f})^{2}}}{(\Gamma_{T}/\Gamma_{t-1}-1)(1-\mu_{f}\cdot\frac{2}{\mu_{f}(t+1+4\kappa_{f})})}\right]$
$=$ $\frac{2LD_{\psi}(x,x^{1})}{(1+2\kappa_{f})\Gamma_{T}-2\kappa_{f}}+\frac{2(M^{2}+\sigma^{2})}{\mu_{f}}\left[\frac{1}{(1+2\kappa_{f})((1+2\kappa_{f})\Gamma_{T}-2\kappa_{f})}\right$.
$\qquad\qquad\qquad\qquad\qquad\qquad\left.+\sum_{t=2}^{T}\frac{4}{(\Gamma_{T}/\Gamma_{t-1}-1)(t-1+4\kappa_{f})(t+1+4\kappa_{f})}\right]$
$=$ $\frac{4(1+4\kappa_{f})LD_{\psi}(x,x^{1})}{T(T+1+8\kappa_{f})}$
$+\frac{2(M^{2}+\sigma^{2})}{\mu_{f}}\left[\frac{2(1+4\kappa_{f})}{(1+2\kappa_{f})T(T+1+8\kappa_{f})}+\sum_{t=2}^{T}\frac{4(t+4\kappa_{f})}{(T+1-t)(T+t+8\kappa_{f})(t+1+4\kappa_{f})}\right]$
$=$ $\frac{4(1+4\kappa_{f})LD_{\psi}(x,x^{1})}{T(T+1+8\kappa_{f})}+\frac{2(M^{2}+\sigma^{2})}{\mu_{f}}\sum_{t=1}^{T}\frac{4(t+4\kappa_{f})}{(T+1-t)(T+t+8\kappa_{f})(t+1+4\kappa_{f})}$
$\leq$ $\frac{4(1+4\kappa_{f})LD_{\psi}(x,x^{1})}{T(T+1+8\kappa_{f})}+\frac{2(M^{2}+\sigma^{2})}{\mu_{f}}\sum_{t=1}^{T}\frac{4}{2T+1+8\kappa_{f}}\left(\frac{1}{T+1-t}+\frac{1}{T+t+8\kappa_{f}}\right)$
$\leq$ $\frac{4(1+4\kappa_{f})LD_{\psi}(x,x^{1})}{T(T+1+8\kappa_{f})}+\frac{2(M^{2}+\sigma^{2})}{\mu_{f}}\cdot\frac{4(1+\log T+\log\frac{2T+8\kappa_{f}}{T+8\kappa_{f}})}{2T+1+8\kappa_{f}}$
$\leq$ $\frac{4(1+4\kappa_{f})LD_{\psi}(x,x^{1})}{T(T+1+8\kappa_{f})}+\frac{8(M^{2}+\sigma^{2})(1+\log 2T)}{\mu_{f}(2T+1+8\kappa_{f})}=O\left(\frac{L(1+\kappa_{f})D_{\psi}(x,x^{1})}{T(T+\kappa_{f})}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{f}(T+\kappa_{f})}\right).$ (39)
Finally, if $T$ is known, recall that we choose
$\eta_{t\in\left[T\right]}=\begin{cases}\frac{1}{\mu_{f}(1+2\kappa_{f})}&t=1\\ \frac{1}{\mu_{f}(\eta+2\kappa_{f})}&2\leq t\leq\tau\\ \frac{2}{\mu_{f}(t-\tau+2+4\kappa_{f})}&t\geq\tau+1\end{cases}.$
Note that we have $\eta_{1}=\frac{1}{\mu_{f}(1+2\kappa_{f})}$ and
$\Gamma_{t\in\left[T\right]}=\prod_{s=2}^{t}\frac{1}{1-\mu_{f}\eta_{s}}=\begin{cases}\left(\frac{\eta+2\kappa_{f}}{\eta+2\kappa_{f}-1}\right)^{t-1}&t\leq\tau\\ \left(\frac{\eta+2\kappa_{f}}{\eta+2\kappa_{f}-1}\right)^{\tau-1}\frac{(t-\tau+1+4\kappa_{f})(t-\tau+2+4\kappa_{f})}{(1+4\kappa_{f})(2+4\kappa_{f})}&t\geq\tau+1\end{cases}.$
So we know
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq$ $\frac{(1-\mu_{f}\eta_{1})D_{\psi}(x,x^{1})}{\eta_{1}+\frac{\Gamma_{T}-1}{\mu_{f}}}+2(M^{2}+\sigma^{2})\left[\frac{\eta_{1}^{2}}{\eta_{1}+\frac{\Gamma_{T}-1}{\mu_{f}}}+\sum_{t=2}^{T}\frac{\mu_{f}\eta_{t}^{2}}{(\Gamma_{T}/\Gamma_{t-1}-1)(1-\mu_{f}\eta_{t})}\right]$
$=$ $\frac{2LD_{\psi}(x,x^{1})}{(1+2\kappa_{f})\Gamma_{T}-2\kappa_{f}}+\frac{2(M^{2}+\sigma^{2})}{\mu_{f}}\cdot\frac{1}{(1+2\kappa_{f})((1+2\kappa_{f})\Gamma_{T}-2\kappa_{f})}$
$+2\mu_{f}(M^{2}+\sigma^{2})\left[\underbrace{\sum_{t=2}^{\tau}\frac{\eta_{t}^{2}}{(\Gamma_{T}/\Gamma_{t-1}-1)(1-\mu_{f}\eta_{t})}}_{\text{I}}+\underbrace{\sum_{t=\tau+1}^{T}\frac{\eta_{t}^{2}}{(\Gamma_{T}/\Gamma_{t-1}-1)(1-\mu_{f}\eta_{t})}}_{\text{II}}\right].$ (40)
Note that we can bound
$\frac{1}{(1+2\kappa_{f})\Gamma_{T}-2\kappa_{f}}=$ $\frac{1}{\left(\frac{\eta+2\kappa_{f}}{\eta+2\kappa_{f}-1}\right)^{\tau-1}\frac{(T-\tau+1+4\kappa_{f})(T-\tau+2+4\kappa_{f})}{2(1+4\kappa_{f})}-2\kappa_{f}}$
$\overset{(a)}{\leq}$ $\frac{1}{\left(\frac{\eta+2\kappa_{f}}{\eta+2\kappa_{f}-1}\right)^{\tau-1}(1+2\kappa_{f})-2\kappa_{f}}=\frac{1}{2\kappa_{f}\left[\left(\frac{\eta+2\kappa_{f}}{\eta+2\kappa_{f}-1}\right)^{\tau-1}-1\right]+\left(\frac{\eta+2\kappa_{f}}{\eta+2\kappa_{f}-1}\right)^{\tau-1}}$
$\overset{(b)}{\leq}$ $\left(1-\frac{1}{\eta+2\kappa_{f}}\right)^{\tau-1}\leq\exp\left(-\frac{\tau-1}{\eta+2\kappa_{f}}\right)=\exp\left(\frac{-\tau}{\eta+2\kappa_{f}}+\frac{1}{\eta+2\kappa_{f}}\right)$
$\overset{(c)}{\leq}$ $\exp\left(-\frac{T}{2(\eta+2\kappa_{f})}+1\right),$ (41)
where $(a)$ holds due to $T-\tau\geq 0$, $(b)$ is by $\kappa_{f}\geq 0$ and $\left(\frac{\eta+2\kappa_{f}}{\eta+2\kappa_{f}-1}\right)^{\tau-1}\geq 1$, $(c)$ is from $\tau\geq\frac{T}{2}$, $\eta+\kappa_{f}>1$ and $\kappa_{f}\geq 0$. We can also bound
$\frac{1}{(1+2\kappa_{f})((1+2\kappa_{f})\Gamma_{T}-2\kappa_{f})}$
$=$ $\frac{1}{(1+2\kappa_{f})\left[\left(\frac{\eta+2\kappa_{f}}{\eta+2\kappa_{f}-1}\right)^{\tau-1}\frac{(T-\tau+1+4\kappa_{f})(T-\tau+2+4\kappa_{f})}{2(1+4\kappa_{f})}-2\kappa_{f}\right]}$
$\leq$ $\frac{1}{(1+2\kappa_{f})\left[\frac{(T-\tau+1+4\kappa_{f})(T-\tau+2+4\kappa_{f})}{2(1+4\kappa_{f})}-2\kappa_{f}\right]}=\frac{2(1+4\kappa_{f})}{(1+2\kappa_{f})(T-\tau+1)(T-\tau+2+8\kappa_{f})}$
$\overset{(d)}{\leq}$ $\frac{2(1+4\kappa_{f})}{(1+2\kappa_{f})(T-\frac{T+1}{2}+1)(T-\frac{T+1}{2}+2+8\kappa_{f})}=\frac{8(1+4\kappa_{f})}{(1+2\kappa_{f})(T+1)(T+3+16\kappa_{f})}$
$\overset{(e)}{\leq}$ $\frac{8}{T+3+16\kappa_{f}},$ (42)
where $(d)$ is by $\tau\leq\frac{T+1}{2}$ and $(e)$ is by $T\geq 1$. Besides, there is
$\text{I}=$ $\frac{1}{\mu_{f}^{2}(\eta+2\kappa_{f})(\eta+2\kappa_{f}-1)}\sum_{t=2}^{\tau}\frac{1}{\left(\frac{\eta+2\kappa_{f}}{\eta+2\kappa_{f}-1}\right)^{\tau-t+1}\frac{(T-\tau+1+4\kappa_{f})(T-\tau+2+4\kappa_{f})}{(1+4\kappa_{f})(2+4\kappa_{f})}-1}$
$\overset{(f)}{\leq}$ $\frac{1}{\mu_{f}^{2}(\eta+2\kappa_{f})(\eta+2\kappa_{f}-1)\frac{\eta+2\kappa_{f}}{\eta+2\kappa_{f}-1}}\sum_{t=2}^{\tau}\frac{1}{\frac{(T-\tau+1+4\kappa_{f})(T-\tau+2+4\kappa_{f})}{(1+4\kappa_{f})(2+4\kappa_{f})}-1}$
$=$ $\frac{1}{\mu_{f}^{2}(\eta+2\kappa_{f})^{2}}\sum_{t=2}^{\tau}\frac{(1+4\kappa_{f})(2+4\kappa_{f})}{(T-\tau)(T-\tau+3+8\kappa_{f})}=\frac{(1+4\kappa_{f})(2+4\kappa_{f})}{\mu_{f}^{2}(\eta+2\kappa_{f})^{2}}\cdot\frac{\tau-1}{(T-\tau)(T-\tau+3+8\kappa_{f})}$
$\overset{(g)}{\leq}$ $\frac{(1+4\kappa_{f})(2+4\kappa_{f})}{\mu_{f}^{2}(\eta+2\kappa_{f})^{2}}\cdot\frac{\frac{T+1}{2}-1}{(T-\frac{T+1}{2})(T-\frac{T+1}{2}+3+8\kappa_{f})}$
$=$ $\frac{(1+4\kappa_{f})(2+4\kappa_{f})}{\mu_{f}^{2}(\eta+2\kappa_{f})^{2}}\cdot\frac{2}{T+5+16\kappa_{f}}\overset{(h)}{\leq}\frac{32}{\mu_{f}^{2}(T+5+16\kappa_{f})},$ (43)
where $(f)$ is by $\tau-t\geq 0$ (w.l.o.g., we can assume $\tau\geq 2$, otherwise, there is $\text{I}=0\leq\frac{8}{\mu_{f}^{2}(T+5+16\kappa_{f})}$) and $\frac{\eta+2\kappa_{f}}{\eta+2\kappa_{f}-1}\geq 1$, $(g)$ is due to $\tau\leq\frac{T+1}{2}$ and $(h)$ is by $\eta+\kappa_{f}>1$. We also have
$\text{II}=$ $\frac{4}{\mu_{f}^{2}}\sum_{t=\tau+1}^{T}\frac{1}{\left[\frac{(T-\tau+1+4\kappa_{f})(T-\tau+2+4\kappa_{f})}{(t-\tau+4\kappa_{f})(t-\tau+1+4\kappa_{f})}-1\right](t-\tau+2+4\kappa_{f})(t-\tau+4\kappa_{f})}$
$=$ $\frac{4}{\mu_{f}^{2}}\sum_{t=1}^{T-\tau}\frac{1}{\left[\frac{(T-\tau+1+4\kappa_{f})(T-\tau+2+4\kappa_{f})}{(t+4\kappa_{f})(t+1+4\kappa_{f})}-1\right](t+2+4\kappa_{f})(t+4\kappa_{f})}$
$=$ $\frac{4}{\mu_{f}^{2}}\sum_{t=1}^{T-\tau}\frac{t+1+4\kappa_{f}}{\left[(T-\tau+1+4\kappa_{f})(T-\tau+2+4\kappa_{f})-(t+4\kappa_{f})(t+1+4\kappa_{f})\right](t+2+4\kappa_{f})}$
$\leq$ $\frac{4}{\mu_{f}^{2}}\sum_{t=1}^{T-\tau}\frac{1}{(T-\tau+1+4\kappa_{f})(T-\tau+2+4\kappa_{f})-(t+4\kappa_{f})(t+1+4\kappa_{f})}$
$=$ $\frac{4}{\mu_{f}^{2}}\sum_{t=1}^{T-\tau}\frac{1}{(T-\tau+1-t)(T-\tau+2+8\kappa_{f}+t)}$
$=$ $\frac{4}{\mu_{f}^{2}}\sum_{t=1}^{T-\tau}\frac{1}{2T-2\tau+3+8\kappa_{f}}\left(\frac{1}{T-\tau+1-t}+\frac{1}{T-\tau+2+8\kappa_{f}+t}\right)$
$\leq$ $\frac{4(1+\log(T-\tau)+\log\frac{2T-2\tau+2+8\kappa_{f}}{T-\tau+2+8\kappa_{f}})}{\mu_{f}^{2}(2T-2\tau+3+8\kappa_{f})}\leq\frac{4(1+\log T)}{\mu_{f}^{2}(T+2+8\kappa_{f})},$ (44)
where we use $\frac{T}{2}\leq\tau\leq\frac{T+1}{2}$ in the last inequality.
Plugging (41), (42), (43) and (44) into (40), we have
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq$ $2LD_{\psi}(x,x^{1})\exp\left(-\frac{T}{2(\eta+2\kappa_{f})}+1\right)+\frac{2(M^{2}+\sigma^{2})}{\mu_{f}}\cdot\frac{8}{T+3+16\kappa_{f}}$
$+2\mu_{f}(M^{2}+\sigma^{2})\left[\frac{32}{\mu_{f}^{2}(T+5+16\kappa_{f})}+\frac{4(1+\log T)}{\mu_{f}^{2}(T+2+8\kappa_{f})}\right]$
$=$ $O\left(\frac{LD_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2\eta+4\kappa_{f}}\right)}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{f}(T+\kappa_{f})}\right).$ (45)
Theorem D.2.
Under Assumptions 2.-4. and 5B. with $\mu_{f}>0$ and $\mu_{h}=0$, let $\kappa_{f}\coloneqq\frac{L}{\mu_{f}}\geq 0$ and $\delta\in(0,1)$, for any $x\in\mathcal{X}$:
If $T$ is unknown, by taking either $\eta_{t\in\left[T\right]}=\frac{1}{\mu_{f}(t+2\kappa_{f})}$ or $\eta_{t\in\left[T\right]}=\frac{2}{\mu_{f}(t+1+4\kappa_{f})}$, then with probability at least $1-\delta$, there is
$F(x^{T+1})-F(x)\leq\begin{cases}O\left(\frac{\mu_{f}(1+\kappa_{f})D_{\psi}(x,x^{1})}{T}+\frac{(M^{2}+\sigma^{2}\log\frac{1}{\delta})\log T}{\mu_{f}(T+\kappa_{f})}\right)&\eta_{t\in\left[T\right]}=\frac{1}{\mu_{f}(t+2\kappa_{f})}\\ O\left(\frac{\mu_{f}(1+\kappa_{f})^{2}D_{\psi}(x,x^{1})}{T(T+\kappa_{f})}+\frac{(M^{2}+\sigma^{2}\log\frac{1}{\delta})\log T}{\mu_{f}(T+\kappa_{f})}\right)&\eta_{t\in\left[T\right]}=\frac{2}{\mu_{f}(t+1+4\kappa_{f})}\end{cases}.$
If $T$ is known, by taking $\eta_{t\in\left[T\right]}=\begin{cases}\frac{1}{\mu_{f}(1+2\kappa_{f})}&t=1\\ \frac{1}{\mu_{f}(\eta+2\kappa_{f})}&2\leq t\leq\tau\\ \frac{2}{\mu_{f}(t-\tau+2+4\kappa_{f})}&t\geq\tau+1\end{cases}$ where $\eta\geq 0$ can be any number satisfying $\eta+\kappa_{f}>1$ and $\tau\coloneqq\left\lceil\frac{T}{2}\right\rceil$, then with probability at least $1-\delta$, there is
$F(x^{T+1})-F(x)\leq O\left(\left(1\lor\frac{1}{\eta+2\kappa_{f}-1}\right)\left[\frac{\mu_{f}(1+\kappa_{f})D_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2\eta+4\kappa_{f}}\right)}+\frac{(M^{2}+\sigma^{2}\log\frac{1}{\delta})\log T}{\mu_{f}(T+\kappa_{f})}\right]\right).$
Proof.
When $\mu_{f}>0$ and $\mu_{h}=0$, suppose the condition of $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}$ in Lemma 4.3 holds, we will have with probability at least $1-\delta$
$F(x^{T+1})-F(x)\leq 2\left(1+\max_{2\leq t\leq T}\frac{1}{1-\mu_{f}\eta_{t}}\right)\left[\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+\left(M^{2}+\sigma^{2}\left(1+2\log\frac{2}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}\right].$ (46)
Now observe that
$\gamma_{t\in\left[T\right]}=\eta_{t}\prod_{s=2}^{t}\frac{1+\mu_{h}\eta_{s-1}}{1-\mu_{f}\eta_{s}}=\eta_{t}\prod_{s=2}^{t}\frac{1}{1-\mu_{f}\eta_{s}}=\eta_{t}\Gamma_{t}=\begin{cases}\frac{\Gamma_{t}-\Gamma_{t-1}}{\mu_{f}}&t\geq 2\\ \eta_{1}&t=1\end{cases},$
where $\Gamma_{t\in\left[T\right]}\coloneqq\prod_{s=2}^{t}\frac{1}{1-\mu_{f}\eta_{s}}$. Hence, (46) can be rewritten as
$F(x^{T+1})-F(x)\leq 2\left(1+\max_{2\leq t\leq T}\frac{1}{1-\mu_{f}\eta_{t}}\right)\left(\frac{D_{\psi}(x,x^{1})}{\eta_{1}+\frac{\Gamma_{T}-1}{\mu_{f}}}+\left(M^{2}+\sigma^{2}\left(1+2\log\frac{2}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}\right).$ (47)
Now let us check the condition of $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}$ for our three choices respectively:
$\eta_{t\in\left[T\right]}=$ $\frac{1}{\mu_{f}(t+2\kappa_{f})}\leq\frac{1}{\mu_{f}+2L}\leq\frac{1}{2L\lor\mu_{f}};$
$\eta_{t\in\left[T\right]}=$ $\frac{2}{\mu_{f}(t+1+4\kappa_{f})}\leq\frac{1}{\mu_{f}+2L}\leq\frac{1}{2L\lor\mu_{f}};$
$\eta_{t\in\left[T\right]}=$ $\begin{cases}\frac{1}{\mu_{f}(1+2\kappa_{f})}=\frac{1}{\mu_{f}+2L}\leq\frac{1}{2L\lor\mu_{f}}&t=1\\ \frac{1}{\mu_{f}(\eta+2\kappa_{f})}\leq\frac{1}{\mu_{f}(2\kappa_{f}\lor(\eta+\kappa_{f}))}\leq\frac{1}{2L\lor\mu_{f}}&2\leq t\leq\tau\\ \frac{2}{\mu_{f}(t-\tau+2+4\kappa_{f})}\leq\frac{1}{\mu_{f}+2L}\leq\frac{1}{2L\lor\mu_{f}}&t\geq\tau+1\end{cases}.$
Therefore, (47) holds for all cases.
First, we consider $\eta_{t\in\left[T\right]}=\frac{1}{\mu_{f}(t+2\kappa_{f})}$. We can find $1+\max_{2\leq t\leq T}\frac{1}{1-\mu_{f}\eta_{t}}=1+\frac{1}{1-\mu_{f}\eta_{2}}\leq 3$, $\eta_{1}=\frac{1}{\mu_{f}(1+2\kappa_{f})}$ and
$\Gamma_{t\in\left[T\right]}=\prod_{s=2}^{t}\frac{1}{1-\mu_{f}\eta_{s}}=\prod_{s=2}^{t}\frac{s+2\kappa_{f}}{s-1+2\kappa_{f}}=\frac{t+2\kappa_{f}}{1+2\kappa_{f}}.$
Hence, using (47), we have
$F(x^{T+1})-F(x)\leq 6\left(\frac{D_{\psi}(x,x^{1})}{\eta_{1}+\frac{\Gamma_{T}-1}{\mu_{f}}}+\left(M^{2}+\sigma^{2}\left(1+2\log\frac{2}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}\right).$
Following similar steps in the proof of (38), we can get
$F(x^{T+1})-F(x)\leq O\left(\frac{\mu_{f}(1+\kappa_{f})D_{\psi}(x,x^{1})}{T}+\frac{(M^{2}+\sigma^{2}\log\frac{1}{\delta})\log T}{\mu_{f}(T+\kappa_{f})}\right).$
Next, for the case of $\eta_{t\in\left[T\right]}=\frac{2}{\mu_{f}(t+1+4\kappa_{f})}$, there are $1+\max_{2\leq t\leq T}\frac{1}{1-\mu_{f}\eta_{t}}=1+\frac{1}{1-\mu_{f}\eta_{2}}\leq 4$, $\eta_{1}=\frac{1}{\mu_{f}(1+2\kappa_{f})}$ and
$\Gamma_{t\in\left[T\right]}=\prod_{s=2}^{t}\frac{1}{1-\mu_{f}\eta_{s}}=\prod_{s=2}^{t}\frac{s+1+4\kappa_{f}}{s-1+4\kappa_{f}}=\frac{(t+4\kappa_{f})(t+1+4\kappa_{f})}{2(1+4\kappa_{f})(1+2\kappa_{f})}.$
Thus, we have
$F(x^{T+1})-F(x)\leq 8\left(\frac{D_{\psi}(x,x^{1})}{\eta_{1}+\frac{\Gamma_{T}-1}{\mu_{f}}}+\left(M^{2}+\sigma^{2}\left(1+2\log\frac{2}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}\right).$
Following similar steps in the proof of (39), we can get
$F(x^{T+1})-F(x)\leq O\left(\frac{\mu_{f}(1+\kappa_{f})^{2}D_{\psi}(x,x^{1})}{T(T+\kappa_{f})}+\frac{(M^{2}+\sigma^{2}\log\frac{1}{\delta})\log T}{\mu_{f}(T+\kappa_{f})}\right).$
Finally, if $T$ is known, we recall the current choice is
$\eta_{t\in\left[T\right]}=\begin{cases}\frac{1}{\mu_{f}(1+2\kappa_{f})}&t=1\\ \frac{1}{\mu_{f}(\eta+2\kappa_{f})}&2\leq t\leq\tau\\ \frac{2}{\mu_{f}(t-\tau+2+4\kappa_{f})}&t\geq\tau+1\end{cases},$
where $\eta>1$ and $\tau=\left\lceil\frac{T}{2}\right\rceil$. Note that we have
$1+\max_{2\leq t\leq T}\frac{1}{1-\mu_{f}\eta_{t}}$ $=1+\frac{1}{1-\mu_{f}(\eta_{2}\lor\eta_{\tau+1})}=2+\frac{1}{\eta\land 1.5+2\kappa_{f}-1}$
$=2+\frac{1}{\eta+2\kappa_{f}-1}\lor\frac{1}{0.5+2\kappa_{f}}\leq 4+\frac{1}{\eta+2\kappa_{f}-1},$
$\eta_{1}=\frac{1}{\mu_{f}(1+2\kappa_{f})}$ and
$\Gamma_{t\in\left[T\right]}=\prod_{s=2}^{t}\frac{1}{1-\mu_{f}\eta_{s}}=\begin{cases}\left(\frac{\eta+2\kappa_{f}}{\eta+2\kappa_{f}-1}\right)^{t-1}&t\leq\tau\\ \left(\frac{\eta+2\kappa_{f}}{\eta+2\kappa_{f}-1}\right)^{\tau-1}\frac{(t-\tau+1+4\kappa_{f})(t-\tau+2+4\kappa_{f})}{(1+4\kappa_{f})(2+4\kappa_{f})}&t\geq\tau+1\end{cases}.$
Thus, there is
$F(x^{T+1})-F(x)\leq\left(8+\frac{2}{\eta+2\kappa_{f}-1}\right)\left(\frac{D_{\psi}(x,x^{1})}{\eta_{1}+\frac{\Gamma_{T}-1}{\mu_{f}}}+\left(M^{2}+\sigma^{2}\left(1+2\log\frac{2}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}\right).$
Following similar steps in the proof of (45), we can obtain
$F(x^{T+1})-F(x)\leq O\left(\left(1\lor\frac{1}{\eta+2\kappa_{f}-1}\right)\left[\frac{\mu_{f}(1+\kappa_{f})D_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2\eta+4\kappa_{f}}\right)}+\frac{(M^{2}+\sigma^{2}\log\frac{1}{\delta})\log T}{\mu_{f}(T+\kappa_{f})}\right]\right).$
D.2 The Case of $\mu_{h}>0$
Theorem D.3.
Under Assumptions 2.-4. and 5A. with $\mu_{f}=0$ and $\mu_{h}>0$, let $\kappa_{h}\coloneqq\frac{L}{\mu_{h}}\geq 0$, for any $x\in\mathcal{X}$:
If $T$ is unknown, by taking $\eta_{t\in\left[T\right]}=\frac{2}{\mu_{h}(t+4\kappa_{h})}$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{\mu_{h}(1+\kappa_{h})^{2}D_{\psi}(x,x^{1})}{T(T+\kappa_{h})}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{h}(T+\kappa_{h})}\right).$
If $T$ is known, by taking $\eta_{t\in\left[T\right]}=\begin{cases}\frac{1}{\mu_{h}(\eta+2\kappa_{h})}&t\leq\tau\\ \frac{2}{\mu_{h}(t-\tau+4\kappa_{h})}&t\geq\tau+1\end{cases}$ where $\eta\geq 0$ can be any number satisfying $\eta+\kappa_{h}>0$ and $\tau\coloneqq\left\lceil\frac{T}{2}\right\rceil$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{\mu_{h}D_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2(1+\eta+2\kappa_{h})}\right)-1}+\left(1\lor\frac{1}{\eta+2\kappa_{h}}\right)\frac{(M^{2}+\sigma^{2})\log T}{\mu_{h}(T+\kappa_{h})}\right).$
Proof.
When $\mu_{f}=0$ and $\mu_{h}>0$, suppose the condition of $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}=\frac{1}{2L}$ in Lemma 4.2 holds, we will have
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{(1-\mu_{f}\eta_{1})D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}$
$=\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}.$ (48)
Observe that
$\gamma_{t\in\left[T\right]}=\eta_{t}\prod_{s=2}^{t}\frac{1+\mu_{h}\eta_{s-1}}{1-\mu_{f}\eta_{s}}=\eta_{t}\prod_{s=2}^{t}(1+\mu_{h}\eta_{s-1})=\eta_{t}\Gamma_{t-1}=\frac{\Gamma_{t}-\Gamma_{t-1}}{\mu_{h}},$
where $\Gamma_{t\in\left\{0\right\}\cup\left[T\right]}\coloneqq\prod_{s=1}^{t}(1+\mu_{h}\eta_{s})$. So (48) can be rewritten as
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq\frac{\mu_{h}D_{\psi}(x,x^{1})}{\Gamma_{T}-1}+2\mu_{h}(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\Gamma_{T}/\Gamma_{t-1}-1}.$ (49)
We can check that
$\eta_{t\in\left[T\right]}$ $=\frac{2}{\mu_{h}(t+4\kappa_{h})}\leq\frac{1}{2\kappa_{h}\mu_{h}}=\frac{1}{2L};$
$\eta_{t\in\left[T\right]}$ $=\begin{cases}\frac{1}{\mu_{h}(\eta+2\kappa_{h})}\leq\frac{1}{2\kappa_{h}\mu_{h}}=\frac{1}{2L}&t\leq\tau\\ \frac{2}{\mu_{h}(t-\tau+4\kappa_{h})}\leq\frac{1}{2\kappa_{h}\mu_{h}}=\frac{1}{2L}&t\geq\tau+1\end{cases}.$
Therefore, (49) is true for all cases.
If $\eta_{t\in\left[T\right]}=\frac{2}{\mu_{h}(t+4\kappa_{h})}$, we have
$\Gamma_{t\in\left\{0\right\}\cup\left[T\right]}=\prod_{s=1}^{t}(1+\mu_{h}\eta_{s})=\frac{(t+1+4\kappa_{h})(t+2+4\kappa_{h})}{(1+4\kappa_{h})(2+4\kappa_{h})}.$
Hence, by (49),
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$
$\leq$ $\frac{\mu_{h}D_{\psi}(x,x^{1})}{\frac{(T+1+4\kappa_{h})(T+2+4\kappa_{h})}{(1+4\kappa_{h})(2+4\kappa_{h})}-1}+\frac{8(M^{2}+\sigma^{2})}{\mu_{h}}\sum_{t=1}^{T}\frac{\frac{1}{(t+4\kappa_{h})^{2}}}{\frac{(T+1+4\kappa_{h})(T+2+4\kappa_{h})}{(t+4\kappa_{h})(t+1+4\kappa_{h})}-1}$
$=$ $\frac{(1+4\kappa_{h})(2+4\kappa_{h})\mu_{h}D_{\psi}(x,x^{1})}{T(T+3+8\kappa_{h})}+\frac{8(M^{2}+\sigma^{2})}{\mu_{h}}\sum_{t=1}^{T}\frac{t+1+4\kappa_{h}}{t+4\kappa_{h}}\cdot\frac{1}{(T+1-t)(T+2+8\kappa_{h}+t)}$
$\leq$ $\frac{(1+4\kappa_{h})(2+4\kappa_{h})\mu_{h}D_{\psi}(x,x^{1})}{T(T+3+8\kappa_{h})}+\frac{16(M^{2}+\sigma^{2})}{\mu_{h}}\sum_{t=1}^{T}\frac{1}{2T+3+8\kappa_{h}}\left(\frac{1}{T+1-t}+\frac{1}{T+2+8\kappa_{h}+t}\right)$
$\leq$ $\frac{(1+4\kappa_{h})(2+4\kappa_{h})\mu_{h}D_{\psi}(x,x^{1})}{T(T+3+8\kappa_{h})}+\frac{16(M^{2}+\sigma^{2})}{\mu_{h}}\cdot\frac{1+\log T+\log\frac{2T+2+8\kappa_{h}}{T+2+8\kappa_{h}}}{2T+3+8\kappa_{h}}$
$\leq$ $\frac{(1+4\kappa_{h})(2+4\kappa_{h})\mu_{h}D_{\psi}(x,x^{1})}{T(T+3+8\kappa_{h})}+\frac{16(M^{2}+\sigma^{2})(1+\log 2T)}{\mu_{h}(2T+3+8\kappa_{h})}$
$=$ $O\left(\frac{\mu_{h}(1+\kappa_{h})^{2}D_{\psi}(x,x^{1})}{T(T+\kappa_{h})}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{h}(T+\kappa_{h})}\right).$ (50)
If $\eta_{t\in\left[T\right]}=\begin{cases}\frac{1}{\mu_{h}(\eta+2\kappa_{h})}&t\leq\tau\\ \frac{2}{\mu_{h}(t-\tau+4\kappa_{h})}&t\geq\tau+1\end{cases}$, we know
$\Gamma_{t\in\left\{0\right\}\cup\left[T\right]}=\prod_{s=1}^{t}(1+\mu_{h}\eta_{s})=\begin{cases}\left(1+\frac{1}{\eta+2\kappa_{h}}\right)^{t}&t\leq\tau\\ \left(1+\frac{1}{\eta+2\kappa_{h}}\right)^{\tau}\frac{(t-\tau+1+4\kappa_{h})(t-\tau+2+4\kappa_{h})}{(1+4\kappa_{h})(2+4\kappa_{h})}&t\geq\tau+1\end{cases}.$
So we can obtain
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$
$\leq$ $\frac{\mu_{h}D_{\psi}(x,x^{1})}{\Gamma_{T}-1}+2\mu_{h}(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\Gamma_{T}/\Gamma_{t-1}-1}$
$=$ $\frac{\mu_{h}D_{\psi}(x,x^{1})}{\left(1+\frac{1}{\eta+2\kappa_{h}}\right)^{\tau}\frac{(T-\tau+1+4\kappa_{h})(T-\tau+2+4\kappa_{h})}{(1+4\kappa_{h})(2+4\kappa_{h})}-1}+2\mu_{h}(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\Gamma_{T}/\Gamma_{t-1}-1}$
$\overset{(a)}{\leq}$ $\frac{\mu_{h}D_{\psi}(x,x^{1})}{\left(1+\frac{1}{\eta+2\kappa_{h}}\right)^{T/2}-1}+2\mu_{h}(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\Gamma_{T}/\Gamma_{t-1}-1}$
$\overset{(b)}{\leq}$ $\frac{\mu_{h}D_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2(1+\eta+2\kappa_{h})}\right)-1}+2\mu_{h}(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\Gamma_{T}/\Gamma_{t-1}-1}$
$=$ $\frac{\mu_{h}D_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2(1+\eta+2\kappa_{h})}\right)-1}+2\mu_{h}(M^{2}+\sigma^{2})\left(\underbrace{\sum_{t=1}^{\tau}\frac{\eta_{t}^{2}}{\Gamma_{T}/\Gamma_{t-1}-1}}_{\text{I}}+\underbrace{\sum_{t=\tau+1}^{T}\frac{\eta_{t}^{2}}{\Gamma_{T}/\Gamma_{t-1}-1}}_{\text{II}}\right),$ (51)
where $(a)$ is by $T-\tau\geq 0$ and $\tau\geq\frac{T}{2}$, $(b)$ is due to
$\left(1+\frac{1}{\eta+2\kappa_{h}}\right)^{T/2}=\exp\left(\frac{T}{2}\log\left(1+\frac{1}{\eta+2\kappa_{h}}\right)\right)\geq\exp\left(\frac{T}{2(1+\eta+2\kappa_{h})}\right).$
Now we bound
I $=\frac{1}{\mu_{h}^{2}(\eta+2\kappa_{h})^{2}}\sum_{t=1}^{\tau}\frac{1}{\left(1+\frac{1}{\eta+2\kappa_{h}}\right)^{\tau-t+1}\frac{(T-\tau+1+4\kappa_{h})(T-\tau+2+4\kappa_{h})}{(1+4\kappa_{h})(2+4\kappa_{h})}-1}$
$\overset{(c)}{\leq}\frac{1}{\mu_{h}^{2}(\eta+2\kappa_{h})^{2}\left(1+\frac{1}{\eta+2\kappa_{h}}\right)}\sum_{t=1}^{\tau}\frac{1}{\frac{(T-\tau+1+4\kappa_{h})(T-\tau+2+4\kappa_{h})}{(1+4\kappa_{h})(2+4\kappa_{h})}-1}$
$=\frac{1}{\mu_{h}^{2}(\eta+2\kappa_{h})(\eta+1+2\kappa_{h})}\sum_{t=1}^{\tau}\frac{(1+4\kappa_{h})(2+4\kappa_{h})}{(T-\tau)(T-\tau+8\kappa_{h}+3)}$
$=\frac{(1+4\kappa_{h})(2+4\kappa_{h})}{\mu_{h}^{2}(\eta+2\kappa_{h})(\eta+1+2\kappa_{h})}\cdot\frac{\tau}{(T-\tau)(T-\tau+8\kappa_{h}+3)}$
$\overset{(d)}{\leq}\frac{2(1+4\kappa_{h})}{\mu_{h}^{2}(\eta+2\kappa_{h})}\cdot\frac{2(T+1)}{(T-1)(T+5+16\kappa_{h})}$
$\overset{(e)}{\leq}\frac{2\left(2+\frac{1}{\eta+2\kappa_{h}}\right)}{\mu_{h}^{2}}\cdot\frac{6}{T+5+16\kappa_{h})}=\frac{12\left(2+\frac{1}{\eta+2\kappa_{h}}\right)}{\mu_{h}^{2}(T+5+16\kappa_{h})},$ (52)
where $(c)$ is by $\tau-t\geq 0$ and $1+\frac{1}{\eta+2\kappa_{h}}\geq 1$. We use $\eta\geq 0$, $\tau\leq\frac{T+1}{2}$ in $(d)$ and $T\geq 2$ in $(e)$. Next, there is
II $=\frac{4}{\mu_{h}^{2}}\sum_{t=\tau+1}^{T}\frac{1}{(t-\tau+4\kappa_{h})^{2}\left(\frac{(T-\tau+1+4\kappa_{h})(T-\tau+2+4\kappa_{h})}{(t-\tau+4\kappa_{h})(t-\tau+1+4\kappa_{h})}-1\right)}$
$=\frac{4}{\mu_{h}^{2}}\sum_{t=1}^{T-\tau}\frac{1}{(t+4\kappa_{h})^{2}\left(\frac{(T-\tau+1+4\kappa_{h})(T-\tau+2+4\kappa_{h})}{(t+4\kappa_{h})(t+1+4\kappa_{h})}-1\right)}$
$=\frac{4}{\mu_{h}^{2}}\sum_{t=1}^{T-\tau}\frac{t+1+4\kappa_{h}}{(t+4\kappa_{h})\left((T-\tau+1+4\kappa_{h})(T-\tau+2+4\kappa_{h})-(t+4\kappa_{h})(t+1+4\kappa_{h})\right)}$
$\leq\frac{8}{\mu_{h}^{2}}\sum_{t=1}^{T-\tau}\frac{1}{(T-\tau+1+4\kappa_{h})(T-\tau+2+4\kappa_{h})-(t+4\kappa_{h})(t+1+4\kappa_{h})}$
$=\frac{8}{\mu_{h}^{2}}\sum_{t=1}^{T-\tau}\frac{1}{(T-\tau+1-t)(T-\tau+2+8\kappa_{h}+t)}$
$=\frac{8}{\mu_{h}^{2}}\sum_{t=1}^{T-\tau}\frac{1}{2T-2\tau+3+8\kappa_{h}}\left(\frac{1}{T-\tau+1-t}+\frac{1}{T-\tau+2+8\kappa_{h}+t}\right)$
$\leq\frac{8}{\mu_{h}^{2}(2T-2\tau+3+8\kappa_{h})}\left(1+\log(T-\tau)+\log\frac{2T-2\tau+2+8\kappa_{h}}{T-\tau+2+8\kappa_{h}}\right)$
$\leq\frac{8(1+\log T)}{\mu_{h}^{2}(T+2+8\kappa_{h})},$ (53)
where we use $\frac{T}{2}\leq\tau\leq\frac{T+1}{2}$ in the last inequality.
Plugging (52) and (53) into (51) to get
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{\mu_{h}D_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2(1+\eta+2\kappa_{h})}\right)-1}+\frac{2(M^{2}+\sigma^{2})}{\mu_{h}}\left(\frac{12\left(2+\frac{1}{\eta+2\kappa_{h}}\right)}{T+5+16\kappa_{h}}+\frac{8(1+\log T)}{T+2+8\kappa_{h}}\right)$
$=O\left(\frac{\mu_{h}D_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2(1+\eta+2\kappa_{h})}\right)-1}+\left(1\lor\frac{1}{\eta+2\kappa_{h}}\right)\frac{(M^{2}+\sigma^{2})\log T}{\mu_{h}(T+\kappa_{h})}\right).$ (54)
Theorem D.4.
Under Assumptions 2.-4. and 5B. with $\mu_{f}=0$ and $\mu_{h}>0$, let $\kappa_{h}\coloneqq\frac{L}{\mu_{h}}\geq 0$ and $\delta\in(0,1)$, for any $x\in\mathcal{X}$:
If $T$ is unknown, by taking $\eta_{t\in\left[T\right]}=\frac{2}{\mu_{h}(t+4\kappa_{h})}$, then with probability at least $1-\delta$, there is
$F(x^{T+1})-F(x)\leq O\left(\frac{\mu_{h}(1+\kappa_{h})^{2}D_{\psi}(x,x^{1})}{T(T+\kappa_{h})}+\frac{(M^{2}+\sigma^{2}\log\frac{1}{\delta})\log T}{\mu_{h}(T+\kappa_{h})}\right).$
If $T$ is known, by taking $\eta_{t\in\left[T\right]}=\begin{cases}\frac{1}{\mu_{h}(\eta+2\kappa_{h})}&t\leq\tau\\ \frac{2}{\mu_{h}(t-\tau+4\kappa_{h})}&t\geq\tau+1\end{cases}$ where $\eta\geq 0$ can be any number satisfying $\eta+\kappa_{h}>0$ and $\tau\coloneqq\left\lceil\frac{T}{2}\right\rceil$, then with probability at least $1-\delta$, there is
$F(x^{T+1})-F(x)\leq O\left(\frac{\mu_{h}D_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2(1+\eta+2\kappa_{h})}\right)-1}+\left(1\lor\frac{1}{\eta+2\kappa_{h}}\right)\frac{(M^{2}+\sigma^{2}\log\frac{1}{\delta})\log T}{\mu_{h}(T+\kappa_{h})}\right).$
Proof.
When $\mu_{f}=0$ and $\mu_{h}>0$, suppose the condition of $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}=\frac{1}{2L}$ in Lemma 4.3 holds, we will have with probability at least $1-\delta$
$F(x^{T+1})-F(x)$ $\leq 2\left(1+\max_{2\leq t\leq T}\frac{1}{1-\mu_{f}\eta_{t}}\right)\left[\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+\left(M^{2}+\sigma^{2}\left(1+2\log\frac{2}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}\right]$
$=\frac{4D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+4\left(M^{2}+\sigma^{2}\left(1+2\log\frac{2}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}.$ (55)
Observe that
$\gamma_{t\in\left[T\right]}=\eta_{t}\prod_{s=2}^{t}\frac{1+\mu_{h}\eta_{s-1}}{1-\mu_{f}\eta_{s}}=\eta_{t}\prod_{s=2}^{t}(1+\mu_{h}\eta_{s-1})=\eta_{t}\Gamma_{t-1}=\frac{\Gamma_{t}-\Gamma_{t-1}}{\mu_{h}},$
where $\Gamma_{t\in\left\{0\right\}\cup\left[T\right]}\coloneqq\prod_{s=1}^{t}(1+\mu_{h}\eta_{s})$. So (55) can be rewritten as
$F(x^{T+1})-F(x)\leq\frac{4\mu_{h}D_{\psi}(x,x^{1})}{\Gamma_{T}-1}+4\mu_{h}\left(M^{2}+\sigma^{2}\left(1+2\log\frac{2}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\Gamma_{T}/\Gamma_{t-1}-1}.$ (56)
We can check that
$\eta_{t\in\left[T\right]}$ $=\frac{2}{\mu_{h}(t+4\kappa_{h})}\leq\frac{1}{2\kappa_{h}\mu_{h}}=\frac{1}{2L};$
$\eta_{t\in\left[T\right]}$ $=\begin{cases}\frac{1}{\mu_{h}(\eta+2\kappa_{h})}\leq\frac{1}{2\kappa_{h}\mu_{h}}=\frac{1}{2L}&t\leq\tau\\ \frac{2}{\mu_{h}(t-\tau+4\kappa_{h})}\leq\frac{1}{2\kappa_{h}\mu_{h}}=\frac{1}{2L}&t\geq\tau+1\end{cases}.$
Therefore, (56) is true for all cases.
If $\eta_{t\in\left[T\right]}=\frac{2}{\mu_{h}(t+4\kappa_{h})}$, following the similar steps in the proof of (50), we can finally get
$F(x^{T+1})-F(x)\leq O\left(\frac{\mu_{h}(1+\kappa_{h})^{2}D_{\psi}(x,x^{1})}{T(T+\kappa_{h})}+\frac{(M^{2}+\sigma^{2}\log\frac{1}{\delta})\log T}{\mu_{h}(T+\kappa_{h})}\right).$
If $\eta_{t\in\left[T\right]}=\begin{cases}\frac{1}{\mu_{h}(\eta+2\kappa_{h})}&t\leq\tau\\ \frac{2}{\mu_{h}(t-\tau+4\kappa_{h})}&t\geq\tau+1\end{cases}$, following the similar steps in the proof of (54), we can finally get
$F(x^{T+1})-F(x)\leq O\left(\frac{\mu_{h}D_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2(1+\eta+2\kappa_{h})}\right)-1}+\left(1\lor\frac{1}{\eta+2\kappa_{h}}\right)\frac{(M^{2}+\sigma^{2}\log\frac{1}{\delta})\log T}{\mu_{h}(T+\kappa_{h})}\right).$
Appendix E Unified Theoretical Analysis under Heavy-Tailed Noise
In this section, we present the analysis under heavy-tailed noise. Unlike Lemma 4.1, the requirement of $\eta_{t\in\left[T\right]}\leq\frac{1}{2L}$ is removed.
Lemma E.1.
Under Assumptions 2. and 3. with $\mu_{f}=\mu_{h}=0$, let $v_{t\in\left\{0\right\}\cup\left[T\right]}>0$ be defined as $v_{t}\coloneqq\frac{\eta_{T}}{\sum_{s=t\lor 1}^{T}\eta_{s}}$, then for any $x\in\mathcal{X}$, we have
$\eta_{T}v_{T}\left(F(x^{T+1})-F(x)\right)\leq$ $v_{0}D_{\psi}(x,x^{1})+\sum_{t=1}^{T}\eta_{t}v_{t-1}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle$
$+\sum_{t=1}^{T}\frac{\eta_{t}^{\frac{p}{2-p}}v_{t}(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}+\eta_{t}^{p}v_{t}4^{p-1}(M^{p}+\left\|\xi^{t}\right\|_{*}^{p})}{p},$
where $\xi^{t}\coloneqq\widehat{g}^{t}-\mathbb{\mathbb{E}}\left[\widehat{g}^{t}\mid\mathcal{F}^{t-1}\right],\forall t\in\left[T\right]$ and $z^{t}\coloneqq\frac{v_{0}}{v_{t}}x+\sum_{s=1}^{t}\frac{v_{s}-v_{s-1}}{v_{t}}x^{s},\forall t\in\left\{0\right\}\cup\left[T\right]$.
Proof.
The proof is similar to the proof of Lemma 4.1. Again, we use the sequence
$z^{t}=\begin{cases}\left(1-\frac{v_{t-1}}{v_{t}}\right)x^{t}+\frac{v_{t-1}}{v_{t}}z^{t-1}&t\in\left[T\right]\\ x&t=0\end{cases}\Leftrightarrow z^{t}=\frac{v_{0}}{v_{t}}x+\sum_{s=1}^{t}\frac{v_{s}-v_{s-1}}{v_{t}}x^{s},\forall t\in\left\{0\right\}\cup\left[T\right].$
As before, we start the proof from the $(L,M)$-smoothness of $f$:
$f(x^{t+1})-f(x^{t})\leq$ $\left\langle g^{t},x^{t+1}-x^{t}\right\rangle+\frac{L}{2}\left\|x^{t+1}-x^{t}\right\|^{2}+M\left\|x^{t+1}-x^{t}\right\|$
$=$ $\left\langle\xi^{t},z^{t}-x^{t}\right\rangle+\underbrace{\left\langle\xi^{t},x^{t}-x^{t+1}\right\rangle}_{\text{I}}+\underbrace{\left\langle\widehat{g}^{t},x^{t+1}-z^{t}\right\rangle}_{\text{II}}$
$+\underbrace{\left\langle g^{t},z^{t}-x^{t}\right\rangle}_{\text{III}}+\underbrace{\frac{L}{2}\left\|x^{t+1}-x^{t}\right\|^{2}+M\left\|x^{t+1}-x^{t}\right\|}_{\text{IV}},$ (57)
where $g^{t}\coloneqq\mathbb{\mathbb{E}}\left[\widehat{g}^{t}|\mathcal{F}^{t-1}\right]\in\partial f(x^{t})$ and $\xi^{t}=\widehat{g}^{t}-g^{t}$. Next we bound these four terms respectively.
  • For term I, by applying Cauchy-Schwarz inequality, Young’s inequality and the $(1,\frac{p}{p-1})$ uniform convexity of $\psi$ (see (1)), we can get the following upper bound
    I $\leq\left\|\xi^{t}\right\|_{*}\left\|x^{t}-x^{t+1}\right\|=(4\eta_{t})^{\frac{p-1}{p}}\left\|\xi^{t}\right\|_{*}\cdot\frac{\left\|x^{t}-x^{t+1}\right\|}{(4\eta_{t})^{\frac{p-1}{p}}}$
    $\leq\frac{(4\eta_{t})^{p-1}\left\|\xi^{t}\right\|_{*}^{p}}{p}+\frac{\left\|x^{t}-x^{t+1}\right\|^{\frac{p}{p-1}}}{\frac{p}{p-1}\cdot 4\eta_{t}}\leq\frac{\eta_{t}^{p-1}4^{p-1}\left\|\xi^{t}\right\|_{*}^{p}}{p}+\frac{D_{\psi}(x^{t+1},x^{t})}{4\eta_{t}}.$ (58)
  • For term II, the same as (7) with $\mu_{h}=0$, we can obtain
    $\text{II}\leq\frac{D_{\psi}(z^{t},x^{t})-D_{\psi}(z^{t},x^{t+1})-D_{\psi}(x^{t+1},x^{t})}{\eta_{t}}+h(z_{t})-h(x_{t+1}).$ (59)
  • For term III, we simply use the convexity of $f$ to get
    $\text{III}\leq f(z^{t})-f(x^{t}).$ (60)
  • For term IV, we have
    IV $=\frac{L}{2}(4\eta_{t})^{\frac{2(p-1)}{p}}\cdot\frac{\left\|x^{t+1}-x^{t}\right\|^{2}}{(4\eta_{t})^{\frac{2(p-1)}{p}}}+(4\eta_{t})^{\frac{p-1}{p}}M\cdot\frac{\left\|x^{t+1}-x^{t}\right\|}{(4\eta_{t})^{\frac{p-1}{p}}}$
    $\leq\frac{\left[\frac{L}{2}(4\eta_{t})^{\frac{2(p-1)}{p}}\right]^{\frac{p}{2-p}}}{\frac{p}{2-p}}+\frac{\left\|x^{t+1}-x^{t}\right\|^{\frac{p}{p-1}}}{\frac{p}{2(p-1)}\cdot 4\eta_{t}}+\frac{(4\eta_{t})^{p-1}M^{p}}{p}+\frac{\left\|x^{t+1}-x^{t}\right\|^{\frac{p}{p-1}}}{\frac{p}{p-1}\cdot 4\eta_{t}}$
    $\leq\frac{\eta_{t}^{\frac{2(p-1)}{2-p}}(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}}{p}+\frac{\eta_{t}^{p-1}4^{p-1}M^{p}}{p}+\frac{3D_{\psi}(x^{t+1},x^{t})}{4\eta_{t}},$ (61)
    where the second line is due to Young’s inequality and the last inequality holds by the $(1,\frac{p}{p-1})$ uniform convexity of $\psi$.
By plugging the bounds (58), (59), (60) and (61) into (57), we obtain
$f(x^{t+1})-f(x^{t})\leq$ $\left\langle\xi^{t},z^{t}-x^{t}\right\rangle+\frac{\eta_{t}^{p-1}4^{p-1}\left\|\xi^{t}\right\|_{*}^{p}}{p}+\frac{D_{\psi}(x^{t+1},x^{t})}{4\eta_{t}}$
$+\frac{D_{\psi}(z^{t},x^{t})-D_{\psi}(z^{t},x^{t+1})-D_{\psi}(x^{t+1},x^{t})}{\eta_{t}}+h(z_{t})-h(x_{t+1})$
$+f(z^{t})-f(x^{t})+\frac{\eta_{t}^{\frac{2(p-1)}{2-p}}(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}}{p}+\frac{\eta_{t}^{p-1}4^{p-1}M^{p}}{p}+\frac{3D_{\psi}(x^{t+1},x^{t})}{4\eta_{t}}.$
Rearranging the terms to get
$F(x^{t+1})-F(z^{t})\leq$ $\left\langle\xi^{t},z^{t}-x^{t}\right\rangle+\frac{D_{\psi}(z^{t},x^{t})-D_{\psi}(z^{t},x^{t+1})}{\eta_{t}}$
$+\frac{\eta_{t}^{\frac{2(p-1)}{2-p}}(2-p)L^{\frac{p}{2-p}}2^{\frac{3p-4}{2-p}}+\eta_{t}^{p-1}4^{p-1}(M^{p}+\left\|\xi^{t}\right\|_{*}^{p})}{p}$
$\leq$ $\frac{v_{t-1}}{v_{t}}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle+\frac{\frac{v_{t-1}}{v_{t}}D_{\psi}(z^{t-1},x^{t})-D_{\psi}(z^{t},x^{t+1})}{\eta_{t}}$
$+\frac{\eta_{t}^{\frac{2(p-1)}{2-p}}(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}+\eta_{t}^{p-1}4^{p-1}(M^{p}+\left\|\xi^{t}\right\|_{*}^{p})}{p},$ (62)
where the last line is again by $z^{t}-x^{t}=\frac{v_{t-1}}{v_{t}}\left(z^{t-1}-x^{t}\right)$ and $D_{\psi}(z^{t},x^{t})\leq\frac{v_{t-1}}{v_{t}}D_{\psi}(z^{t-1},x^{t})$.
Multiplying both sides of (62) by $\eta_{t}v_{t}$ (these two terms are non-negative) and summing up from $t=1$ to $T$, we obtain
$\sum_{t=1}^{T}\eta_{t}v_{t}\left(F(x^{t+1})-F(z^{t})\right)\leq$ $\sum_{t=1}^{T}\eta_{t}v_{t-1}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle+\sum_{t=1}^{T}v_{t-1}D_{\psi}(z^{t-1},x^{t})-v_{t}D_{\psi}(z^{t},x^{t+1})$
$+\sum_{t=1}^{T}\frac{\eta_{t}^{\frac{p}{2-p}}v_{t}(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}+\eta_{t}^{p}v_{t}4^{p-1}(M^{p}+\left\|\xi^{t}\right\|_{*}^{p})}{p}$
$=$ $v_{0}D_{\psi}(x,x^{1})-v_{T}D_{\psi}(z^{T},x^{T+1})+\sum_{t=1}^{T}\eta_{t}v_{t-1}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle$
$+\sum_{t=1}^{T}\frac{\eta_{t}^{\frac{p}{2-p}}v_{t}(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}+\eta_{t}^{p}v_{t}4^{p-1}(M^{p}+\left\|\xi^{t}\right\|_{*}^{p})}{p}.$ (63)
Similar to the proof of Lemma 4.1, we can prove
$\sum_{t=1}^{T}\eta_{t}v_{t}\left(F(x^{t+1})-F(z^{t})\right)\geq\eta_{T}v_{T}\left(F(x^{T+1})-F(x)\right),$ (64)
Combining (63) and (64), there is
$\eta_{T}v_{T}\left(F(x^{T+1})-F(x)\right)\leq$ $v_{0}D_{\psi}(x,x^{1})+\sum_{t=1}^{T}\eta_{t}v_{t-1}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle$
$+\sum_{t=1}^{T}\frac{\eta_{t}^{\frac{p}{2-p}}v_{t}(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}+\eta_{t}^{p}v_{t}4^{p-1}(M^{p}+\left\|\xi^{t}\right\|_{*}^{p})}{p}.$
Now we are ready to state the general in-expectation convergence result in Lemma E.2.
Lemma E.2.
Under Assumptions 2.-4. and 5C. with $\mu_{f}=\mu_{h}=0$, then for any $x\in\mathcal{X}$ we have
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\eta_{t}}+\frac{(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}}{p}\sum_{t=1}^{T}\frac{\eta_{t}^{\frac{p}{2-p}}}{\sum_{s=t}^{T}\eta_{s}}+\frac{4^{p-1}(M^{p}+\sigma^{p})}{p}\sum_{t=1}^{T}\frac{\eta_{t}^{p}}{\sum_{s=t}^{T}\eta_{s}}.$
Proof.
We invoke Lemma E.1 to get
$\eta_{T}v_{T}\left(F(x^{T+1})-F(x)\right)\leq$ $v_{0}D_{\psi}(x,x^{1})+\sum_{t=1}^{T}\eta_{t}v_{t-1}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle$
$+\sum_{t=1}^{T}\frac{\eta_{t}^{\frac{p}{2-p}}v_{t}(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}+\eta_{t}^{p}v_{t}4^{p-1}(M^{p}+\left\|\xi^{t}\right\|_{*}^{p})}{p}.$
Take expectations on both sides to obtain
$\eta_{T}v_{T}\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq$ $v_{0}D_{\psi}(x,x^{1})+\sum_{t=1}^{T}\eta_{t}v_{t-1}\mathbb{\mathbb{E}}\left[\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle\right]$
$+\sum_{t=1}^{T}\frac{\eta_{t}^{\frac{p}{2-p}}v_{t}(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}+\eta_{t}^{p}v_{t}4^{p-1}(M^{p}+\mathbb{\mathbb{E}}\left[\left\|\xi^{t}\right\|_{*}^{p}\right])}{p}$
$\leq$ $v_{0}D_{\psi}(x,x^{1})+\sum_{t=1}^{T}\frac{\eta_{t}^{\frac{p}{2-p}}v_{t}(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}+\eta_{t}^{p}v_{t}4^{p-1}(M^{p}+\sigma^{p})}{p},$
where the last line is due to $\mathbb{\mathbb{E}}\left[\left\|\xi^{t}\right\|_{*}^{p}\right]=\mathbb{\mathbb{E}}\left[\mathbb{\mathbb{E}}\left[\left\|\xi^{t}\right\|_{*}^{p}\mid\mathcal{F}^{t-1}\right]\right]\leq\sigma^{p}$ (Assumption 5C.) and $\mathbb{\mathbb{E}}\left[\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle\right]=\mathbb{\mathbb{E}}\left[\left\langle\mathbb{\mathbb{E}}\left[\xi^{t}|\mathcal{F}^{t-1}\right],z^{t-1}-x^{t}\right\rangle\right]=0$ ($z^{t-1}-x^{t}\in\mathcal{F}^{t-1}$ and Assumption 4.). Finally, we divide both sides by $\eta_{T}v_{T}$ and plug in $v_{t\in\left\{0\right\}\cup\left[T\right]}=\frac{\eta_{T}}{\sum_{s=t\lor 1}^{T}\eta_{s}}$ to finish the proof. ∎
Appendix F General Convex Functions under Heavy-Tailed Noise
In this section, we provide the full version of Theorem 5.2 by using Lemma E.2,
Theorem F.1.
Under Assumptions 2.-4. and 5C. with $\mu_{f}=\mu_{h}=0$, for any $x\in\mathcal{X}$:
If $T$ is unknown, by taking $\eta_{t\in\left[T\right]}=\frac{\eta_{*}}{t^{\frac{2-p}{p}}}\land\frac{\eta}{t^{\frac{1}{p}}}$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq$ $O\left(\frac{1}{T^{2-\frac{2}{p}}}\left(\frac{D_{\psi}(x,x^{1})}{\eta_{*}}+\eta_{*}^{\frac{2p-2}{2-p}}L^{\frac{p}{2-p}}\log T+\frac{\eta^{p}(M^{p}+\sigma^{p})}{\eta_{*}}\log T\right)\right$.
$+\left.\frac{1}{T^{1-\frac{1}{p}}}\left(\frac{D_{\psi}(x,x^{1})}{\eta}+\eta^{p-1}(M^{p}+\sigma^{p})\log T+\frac{\eta_{*}^{\frac{p}{2-p}}L^{\frac{p}{2-p}}}{\eta}\log T\right)\right).$
In particular, by choosing $\eta_{*}=\Theta\left(\frac{\left[D_{\psi}(x,x^{1})\right]^{\frac{2-p}{p}}}{L}\right)$ and $\eta=\Theta\left(\left[\frac{D_{\psi}(x,x^{1})}{M^{p}+\sigma^{p}}\right]^{\frac{1}{p}}\right)$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{L\left[D_{\psi}(x,x^{1})\right]^{2-\frac{2}{p}}\log T}{T^{2-\frac{2}{p}}}+\frac{(M+\sigma)\left[D_{\psi}(x,x^{1})\right]^{1-\frac{1}{p}}\log T}{T^{1-\frac{1}{p}}}\right).$
If $T$ is known, by taking $\eta_{t\in\left[T\right]}=\frac{\eta_{*}}{T^{\frac{2-p}{p}}}\land\frac{\eta}{T^{\frac{1}{p}}}$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq$ $O\left(\frac{1}{T^{2-\frac{2}{p}}}\left(\frac{D_{\psi}(x,x^{1})}{\eta_{*}}+\eta_{*}^{\frac{2p-2}{2-p}}L^{\frac{p}{2-p}}\log T\right)\right$.
$+\left.\frac{1}{T^{1-\frac{1}{p}}}\left(\frac{D_{\psi}(x,x^{1})}{\eta}+\eta^{p-1}(M^{p}+\sigma^{p})\log T\right)\right).$
In particular, by choosing $\eta_{*}=\Theta\left(\frac{1}{L}\left[\frac{D_{\psi}(x,x^{1})}{\log T}\right]^{\frac{2-p}{p}}\right)$ and $\eta=\Theta\left(\left[\frac{D_{\psi}(x,x^{1})}{(M^{p}+\sigma^{p})\log T}\right]^{\frac{1}{p}}\right)$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{L\left[D_{\psi}(x,x^{1})\right]^{2-\frac{2}{p}}(\log T)^{\frac{2}{p}-1}}{T^{2-\frac{2}{p}}}+\frac{(M+\sigma)\left[D_{\psi}(x,x^{1})\right]^{1-\frac{1}{p}}(\log T)^{\frac{1}{p}}}{T^{1-\frac{1}{p}}}\right).$
Proof.
By Lemma E.2, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\eta_{t}}+\frac{(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}}{p}\sum_{t=1}^{T}\frac{\eta_{t}^{\frac{p}{2-p}}}{\sum_{s=t}^{T}\eta_{s}}+\frac{4^{p-1}(M^{p}+\sigma^{p})}{p}\sum_{t=1}^{T}\frac{\eta_{t}^{p}}{\sum_{s=t}^{T}\eta_{s}}.$
Hence, we only need to plug in different step sizes. Before doing so, we first prove the following simple fact for any $a\in(0,1)$,
$\sum_{t=1}^{T}\frac{\sum_{s=t}^{T}s^{a}}{t(T-t+1)^{2}}\leq\sum_{t=1}^{T}\frac{T^{a}}{t(T-t+1)}=\frac{1}{T^{1-a}}\sum_{t=1}^{T}\frac{1}{t}+\frac{1}{T-t+1}\leq\frac{2(1+\log T)}{T^{1-a}}.$ (65)
If $\eta_{t\in\left[T\right]}=\frac{\eta_{*}}{t^{\frac{2-p}{p}}}\land\frac{\eta}{t^{\frac{1}{p}}}$, we have
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq$ $\frac{D_{\psi}(x,x^{1})}{T^{2}}\left(\sum_{t=1}^{T}\frac{1}{\eta_{t}}\right)+\frac{(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}}{p}\sum_{t=1}^{T}\frac{\eta_{t}^{\frac{p}{2-p}}}{(T-t+1)^{2}}\left(\sum_{s=t}^{T}\frac{1}{\eta_{s}}\right)$
$+\frac{4^{p-1}(M^{p}+\sigma^{p})}{p}\sum_{t=1}^{T}\frac{\eta_{t}^{p}}{(T-t+1)^{2}}\left(\sum_{s=t}^{T}\frac{1}{\eta_{s}}\right).$
We can bound
$\sum_{t=1}^{T}\frac{1}{\eta_{t}}$ $\leq\sum_{t=1}^{T}\frac{t^{\frac{2-p}{p}}}{\eta_{*}}+\frac{t^{\frac{1}{p}}}{\eta}\leq\frac{T^{\frac{2-p}{p}}+\int_{1}^{T}t^{\frac{2-p}{p}}\mathrm{d}t}{\eta_{*}}+\frac{T^{\frac{1}{p}}+\int_{1}^{T}t^{\frac{1}{p}}\mathrm{d}t}{\eta}$
$\leq\frac{T^{\frac{2-p}{p}}+\frac{p}{2}T^{\frac{2}{p}}}{\eta_{*}}+\frac{T^{\frac{1}{p}}+\frac{p}{p+1}T^{\frac{1}{p}+1}}{\eta}\overset{(a)}{\leq}\frac{T^{\frac{2}{p}}}{\eta_{*}}+\frac{5T^{\frac{1}{p}+1}}{3\eta},$
where $(a)$ is by $p\leq 2$ and $T\geq 1$. For the other two terms, we have
$\sum_{t=1}^{T}\frac{\eta_{t}^{\frac{p}{2-p}}}{(T-t+1)^{2}}\left(\sum_{s=t}^{T}\frac{1}{\eta_{s}}\right)$ $\leq\sum_{t=1}^{T}\frac{\eta_{t}^{\frac{p}{2-p}}}{(T-t+1)^{2}}\left(\sum_{s=t}^{T}\frac{s^{\frac{2-p}{p}}}{\eta_{*}}+\frac{s^{\frac{1}{p}}}{\eta}\right)$
$\leq\eta_{*}^{\frac{2p-2}{2-p}}\sum_{t=1}^{T}\frac{\sum_{s=t}^{T}s^{\frac{2-p}{p}}}{t(T-t+1)^{2}}+\frac{\eta_{*}^{\frac{p}{2-p}}}{\eta}\sum_{t=1}^{T}\frac{\sum_{s=t}^{T}s^{\frac{1}{p}}}{t(T-t+1)^{2}}$
[…]
$\sum_{t=1}^{T}\frac{\eta_{t}^{p}}{(T-t+1)^{2}}\left(\sum_{s=t}^{T}\frac{1}{\eta_{s}}\right)$ $\leq\sum_{t=1}^{T}\frac{\eta_{t}^{p}}{(T-t+1)^{2}}\left(\sum_{s=t}^{T}\frac{s^{\frac{2-p}{p}}}{\eta_{*}}+\frac{s^{\frac{1}{p}}}{\eta}\right)$
$\leq\frac{\eta^{p}}{\eta_{*}}\sum_{t=1}^{T}\frac{\sum_{s=t}^{T}s^{\frac{2-p}{p}}}{t(T-t+1)^{2}}+\eta^{p-1}\sum_{t=1}^{T}\frac{\sum_{s=t}^{T}s^{\frac{1}{p}}}{t(T-t+1)^{2}}$
[…]
Hence, we have
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq$ $\frac{D_{\psi}(x,x^{1})}{\eta_{*}T^{2-\frac{2}{p}}}+\frac{5D_{\psi}(x,x^{1})}{3\eta T^{1-\frac{1}{p}}}$
$+\frac{(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}}{p}\left[\eta_{*}^{\frac{2p-2}{2-p}}\frac{2(1+\log T)}{T^{2-\frac{2}{p}}}+\frac{\eta_{*}^{\frac{p}{2-p}}}{\eta}\cdot\frac{2(1+\log T)}{T^{1-\frac{1}{p}}}\right]$
$+\frac{4^{p-1}(M^{p}+\sigma^{p})}{p}\left[\frac{\eta^{p}}{\eta_{*}}\cdot\frac{2(1+\log T)}{T^{2-\frac{2}{p}}}+\eta^{p-1}\frac{2(1+\log T)}{T^{1-\frac{1}{p}}}\right]$
$=$ $O\left(\frac{1}{T^{2-\frac{2}{p}}}\left(\frac{D_{\psi}(x,x^{1})}{\eta_{*}}+\eta_{*}^{\frac{2p-2}{2-p}}L^{\frac{p}{2-p}}\log T+\frac{\eta^{p}(M^{p}+\sigma^{p})}{\eta_{*}}\log T\right)\right$.
$+\left.\frac{1}{T^{1-\frac{1}{p}}}\left(\frac{D_{\psi}(x,x^{1})}{\eta}+\eta^{p-1}(M^{p}+\sigma^{p})\log T+\frac{\eta_{*}^{\frac{p}{2-p}}L^{\frac{p}{2-p}}}{\eta}\log T\right)\right).$
By plugging in $\eta_{*}=\Theta\left(\frac{\left[D_{\psi}(x,x^{1})\right]^{\frac{2-p}{p}}}{L}\right)$ and $\eta=\Theta\left(\left[\frac{D_{\psi}(x,x^{1})}{M^{p}+\sigma^{p}}\right]^{\frac{1}{p}}\right)$, we get the desired bound.
If $\eta_{t\in\left[T\right]}=\frac{\eta_{*}}{T^{\frac{2-p}{p}}}\land\frac{\eta}{T^{\frac{1}{p}}}$, we will obtain
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq$ $\frac{D_{\psi}(x,x^{1})}{T}\left(\frac{T^{\frac{2-p}{p}}}{\eta_{*}}\lor\frac{T^{\frac{1}{p}}}{\eta}\right)+\frac{(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}}{p}\left(\frac{\eta_{*}}{T^{\frac{2-p}{p}}}\land\frac{\eta}{T^{\frac{1}{p}}}\right)^{\frac{2p-2}{2-p}}\sum_{t=1}^{T}\frac{1}{T-t+1}$
$+\frac{4^{p-1}(M^{p}+\sigma^{p})}{p}\left(\frac{\eta_{*}}{T^{\frac{2-p}{p}}}\land\frac{\eta}{T^{\frac{1}{p}}}\right)^{p-1}\sum_{t=1}^{T}\frac{1}{T-t+1}$
$\leq$ $\frac{1}{T^{\frac{2p-2}{p}}}\left[\frac{D_{\psi}(x,x^{1})}{\eta_{*}}+\frac{\eta_{*}^{\frac{2p-2}{2-p}}(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}\log T}{p}\right]$
$+\frac{1}{T^{2-\frac{2}{p}}}\left[\frac{D_{\psi}(x,x^{1})}{\eta}+\frac{\eta^{p-1}4^{p-1}(M^{p}+\sigma^{p})\log T}{p}\right]$
$=$ $O\left(\frac{1}{T^{2-\frac{2}{p}}}\left(\frac{D_{\psi}(x,x^{1})}{\eta_{*}}+\eta_{*}^{\frac{2p-2}{2-p}}L^{\frac{p}{2-p}}\log T\right)\right$.
$+\left.\frac{1}{T^{1-\frac{1}{p}}}\left(\frac{D_{\psi}(x,x^{1})}{\eta}+\eta^{p-1}(M^{p}+\sigma^{p})\log T\right)\right).$
By plugging in $\eta_{*}=\Theta\left(\frac{1}{L}\left[\frac{D_{\psi}(x,x^{1})}{\log T}\right]^{\frac{2-p}{p}}\right)$ and $\eta=\Theta\left(\left[\frac{D_{\psi}(x,x^{1})}{(M^{p}+\sigma^{p})\log T}\right]^{\frac{1}{p}}\right)$, we get the desired bound. ∎
F.1 Optimal Rate under Heavy-Tailed Noise
The full version of Theorem 5.3 is stated as follows.
Theorem F.2.
Under Assumptions 2.-4. and 5C. with $\mu_{f}=\mu_{h}=0$, for any $x\in\mathcal{X}$, if $T$ is known, by taking $\eta_{t\in\left[T\right]}=\frac{\eta_{*}(T-t+1)}{T^{\frac{2}{p}}}\land\frac{\eta(T-t+1)}{T^{\frac{1}{p}+1}}$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{1}{T^{2-\frac{2}{p}}}\left(\frac{D_{\psi}(x,x^{1})}{\eta_{*}}+\eta_{*}^{\frac{2p-2}{2-p}}L^{\frac{p}{2-p}}\right)+\frac{1}{T^{1-\frac{1}{p}}}\left(\frac{D_{\psi}(x,x^{1})}{\eta}+\eta^{p-1}(M^{p}+\sigma^{p})\right)\right).$
In particular, by choosing $\eta_{*}=\Theta\left(\frac{\left[D_{\psi}(x,x^{1})\right]^{\frac{2-p}{p}}}{L}\right)$ and $\eta=\Theta\left(\left[\frac{D_{\psi}(x,x^{1})}{M^{p}+\sigma^{p}}\right]^{\frac{1}{p}}\right)$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{L\left[D_{\psi}(x,x^{1})\right]^{2-\frac{2}{p}}}{T^{2-\frac{2}{p}}}+\frac{(M+\sigma)\left[D_{\psi}(x,x^{1})\right]^{1-\frac{1}{p}}}{T^{1-\frac{1}{p}}}\right).$
Proof.
By Lemma E.2, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\eta_{t}}+\frac{(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}}{p}\sum_{t=1}^{T}\frac{\eta_{t}^{\frac{p}{2-p}}}{\sum_{s=t}^{T}\eta_{s}}+\frac{4^{p-1}(M^{p}+\sigma^{p})}{p}\sum_{t=1}^{T}\frac{\eta_{t}^{p}}{\sum_{s=t}^{T}\eta_{s}}.$
For $\eta_{t\in\left[T\right]}=\frac{\eta_{*}(T-t+1)}{T^{\frac{2}{p}}}\land\frac{\eta(T-t+1)}{T^{\frac{1}{p}+1}}$, we can compute
$\sum_{t=1}^{T}\eta_{t}=\left(\frac{\eta_{*}}{T^{\frac{2}{p}}}\land\frac{\eta}{T^{\frac{1}{p}+1}}\right)\sum_{t=1}^{T}T-t+1\geq\frac{\eta_{*}}{2}T^{2-\frac{2}{p}}\land\frac{\eta}{2}T^{1-\frac{1}{p}};$
$\sum_{t=1}^{T}\frac{\eta_{t}^{\frac{p}{2-p}}}{\sum_{s=t}^{T}\eta_{s}}$ $=\left(\frac{\eta_{*}}{T^{\frac{2}{p}}}\land\frac{\eta}{T^{\frac{1}{p}+1}}\right)^{\frac{2p-2}{2-p}}\sum_{t=1}^{T}\frac{(T-t+1)^{\frac{p}{2-p}}}{\sum_{s=t}^{T}T-s+1}\leq\frac{2\eta_{*}^{\frac{2p-2}{2-p}}}{T^{\frac{4p-4}{p(2-p)}}}\sum_{t=1}^{T}(T-t+1)^{\frac{3p-4}{2-p}}$
$\leq\frac{2\eta_{*}^{\frac{2p-2}{2-p}}}{T^{\frac{4p-4}{p(2-p)}}}\begin{cases}\int_{1}^{T+1}t^{\frac{3p-4}{2-p}}\mathrm{d}t&p\in\left[\frac{4}{3},2\right)\\ \int_{0}^{T}t^{\frac{3p-4}{2-p}}\mathrm{d}t&p\in\left(1,\frac{4}{3}\right)\end{cases}\leq\frac{2^{\frac{2p-2}{2-p}}(2-p)\eta_{*}^{\frac{2p-2}{2-p}}}{(p-1)T^{2-\frac{2}{p}}};$
$\sum_{t=1}^{T}\frac{\eta_{t}^{p}}{\sum_{s=t}^{T}\eta_{s}}$ $=\left(\frac{\eta_{*}}{T^{\frac{2}{p}}}\land\frac{\eta}{T^{\frac{1}{p}+1}}\right)^{p-1}\sum_{t=1}^{T}\frac{(T-t+1)^{p}}{\sum_{s=t}^{T}T-s+1}\leq\frac{2\eta^{p-1}}{T^{\frac{p^{2}-1}{p}}}\sum_{t=1}^{T}(T-t+1)^{p-2}$
$\leq\frac{2\eta^{p-1}}{T^{\frac{p^{2}-1}{p}}}\int_{0}^{T}t^{p-2}\mathrm{d}t=\frac{2\eta^{p-1}}{(p-1)T^{1-\frac{1}{p}}}.$
Hence, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{2D_{\psi}(x,x^{1})}{\eta_{*}T^{2-\frac{2}{p}}}+\frac{2D_{\psi}(x,x^{1})}{\eta T^{1-\frac{1}{p}}}+\frac{(2-p)^{2}2^{\frac{5p-6}{2-p}}L^{\frac{p}{2-p}}\eta_{*}^{\frac{2p-2}{2-p}}}{p(p-1)T^{2-\frac{2}{p}}}+\frac{4^{p-\frac{1}{2}}(M^{p}+\sigma^{p})\eta^{p-1}}{p(p-1)T^{1-\frac{1}{p}}}$
$=O\left(\frac{1}{T^{2-\frac{2}{p}}}\left(\frac{D_{\psi}(x,x^{1})}{\eta_{*}}+\eta_{*}^{\frac{2p-2}{2-p}}L^{\frac{p}{2-p}}\right)+\frac{1}{T^{1-\frac{1}{p}}}\left(\frac{D_{\psi}(x,x^{1})}{\eta}+\eta^{p-1}(M^{p}+\sigma^{p})\right)\right).$
By plugging in $\eta_{*}=\Theta\left(\frac{\left[D_{\psi}(x,x^{1})\right]^{\frac{2-p}{p}}}{L}\right)$ and $\eta=\Theta\left(\left[\frac{D_{\psi}(x,x^{1})}{M^{p}+\sigma^{p}}\right]^{\frac{1}{p}}\right)$, we get the desired bound. ∎
Appendix G Unified Theoretical Analysis under Sub-Weibull Noise
Lemma G.1 can be viewed as an any-time version of Lemma 4.1. However, this intriguing any-time property only holds for the minimizer $x^{*}$ unlike Lemma 4.1 being true for any $x\in\mathcal{X}$. Besides, the weight sequence $w_{t\in\left[T\right]}$ used in Lemma 4.1 is dropped. The reason is that we need an almost surely bounded random vector to use Lemma 6.1 (for the case $p=1$) to bound $\gamma_{t}v_{t-1}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle$. However, it is unclear whether $\left\|\gamma_{t}v_{t-1}(z^{t-1}-x^{t})\right\|$ is bounded. Hence, the weight sequence technique of Liu et al. (2023b) may fail.
Lemma G.1.
Under Assumptions 1.-3., suppose $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}$ and let $\gamma_{t\in\left[T\right]}\coloneqq\eta_{t}\prod_{s=2}^{t}\frac{1+\mu_{h}\eta_{s-1}}{1-\mu_{f}\eta_{s}}$ and $v_{t\in\left\{0\right\}\cup\left[T\right]}>0$ be defined as $v_{t}\coloneqq\frac{\gamma_{T}}{\sum_{s=t\lor 1}^{T}\gamma_{s}}$, then for any $s\in\left[T\right]$, we have
$\gamma_{s}v_{s}\left(F(x^{s+1})-F(x^{*})\right)+\gamma_{s+1}\eta_{s+1}^{-1}(1-\mu_{f}\eta_{s+1})v_{s}D_{\psi}(z^{s},x^{s+1})$
$\leq$ $(1-\mu_{f}\eta_{1})v_{0}D_{\psi}(x^{*},x^{1})+\sum_{t=1}^{s}2\gamma_{t}\eta_{t}v_{t}(M^{2}+\left\|\xi^{t}\right\|_{*}^{2})+\sum_{t=1}^{s}\gamma_{t}v_{t-1}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle,$
where $\xi^{t}\coloneqq\widehat{g}^{t}-\mathbb{\mathbb{E}}\left[\widehat{g}^{t}\mid\mathcal{F}^{t-1}\right],\forall t\in\left[T\right]$ and $z^{t}\coloneqq\frac{v_{0}}{v_{t}}x^{*}+\sum_{s=1}^{t}\frac{v_{s}-v_{s-1}}{v_{t}}x^{s},\forall t\in\left\{0\right\}\cup\left[T\right]$.
Proof.
The proof is similar to the proof of Lemma 4.1. Again, we use the sequence
$z^{t}=\begin{cases}\left(1-\frac{v_{t-1}}{v_{t}}\right)x^{t}+\frac{v_{t-1}}{v_{t}}z^{t-1}&t\in\left[T\right]\\ x^{*}&t=0\end{cases}\Leftrightarrow z^{t}=\frac{v_{0}}{v_{t}}x^{*}+\sum_{s=1}^{t}\frac{v_{s}-v_{s-1}}{v_{t}}x^{s},\forall t\in\left\{0\right\}\cup\left[T\right].$
Following the same step when proving Lemma 4.1, we have for any $t\geq 1$
$F(x^{t+1})-F(z^{t})\leq$ $(\eta_{t}^{-1}-\mu_{f})\frac{v_{t-1}}{v_{t}}D_{\psi}(z^{t-1},x^{t})-(\eta_{t}^{-1}+\mu_{h})D_{\psi}(z^{t},x^{t+1})$
$+2\eta_{t}(M^{2}+\left\|\xi^{t}\right\|_{*}^{2})+\frac{v_{t-1}}{v_{t}}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle.$ (66)
Multiplying both sides by $\gamma_{t}v_{t}$ (these two terms are non-negative) and summing up from $t=1$ to $s$ where $s\in\left[T\right]$, we obtain
$\sum_{t=1}^{s}\gamma_{t}v_{t}\left(F(x^{t+1})-F(z^{t})\right)\leq$ $(1-\mu_{f}\eta_{1})v_{0}D_{\psi}(x^{*},x^{1})-\gamma_{s}(\eta_{s}^{-1}+\mu_{h})v_{s}D_{\psi}(z^{s},x^{s+1})$
$+\sum_{t=1}^{s}2\gamma_{t}\eta_{t}v_{t}(M^{2}+\left\|\xi^{t}\right\|_{*}^{2})+\sum_{t=1}^{s}\gamma_{t}v_{t-1}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle.$ (67)
Next, by the convexity of $F$ and the definition of $z^{t}$, we can find
$\sum_{t=1}^{s}\gamma_{t}v_{t}\left(F(x^{t+1})-F(z^{t})\right)\geq$ $\gamma_{s}v_{s}\left(F(x^{s+1})-F(x^{*})\right)-\left(\sum_{\ell=1}^{s}\gamma_{\ell}\right)\left(v_{1}-v_{0}\right)\left(F(x^{1})-F(x^{*})\right)$
$+\sum_{t=2}^{s}\left[\gamma_{t-1}v_{t-1}-\left(\sum_{\ell=t}^{s}\gamma_{\ell}\right)\left(v_{t}-v_{t-1}\right)\right]\left(F(x^{t})-F(x^{*})\right).$
Note that $v_{1}=v_{0}$ and for $2\leq t\leq s$ there is
$\gamma_{t-1}v_{t-1}-\left(\sum_{\ell=t}^{s}\gamma_{\ell}\right)\left(v_{t}-v_{t-1}\right)$ $=\left(\sum_{\ell=t-1}^{s}\gamma_{\ell}\right)v_{t-1}-\left(\sum_{\ell=t}^{s}\gamma_{\ell}\right)v_{t}=\gamma_{T}\left(\frac{\sum_{\ell=t-1}^{s}\gamma_{\ell}}{\sum_{\ell=t-1}^{T}\gamma_{\ell}}-\frac{\sum_{\ell=t}^{s}\gamma_{\ell}}{\sum_{\ell=t}^{T}\gamma_{\ell}}\right)$
$=\frac{\gamma_{T}\gamma_{t-1}\left(\sum_{\ell=s+1}^{T}\gamma_{\ell}\right)}{\left(\sum_{\ell=t-1}^{T}\gamma_{\ell}\right)\left(\sum_{\ell=t}^{T}\gamma_{\ell}\right)}\geq 0.$
Hence, by using $F(x^{t})-F(x^{*})\geq 0$, we still have
$\sum_{t=1}^{s}\gamma_{t}v_{t}\left(F(x^{t+1})-F(z^{t})\right)\geq\gamma_{s}v_{s}\left(F(x^{s+1})-F(x^{*})\right).$ (68)
Combining (67) and (68), we conclude for any $s\in\left[T\right]$
$\gamma_{s}v_{s}\left(F(x^{s+1})-F(x^{*})\right)+\gamma_{s}(\eta_{s}^{-1}+\mu_{h})v_{s}D_{\psi}(z^{s},x^{s+1})$
$\leq$ $(1-\mu_{f}\eta_{1})v_{0}D_{\psi}(x^{*},x^{1})+\sum_{t=1}^{s}2\gamma_{t}\eta_{t}v_{t}(M^{2}+\left\|\xi^{t}\right\|_{*}^{2})+\sum_{t=1}^{s}\gamma_{t}v_{t-1}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle.$
The proof is finished by noticing $\gamma_{s}(\eta_{s}^{-1}+\mu_{h})=\gamma_{s+1}\eta_{s+1}^{-1}(1-\mu_{f}\eta_{s+1}),\forall s\in\left[T\right]$. ∎
As described above, our key problem now is to deal with the term $\sum_{t=1}^{s}\gamma_{t}v_{t-1}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle$ (the term $\sum_{t=1}^{s}\gamma_{t}\eta_{t}v_{t}\left\|\xi^{t}\right\|_{*}^{2}$ can be bounded by Lemma G.3). Inspired by Ivgi et al. (2023) and Liu and Zhou (2023), we will introduce another sequence $y_{t\in\left[T\right]}$ (see (70) for its definition). By a transformation, we only need to bound $\sqrt{y_{t}}\max_{\ell\in\left[s\right]}\left|\sum_{t=1}^{\ell}\gamma_{t}v_{t-1}\left\langle\xi^{t},\frac{z^{t-1}-x^{t}}{\sqrt{y_{t}}}\right\rangle\right|$ (this is done by Lemma C.2 in Ivgi et al. (2023)). With the carefully designed $y_{t}$, we will make sure that $\gamma_{t}v_{t-1}\frac{z^{t-1}-x^{t}}{\sqrt{y_{t}}}$ is bounded almost surely (the choice of $y_{t}$ is inspired by Liu and Zhou (2023), especially from their Theorem 6). Hence, the term $\left|\sum_{t=1}^{\ell}\gamma_{t}v_{t-1}\left\langle\xi^{t},\frac{z^{t-1}-x^{t}}{\sqrt{y_{t}}}\right\rangle\right|$ can be controlled in a high-probability way (see Lemma G.4). Moreover, we use an induction-based proof like Liu and Zhou (2023) to obtain the following general lemma. We refer the interested reader to our proof for details.
Note that Lemma G.2 can also be used to get the high-probability convergence results for strongly convex objectives. We omit that case for simplicity.
Lemma G.2.
Under Assumptions 1.-4. and 5D., suppose $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}$ and let $\gamma_{t\in\left[T\right]}\coloneqq\eta_{t}\prod_{s=2}^{t}\frac{1+\mu_{h}\eta_{s-1}}{1-\mu_{f}\eta_{s}}$, then for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
$F(x^{T+1})-F(x^{*})\leq\max_{2\leq t\leq T}\frac{2}{1-\mu_{f}\eta_{t}}\left[\frac{D_{\psi}(x^{*},x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+2\left(M^{2}+\sigma^{2}C(\delta,p)\right)\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}\right],$
where
$C(\delta,p)\coloneqq\begin{cases}\left(\frac{2e}{p}\lor e\log\frac{2e}{\delta}\right)^{\frac{2}{p}}+\left(4\sqrt{3}+3\sqrt{2}\right)^{2}\log^{\frac{2}{p}}\frac{4}{\delta}=O\left(1+\log^{\frac{2}{p}}\frac{1}{\delta}\right)&p\in\left[1,2\right)\\ \left(\frac{2e}{p}\lor e\log\frac{2e}{\delta}\right)^{\frac{2}{p}}+\frac{64\left[1\lor\log^{\frac{p+2}{p}}\left(4\left[3+\left(\frac{3}{p}\right)^{\frac{2}{p}}2\right]/\delta\right)\right]}{\log^{\frac{2}{p}}2}=O\left(1+\log^{1+\frac{2}{p}}\frac{1}{\delta}\right)&p\in\left(0,1\right)\end{cases}.$ (69)
Proof.
We first introduce the following sequence
$y_{t\in\left[T\right]}\coloneqq\max_{\ell\in\left[t\right]}\gamma_{\ell}\eta_{\ell}^{-1}\left(1-\mu_{f}\eta_{\ell}\mathds{1}\left[\ell\geq 2\right]\right)v_{\ell-1}D_{\psi}(z^{\ell-1},x^{\ell}).$ (70)
Now, fix a time $s\in\left[T\right]$, we invoke Lemma G.1 to get
$\gamma_{s}v_{s}\left(F(x^{s+1})-F(x^{*})\right)+\gamma_{s+1}\eta_{s+1}^{-1}(1-\mu_{f}\eta_{s+1})v_{s}D_{\psi}(z^{s},x^{s+1})$
$\leq$ $(1-\mu_{f}\eta_{1})v_{0}D_{\psi}(x^{*},x^{1})+\sum_{t=1}^{s}2\gamma_{t}\eta_{t}v_{t}(M^{2}+\left\|\xi^{t}\right\|_{*}^{2})+\sum_{t=1}^{s}\gamma_{t}v_{t-1}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle.$
We can bound
$\sum_{t=1}^{s}\gamma_{t}v_{t-1}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle=\sum_{t=1}^{s}\gamma_{t}v_{t-1}\left\langle\xi^{t},\frac{z^{t-1}-x^{t}}{\sqrt{y_{t}}}\right\rangle\sqrt{y_{t}}$
$\overset{(a)}{\leq}$ $2\sqrt{y_{s}}\max_{\ell\in\left[s\right]}\left|\sum_{t=1}^{\ell}\gamma_{t}v_{t-1}\left\langle\xi^{t},\frac{z^{t-1}-x^{t}}{\sqrt{y_{t}}}\right\rangle\right|\overset{(b)}{\leq}\frac{y_{s}}{2}+2\left(\max_{\ell\in\left[s\right]}\left|\sum_{t=1}^{\ell}\gamma_{t}v_{t-1}\left\langle\xi^{t},\frac{z^{t-1}-x^{t}}{\sqrt{y_{t}}}\right\rangle\right|\right)^{2},$
where $(a)$ is by Lemma C.2 of Ivgi et al. (2023), $(b)$ is due to AM-GM inequality. Hence, for any $s\in\left[T\right]$, the following inequality holds
$\gamma_{s}v_{s}\left(F(x^{s+1})-F(x^{*})\right)+\gamma_{s+1}\eta_{s+1}^{-1}(1-\mu_{f}\eta_{s+1})v_{s}D_{\psi}(z^{s},x^{s+1})$
$\leq$ $(1-\mu_{f}\eta_{1})v_{0}D_{\psi}(x^{*},x^{1})+\sum_{t=1}^{s}2\gamma_{t}\eta_{t}v_{t}(M^{2}+\left\|\xi^{t}\right\|_{*}^{2})+\frac{y_{s}}{2}+2\left(\max_{\ell\in\left[s\right]}\left|\sum_{t=1}^{\ell}\gamma_{t}v_{t-1}\left\langle\xi^{t},\frac{z^{t-1}-x^{t}}{\sqrt{y_{t}}}\right\rangle\right|\right)^{2}$
$\leq$ $(1-\mu_{f}\eta_{1})v_{0}D_{\psi}(x^{*},x^{1})+\sum_{t=1}^{T}2\gamma_{t}\eta_{t}v_{t}(M^{2}+\left\|\xi^{t}\right\|_{*}^{2})+\frac{y_{s}}{2}+2\left(\max_{\ell\in\left[T\right]}\left|\sum_{t=1}^{\ell}\gamma_{t}v_{t-1}\left\langle\xi^{t},\frac{z^{t-1}-x^{t}}{\sqrt{y_{t}}}\right\rangle\right|\right)^{2}.$ (71)
Next, we use induction to prove for any $s\in\left[T\right]$
$y_{s}\leq 2v_{0}D_{\psi}(x^{*},x^{1})+4\sum_{t=1}^{T}\gamma_{t}\eta_{t}v_{t}(M^{2}+\left\|\xi^{t}\right\|_{*}^{2})+4\left(\max_{\ell\in\left[T\right]}\left|\sum_{t=1}^{\ell}\gamma_{t}v_{t-1}\left\langle\xi^{t},\frac{z^{t-1}-x^{t}}{\sqrt{y_{t}}}\right\rangle\right|\right)^{2}.$ (72)
[…]
$\gamma_{s+1}\eta_{s+1}^{-1}(1-\mu_{f}\eta_{s+1})v_{s}D_{\psi}(z^{s},x^{s+1})$
$\leq$ $2v_{0}D_{\psi}(x^{*},x^{1})+\sum_{t=1}^{T}4\gamma_{t}\eta_{t}v_{t}(M^{2}+\left\|\xi^{t}\right\|_{*}^{2})+4\left(\max_{\ell\in\left[T\right]}\left|\sum_{t=1}^{\ell}\gamma_{t}v_{t-1}\left\langle\xi^{t},\frac{z^{t-1}-x^{t}}{\sqrt{y_{t}}}\right\rangle\right|\right)^{2},$
which gives us
$y_{s+1}$ $=y_{s}\lor\gamma_{s+1}\eta_{s+1}^{-1}(1-\mu_{f}\eta_{s+1})v_{s}D_{\psi}(z^{s},x^{s+1})$
$\leq 2v_{0}D_{\psi}(x^{*},x^{1})+\sum_{t=1}^{T}4\gamma_{t}\eta_{t}v_{t}(M^{2}+\left\|\xi^{t}\right\|_{*}^{2})+4\left(\max_{\ell\in\left[T\right]}\left|\sum_{t=1}^{\ell}\gamma_{t}v_{t-1}\left\langle\xi^{t},\frac{z^{t-1}-x^{t}}{\sqrt{y_{t}}}\right\rangle\right|\right)^{2}.$
Hence, the induction is completed.
Applying $s=T$ to (71) and (72), we have
$\gamma_{T}v_{T}\left(F(x^{T+1})-F(x^{*})\right)\leq(2-\mu_{f}\eta_{1})v_{0}D_{\psi}(x^{*},x^{1})+\sum_{t=1}^{T}4\gamma_{t}\eta_{t}v_{t}(M^{2}+\left\|\xi^{t}\right\|_{*}^{2})+4\left(\max_{s\in\left[T\right]}\left|\sum_{t=1}^{s}\gamma_{t}v_{t-1}\left\langle\xi^{t},\frac{z^{t-1}-x^{t}}{\sqrt{y_{t}}}\right\rangle\right|\right)^{2}.$
Using Lemmas G.3 and G.4, with probability at least $1-\delta$, there are
$\sum_{t=1}^{T}\gamma_{t}\eta_{t}v_{t}\left\|\xi^{t}\right\|_{*}^{2}\leq\sum_{t=1}^{T}\gamma_{t}\eta_{t}v_{t}\sigma^{2}\left(\frac{2e}{p}\lor e\log\frac{2e}{\delta}\right)^{\frac{2}{p}},$
and
$\left(\max_{s\in\left[T\right]}\left|\sum_{t=1}^{s}\gamma_{t}v_{t-1}\left\langle\xi^{t},\frac{z^{t-1}-x^{t}}{\sqrt{y_{t}}}\right\rangle\right|\right)^{2}\leq\max_{2\leq t\leq T}\frac{\widetilde{C}(\delta,p)}{1-\mu_{f}\eta_{t}}\sum_{t=1}^{T}\gamma_{t}\eta_{t}v_{t}\sigma^{2},$
where
$\widetilde{C}(\delta,p)\coloneqq\begin{cases}\left(4\sqrt{3}+3\sqrt{2}\right)^{2}\log^{\frac{2}{p}}\frac{4}{\delta}&p\in\left[1,2\right)\\ \frac{64\left[1\lor\log^{\frac{p+2}{p}}\left(4\left[3+\left(\frac{3}{p}\right)^{\frac{2}{p}}2\right]/\delta\right)\right]}{\log^{\frac{2}{p}}2}&p\in\left(0,1\right)\end{cases}.$
Thus, we obtain
$\gamma_{T}v_{T}\left(F(x^{T+1})-F(x^{*})\right)$
$\leq$ $(2-\mu_{f}\eta_{1})v_{0}D_{\psi}(x^{*},x^{1})+4\left[M^{2}+\sigma^{2}\left(\left(\frac{2e}{p}\lor e\log\frac{2e}{\delta}\right)^{\frac{2}{p}}+\max_{2\leq t\leq T}\frac{\widetilde{C}(\delta,p)}{1-\mu_{f}\eta_{t}}\right)\right]\sum_{t=1}^{T}\gamma_{t}\eta_{t}v_{t}$
$\leq$ $\max_{2\leq t\leq T}\frac{2}{1-\mu_{f}\eta_{t}}\left[v_{0}D_{\psi}(x^{*},x^{1})+2\left(M^{2}+\sigma^{2}C(\delta,p)\right)\sum_{t=1}^{T}\gamma_{t}\eta_{t}v_{t}\right].$
Dividing both sides by $\gamma_{T}v_{T}$ and plugging in $v_{t\in\left\{0\right\}\cup\left[T\right]}=\frac{\gamma_{T}}{\sum_{s=t\lor 1}^{T}\gamma_{s}}$, the proof is finished. ∎
The proof of the following lemma is inspired by Vladimirova et al. (2020).
Lemma G.3.
Under Assumption 5D., for any $\delta\in(0,1)$, with probability at least $1-\delta/2$, we have
$\sum_{t=1}^{T}\gamma_{t}\eta_{t}v_{t}\left\|\xi^{t}\right\|_{*}^{2}\leq\sum_{t=1}^{T}\gamma_{t}\eta_{t}v_{t}\sigma^{2}\left(\frac{2e}{p}\lor e\log\frac{2e}{\delta}\right)^{\frac{2}{p}},$
where $\gamma_{t}$, $\eta_{t}$, $v_{t}$ are the same as in Lemma G.1.
Proof.
Let $\Xi_{t}\coloneqq\gamma_{t}\eta_{t}v_{t}\left\|\xi^{t}\right\|_{*}^{2}$ and $\Lambda_{t}\coloneqq\gamma_{t}\eta_{t}v_{t}\sigma^{2}$ for $t\in\left[T\right]$. By Lemma A.2, for any $k\geq 1$, there is
$\mathbb{\mathbb{E}}\left[\left|\Xi_{t}\right|^{k}\mid\mathcal{F}^{t}\right]\leq e\Lambda_{t}^{k}\left(\frac{2k}{p}\right)^{\frac{2k}{p}}\Rightarrow\left\|\Xi_{t}\right\|_{k}\leq\Lambda_{t}e^{\frac{1}{k}}\left(\frac{2k}{p}\right)^{\frac{2}{p}},$
where $\left\|\cdot\right\|_{k}\coloneqq\left(\mathbb{\mathbb{E}}\left[\left|\cdot\right|^{k}\right]\right)^{\frac{1}{k}}$. Hence, for any $k\geq 1$,
$\left\|\sum_{t=1}^{T}\Xi_{t}\right\|_{k}\leq\sum_{t=1}^{T}\left\|\Xi_{t}\right\|_{k}\leq\left(\sum_{t=1}^{T}\Lambda_{t}\right)e^{\frac{1}{k}}\left(\frac{2k}{p}\right)^{\frac{2}{p}}.$
Let $k\coloneqq 1\lor\left(\frac{p}{2}\log\frac{2e}{\delta}\right)$ and $\lambda\coloneqq\left(\sum_{t=1}^{T}\Lambda_{t}\right)\left(\frac{2e}{\delta}\right)^{\frac{1}{k}}\left(\frac{2k}{p}\right)^{\frac{2}{p}}$, then by Markov’s inequality
$\Pr\left[\sum_{t=1}^{T}\gamma_{t}\eta_{t}v_{t}\left\|\xi^{t}\right\|_{*}^{2}>\lambda\right]=\Pr\left[\left(\sum_{t=1}^{T}\Xi_{t}\right)^{k}>\lambda^{k}\right]\leq\frac{\left\|\sum_{t=1}^{T}\Xi_{t}\right\|_{k}^{k}}{\lambda^{k}}\leq\frac{\delta}{2},$
which implies with probability at least $1-\delta/2$
$\sum_{t=1}^{T}\gamma_{t}\eta_{t}v_{t}\left\|\xi^{t}\right\|_{*}^{2}\leq\left(\sum_{t=1}^{T}\Lambda_{t}\right)\left(\frac{2e}{\delta}\right)^{\frac{1}{k}}\left(\frac{2k}{p}\right)^{\frac{2}{p}}\overset{(a)}{\leq}\left(\sum_{t=1}^{T}\Lambda_{t}\right)\left(\frac{2ek}{p}\right)^{\frac{2}{p}}=\sum_{t=1}^{T}\gamma_{t}\eta_{t}v_{t}\sigma^{2}\left(\frac{2e}{p}\lor e\log\frac{2e}{\delta}\right)^{\frac{2}{p}},$
where $(a)$ is due to $\frac{2e}{\delta}\leq\exp\left(\frac{2k}{p}\right)$. ∎
Lemma G.4 is stated for two cases, $p\in\left[1,2\right)$ and $p\in\left(0,1\right)$. To prove the first case, we only need to use Lemma 6.1. The other case of $p\in\left(0,1\right)$ is tricky. To deal with it, we need Theorem 1 in Li (2018).
Lemma G.4.
Under Assumptions 4. and 5D., for any $\delta\in(0,1)$, with probability at least $1-\delta/2$, for any $s\in\left[T\right]$, we have
$\left|\sum_{t=1}^{s}\gamma_{t}v_{t-1}\left\langle\xi^{t},\frac{z^{t-1}-x^{t}}{\sqrt{y_{t}}}\right\rangle\right|\leq\begin{cases}4\sqrt{3\log\frac{4}{\delta}\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}v_{t}\sigma^{2}}{1-\mu_{f}\eta_{t}\mathds{1}\left[t\geq 2\right]}}+3\sqrt{2}\left(\sum_{t=1}^{T}\left(\frac{\gamma_{t}\eta_{t}v_{t}\sigma^{2}}{1-\mu_{f}\eta_{t}\mathds{1}\left[t\geq 2\right]}\right)^{\frac{q}{2}}\right)^{\frac{1}{q}}\log^{\frac{1}{p}}\frac{4}{\delta}&p\in\left[1,2\right)\\ \frac{8}{\log^{\frac{1}{p}}2}\left[1\lor\log^{\frac{p+2}{2p}}\left(4\left[3+\left(\frac{3}{p}\right)^{\frac{2}{p}}2\right]/\delta\right)\right]\sqrt{\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}v_{t}\sigma^{2}}{1-\mu_{f}\eta_{t}\mathds{1}\left[t\geq 2\right]}}&p\in\left(0,1\right)\end{cases}.$
where $\gamma_{t}$, $\eta_{t}$, $v_{t}$ are the same as in Lemma G.1, $y_{t}\coloneqq\max_{\ell\in\left[t\right]}\gamma_{\ell}\eta_{\ell}^{-1}\left(1-\mu_{f}\eta_{\ell}\mathds{1}\left[\ell\geq 2\right]\right)v_{\ell-1}D_{\psi}(z^{\ell-1},x^{\ell})$, and $q\coloneqq\frac{p}{p-1}$ when $p\in\left[1,2\right)$.
Proof.
Let $Z^{t}\coloneqq\gamma_{t}v_{t-1}\frac{z^{t-1}-x^{t}}{\sqrt{y_{t}}},\forall t\in\left[T\right]$, our goal is to bound $\left|\sum_{t=1}^{s}\left\langle\xi^{t},Z^{t}\right\rangle\right|$. A useful observation is that
$\left\|Z^{t}\right\|$ $=\gamma_{t}v_{t-1}\frac{\left\|z^{t-1}-x^{t}\right\|}{\sqrt{y_{t}}}\leq\gamma_{t}v_{t-1}\sqrt{\frac{2D_{\psi}(z^{t-1},x^{t})}{\gamma_{t}\eta_{t}^{-1}\left(1-\mu_{f}\eta_{t}\mathds{1}\left[t\geq 2\right]\right)v_{t-1}D_{\psi}(z^{t-1},x^{t})}}$
$=\sqrt{\frac{2\gamma_{t}\eta_{t}v_{t-1}}{1-\mu_{f}\eta_{t}\mathds{1}\left[t\geq 2\right]}}\overset{v_{t-1}\leq v_{t}}{\leq}\sqrt{\frac{2\gamma_{t}\eta_{t}v_{t}}{1-\mu_{f}\eta_{t}\mathds{1}\left[t\geq 2\right]}}\coloneqq m_{t}.$ (73)
In the following, we split the proof into three cases, $p\in\left(1,2\right)$, $p=1$, and $p\in\left(0,1\right)$.
First, we consider the case $p\in\left(1,2\right)$. Let
$w\coloneqq\sqrt{\frac{\log\frac{4}{\delta}}{6\sigma^{2}\sum_{t=1}^{T}m_{t}^{2}}}\land\left(\frac{\log\frac{4}{\delta}}{2^{q-1}(q-1)\sigma^{q}\left(\sum_{t=1}^{T}m_{t}^{q}\right)}\right)^{\frac{1}{q}}.$ (74)
Now, we define the following non-negative sequence with $R_{0}\coloneqq 1$ and
$R_{s}\coloneqq\exp\left(\sum_{t=1}^{s}w\left\langle\xi^{t},Z^{t}\right\rangle-6w^{2}\sigma^{2}\left\|Z^{t}\right\|^{2}-2^{q-1}w^{q}\sigma^{q}\left\|Z^{t}\right\|^{q}\right)\in\mathcal{F}^{s},\forall s\in\left[T\right].$
We prove that $R_{t}$ is a supermartingale by
$\mathbb{\mathbb{E}}\left[R_{t}\mid\mathcal{F}^{t-1}\right]=R_{t-1}\mathbb{\mathbb{E}}\left[\exp\left(w\left\langle\xi^{t},Z^{t}\right\rangle-6w^{2}\sigma^{2}\left\|Z^{t}\right\|^{2}-2^{q-1}w^{q}\sigma^{q}\left\|Z^{t}\right\|^{q}\right)\mid\mathcal{F}^{t-1}\right]\overset{(a)}{\leq}R_{t-1},$
where $(a)$ is by applying Lemma 6.1 to $\mathbb{\mathbb{E}}\left[\exp\left(w\left\langle\xi^{t},Z^{t}\right\rangle\right)\mid\mathcal{F}^{t-1}\right]$. Then, we consider a stopping time
$\tau\coloneqq\min\left\{s\in\left[T\right]:R_{s}>\frac{4}{\delta}\right\},$
with the convention $\min\emptyset=\infty.$ Note that
$\Pr\left[\exists s\in\left[T\right],R_{s}>\frac{4}{\delta}\right]$ $=\Pr\left[\tau\leq T\right]=\mathbb{\mathbb{E}}\left[\mathds{1}\left[\tau\leq T\right]\right]<\frac{\delta}{4}\mathbb{\mathbb{E}}\left[R_{\tau}\mathds{1}\left[\tau\leq T\right]\right]$
$=\frac{\delta}{4}\mathbb{\mathbb{E}}\left[R_{\tau\land T}\mathds{1}\left[\tau\leq T\right]\right]\leq\frac{\delta}{4}\mathbb{\mathbb{E}}\left[R_{\tau\land T}\right]\overset{(b)}{\leq}\frac{\delta R_{0}}{4}=\frac{\delta}{4},$ (75)
where $(b)$ is due to the optional stopping theorem. Hence, with probability at least $1-\delta/4$, for any $s\in\left[T\right]$ there is $R_{s}\leq\frac{4}{\delta}$, which implies
$\sum_{t=1}^{s}\left\langle\xi^{t},Z^{t}\right\rangle$ $\leq\sum_{t=1}^{T}\left(6w\sigma^{2}\left\|Z^{t}\right\|^{2}+2^{q-1}w^{q-1}\sigma^{q}\left\|Z^{t}\right\|^{q}\right)+\frac{1}{w}\log\frac{4}{\delta}$
[…]
$\leq 2\sigma\sqrt{6\log\frac{4}{\delta}\sum_{t=1}^{T}m_{t}^{2}}+3\sigma\left(\sum_{t=1}^{T}m_{t}^{q}\right)^{\frac{1}{q}}\log^{\frac{1}{p}}\frac{4}{\delta},$
where the last step is due to $\max_{p\in\left(1,2\right)}2^{\frac{1}{p}}p(q-1)^{\frac{1}{q}}=3$. Similarly, we also have that, with probability at least $1-\delta/4$, for any $s\in\left[T\right]$,
$\sum_{t=1}^{s}\left\langle\xi^{t},Z^{t}\right\rangle\leq 2\sigma\sqrt{6\log\frac{4}{\delta}\sum_{t=1}^{T}m_{t}^{2}}+3\sigma\left(\sum_{t=1}^{T}m_{t}^{q}\right)^{\frac{1}{q}}\log^{\frac{1}{p}}\frac{4}{\delta}.$
Hence, we conclude that
$\Pr\left[\max_{s\in\left[T\right]}\left|\left\langle\xi^{t},Z^{t}\right\rangle\right|\leq 2\sigma\sqrt{6\log\frac{4}{\delta}\sum_{t=1}^{T}m_{t}^{2}}+3\sigma\left(\sum_{t=1}^{T}m_{t}^{q}\right)^{\frac{1}{q}}\log^{\frac{1}{p}}\frac{4}{\delta}\right]\geq 1-\frac{\delta}{2}.$ (76)
Next, we consider the case $p=1$. Let
$w\coloneqq\sqrt{\frac{\log\frac{4}{\delta}}{6\sigma^{2}\sum_{t=1}^{T}m_{t}^{2}}}\land\frac{1}{2\sigma\max_{t\in\left[T\right]}m_{t}}.$ (77)
We re-define the following non-negative sequence with $R_{0}\coloneqq 1$ and
$R_{s}\coloneqq\exp\left(\sum_{t=1}^{s}w\left\langle\xi^{t},Z^{t}\right\rangle-6w^{2}\sigma^{2}\left\|Z^{t}\right\|^{2}\right)\in\mathcal{F}^{s},\forall s\in\left[T\right].$
We prove that $R_{t}$ is a supermartingale by
$\mathbb{\mathbb{E}}\left[R_{t}\mid\mathcal{F}^{t-1}\right]=R_{t-1}\mathbb{\mathbb{E}}\left[\exp\left(w\left\langle\xi^{t},Z^{t}\right\rangle-6w^{2}\sigma^{2}\left\|Z^{t}\right\|^{2}\right)\mid\mathcal{F}^{t-1}\right]\overset{(c)}{\leq}R_{t-1},$
[…]
Similarly, we can show the same bound holds with probability at least $1-\delta/4$ for $\sum_{t=1}^{s}\left\langle\xi^{t},Z^{t}\right\rangle,\forall s\in\left[T\right]$. Therefore, we conclude
$\Pr\left[\max_{s\in\left[T\right]}\left|\sum_{t=1}^{s}\left\langle\xi^{t},Z^{t}\right\rangle\right|\leq 2\sigma\sqrt{6\log\frac{4}{\delta}\sum_{t=1}^{T}m_{t}^{2}}+3\sigma\max_{t\in\left[T\right]}m_{t}\log\frac{4}{\delta}\right]\geq 1-\frac{\delta}{2}.$ (78)
So, for $p\in\left[1,2\right)$, (76) and (78) together imply that, with probability at least $1-\delta/2$, for any $s\in\left[T\right]$ (where $\left(\sum_{t=1}^{T}m_{t}^{q}\right)^{\frac{1}{q}}=\max_{t\in\left[T\right]}m_{t}$ when $q=\infty\Leftrightarrow p=1$)
$\left|\sum_{t=1}^{s}\left\langle\xi^{t},Z^{t}\right\rangle\right|\leq 2\sigma\sqrt{6\log\frac{4}{\delta}\sum_{t=1}^{T}m_{t}^{2}}+3\sigma\left(\sum_{t=1}^{T}m_{t}^{q}\right)^{\frac{1}{q}}\log^{\frac{1}{p}}\frac{4}{\delta},$
which gives us (by (73))
$\left|\sum_{t=1}^{s}\gamma_{t}v_{t-1}\left\langle\xi^{t},\frac{z^{t-1}-x^{t}}{\sqrt{y_{t}}}\right\rangle\right|\leq 4\sqrt{3\log\frac{4}{\delta}\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}v_{t}\sigma^{2}}{1-\mu_{f}\eta_{t}\mathds{1}\left[t\geq 2\right]}}+3\sqrt{2}\left(\sum_{t=1}^{T}\left(\frac{\gamma_{t}\eta_{t}v_{t}\sigma^{2}}{1-\mu_{f}\eta_{t}\mathds{1}\left[t\geq 2\right]}\right)^{\frac{q}{2}}\right)^{\frac{1}{q}}\log^{\frac{1}{p}}\frac{4}{\delta}.$
For $p\in\left(0,1\right)$, by Theorem 1 in Li (2018), for any $z>0$
$\Pr\left[\exists s\in\left[T\right],\left|\sum_{t=1}^{s}\gamma_{t}v_{t-1}\left\langle\xi^{t},\frac{z^{t-1}-x^{t}}{\sqrt{y_{t}}}\right\rangle\right|>z\right]\leq 2\left[3+\left(\frac{3}{p}\right)^{\frac{2}{p}}\frac{64\sum_{t=1}^{T}\sigma_{t}^{2}}{z^{2}}\right]\exp\left(-\left(\frac{z^{2}}{32\sum_{t=1}^{T}\sigma_{t}^{2}}\right)^{\frac{p}{p+2}}\right),$
where $\sigma_{t\in\left[T\right]}\coloneqq\frac{m_{t}\sigma}{\log^{\frac{1}{p}}2}$. Thus, we take $z=u^{\frac{p+2}{2p}}\sqrt{32\sum_{t=1}^{T}\sigma_{t}^{2}}$ and $u=1\lor\log\left(4\left[3+\left(\frac{3}{p}\right)^{\frac{2}{p}}2\right]/\delta\right)$ to get
$\Pr\left[\exists s\in\left[T\right],\left|\sum_{t=1}^{s}\gamma_{t}v_{t-1}\left\langle\xi^{t},\frac{z^{t-1}-x^{t}}{\sqrt{y_{t}}}\right\rangle\right|>u^{\frac{p+2}{2p}}\sqrt{32\sum_{t=1}^{T}\sigma_{t}^{2}}\right]\leq 2\left[3+\left(\frac{3}{p}\right)^{\frac{2}{p}}\frac{2}{u^{\frac{p+2}{p}}}\right]\exp\left(-u\right)\leq\frac{\delta}{2},$
which gives us (by (73))
$\left|\sum_{t=1}^{s}\gamma_{t}v_{t-1}\left\langle\xi^{t},\frac{z^{t-1}-x^{t}}{\sqrt{y_{t}}}\right\rangle\right|\leq\frac{8}{\log^{\frac{1}{p}}2}\left[1\lor\log^{\frac{p+2}{2p}}\left(4\left[3+\left(\frac{3}{p}\right)^{\frac{2}{p}}2\right]/\delta\right)\right]\sqrt{\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}v_{t}\sigma^{2}}{1-\mu_{f}\eta_{t}\mathds{1}\left[t\geq 2\right]}}.$
Appendix H General Convex Functions under Sub-Weibull Noise
In this section, we provide the full version of Theorem 6.2.
Theorem H.1.
Under Assumptions 1.-4. and 5D. with $\mu_{f}=\mu_{h}=0$ and let $\delta\in(0,1)$:
If $T$ is unknown, by taking $\eta_{t\in\left[T\right]}=\frac{1}{2L}\land\frac{\eta}{\sqrt{t}}$ with $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x^{*},x^{1})}{M^{2}+\sigma^{2}C(\delta,p)}}\right)$, then with probability at least $1-\delta$, there is
$F(x^{T+1})-F(x^{*})\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{T}+\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x^{*},x^{1})}{\eta}+\eta\left(M^{2}+\sigma^{2}C(\delta,p)\right)\log T\right]\right).$
In particular, by choosing $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x^{*},x^{1})}{M^{2}+\sigma^{2}C(\delta,p)}}\right)$, there is
$F(x^{T+1})-F(x^{*})\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{T}+\frac{(M+\sigma\sqrt{C(\delta,p)})\sqrt{D_{\psi}(x^{*},x^{1})}\log T}{\sqrt{T}}\right).$
If $T$ is known, by taking $\eta_{t\in\left[T\right]}=\frac{1}{2L}\land\frac{\eta}{\sqrt{T}}$ with $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x^{*},x^{1})}{(M^{2}+\sigma^{2}C(\delta,p))\log T}}\right)$, then with probability at least $1-\delta$, there is
$F(x^{T+1})-F(x^{*})\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{T}+\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x^{*},x^{1})}{\eta}+\eta\left(M^{2}+\sigma^{2}C(\delta,p)\right)\log T\right]\right).$
In particular, by choosing $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x^{*},x^{1})}{(M^{2}+\sigma^{2}C(\delta,p))\log T}}\right)$, there is
$F(x^{T+1})-F(x^{*})\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{T}+\frac{(M+\sigma\sqrt{C(\delta,p)})\sqrt{D_{\psi}(x^{*},x^{1})\log T}}{\sqrt{T}}\right).$
Proof.
From Lemma G.2, if $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}$, with probability at least $1-\delta$, there is
$F(x^{T+1})-F(x^{*})\leq\left(\max_{2\leq t\leq T}\frac{2}{1-\mu_{f}\eta_{t}}\right)\left[\frac{D_{\psi}(x^{*},x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+2\left(M^{2}+\sigma^{2}C(\delta,p)\right)\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{t}}\right],$ (79)
where $\gamma_{t\in\left[T\right]}=\eta_{t}\prod_{s=2}^{t}\frac{1+\mu_{h}\eta_{s-1}}{1-\mu_{f}\eta_{s}}$. Note that $\mu_{f}=\mu_{h}=0$ now, hence, both $\eta_{t\in\left[T\right]}=\frac{1}{2L}\land\frac{\eta}{\sqrt{t}}$ and $\eta_{t\in\left[T\right]}=\frac{1}{2L}\land\frac{\eta}{\sqrt{T}}$ satisfy $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}=\frac{1}{2L}$. Besides, $\gamma_{t\in\left[T\right]}$ will degenerate to $\eta_{t\in\left[T\right]}$. Then we can simplify (79) into
$F(x^{T+1})-F(x^{*})\leq\frac{2D_{\psi}(x^{*},x^{1})}{\sum_{t=1}^{T}\eta_{t}}+4\left(M^{2}+\sigma^{2}C(\delta,p)\right)\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\sum_{s=t}^{T}\eta_{s}}.$ (80)
If $\eta_{t\in\left[T\right]}=\frac{1}{2L}\land\frac{\eta}{\sqrt{t}}$, similar to (27), we will have
$F(x^{T+1})-F(x^{*})\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{T}+\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x^{*},x^{1})}{\eta}+\eta\left(M^{2}+\sigma^{2}C(\delta,p)\right)\log T\right]\right).$
By plugging in $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x^{*},x^{1})}{M^{2}+\sigma^{2}C(\delta,p)}}\right)$, we get the desired bound.
If $\eta_{t\in\left[T\right]}=\frac{1}{2L}\land\frac{\eta}{\sqrt{T}}$, similar to (28), we will get
$F(x^{T+1})-F(x^{*})\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{T}+\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x^{*},x^{1})}{\eta}+\eta\left(M^{2}+\sigma^{2}C(\delta,p)\right)\log T\right]\right).$
By plugging in $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x^{*},x^{1})}{(M^{2}+\sigma^{2}C(\delta,p))\log T}}\right)$, we get the desired bound. ∎
H.1 Optimal Rate under Sub-Weibull Noise
In this section, we provide the full version of Theorem 6.3.
Theorem H.2.
Under Assumptions 1.-4. and 5D. with $\mu_{f}=\mu_{h}=0$ and let $\delta\in(0,1)$, if $T$ is known, by taking $\eta_{t\in\left[T\right]}=\frac{T-t+1}{2LT}\land\frac{\eta(T-t+1)}{T^{\frac{3}{2}}}$, then with probability at least $1-\delta$, there is
$F(x^{T+1})-F(x^{*})\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{T}+\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x^{*},x^{1})}{\eta}+\eta\left(M^{2}+\sigma^{2}C(\delta,p)\right)\right]\right).$
In particular, by choosing $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x^{*},x^{1})}{M^{2}+\sigma^{2}C(\delta,p)}}\right)$, there is
$F(x^{T+1})-F(x^{*})\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{T}+\frac{(M+\sigma\sqrt{C(\delta,p)})\sqrt{D_{\psi}(x^{*},x^{1})}}{\sqrt{T}}\right).$
Proof.
From Lemma G.2, if $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}$, with probability at least $1-\delta$, there is
$F(x^{T+1})-F(x^{*})\leq\left(\max_{2\leq t\leq T}\frac{2}{1-\mu_{f}\eta_{t}}\right)\left[\frac{D_{\psi}(x^{*},x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+2\left(M^{2}+\sigma^{2}C(\delta,p)\right)\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}\right],$ (81)
where $\gamma_{t\in\left[T\right]}=\eta_{t}\prod_{s=2}^{t}\frac{1+\mu_{h}\eta_{s-1}}{1-\mu_{f}\eta_{s}}$. Note that $\mu_{f}=\mu_{h}=0$ now, hence, $\eta_{t\in\left[T\right]}=\frac{T-t+1}{2LT}\land\frac{\eta(T-t+1)}{T^{\frac{3}{2}}}$ satisfies $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}=\frac{1}{2L}$. Besides, $\gamma_{t\in\left[T\right]}$ will degenerate to $\eta_{t\in\left[T\right]}$. Then we can simplify (81) into
$F(x^{T+1})-F(x^{*})\leq\frac{2D_{\psi}(x^{*},x^{1})}{\sum_{t=1}^{T}\eta_{t}}+4\left(M^{2}+\sigma^{2}C(\delta,p)\right)\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\sum_{s=t}^{T}\eta_{s}}.$ (82)
For $\eta_{t\in\left[T\right]}=\frac{T-t+1}{2LT}\land\frac{\eta(T-t+1)}{T^{\frac{3}{2}}}$, similar to (33), we will have
$F(x^{T+1})-F(x^{*})\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{T}+\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x^{*},x^{1})}{\eta}+\eta\left(M^{2}+\sigma^{2}C(\delta,p)\right)\right]\right).$
By plugging in $\eta=\Theta\left(\sqrt{\frac{D_{\psi}(x^{*},x^{1})}{M^{2}+\sigma^{2}C(\delta,p)}}\right)$, we get the desired bound. ∎